### v1

# A Framework of SO(3)-equivariant Non-linear Representation Learning and its Application to Electronic-Structure Hamiltonian Prediction

Abstract
We present both a theoretical and a methodological framework that addresses a critical challenge in applying deep learning to physical systems: the reconciliation of non-linear expressiveness with SO(3)-equivariance in predictions of SO(3)-equivariant quantities, such as the electronic-structure Hamiltonian. Inspired by covariant theory in physics, we address this problem by exploring the mathematical relationships between SO(3)-invariant and SO(3)-equivariant quantities and their representations. We first construct theoretical SO(3)-invariant quantities derived from the SO(3)-equivariant regression targets, and use these invariant quantities as supervisory labels to guide the learning of high-quality SO(3)-invariant features. Given that SO(3)-invariance is preserved under non-linear operations, the encoding process for invariant features can extensively utilize non-linear mappings, thereby fully capturing the non-linear patterns inherent in physical systems. Building on this foundation, we propose a gradient-based mechanism to induce SO(3)-equivariant encodings of various degrees from the learned SO(3)-invariant features. This mechanism can incorporate non-linear expressive capabilities into SO(3)-equivariant representations, while theoretically preserving their equivariant properties as we prove. Our approach offers a promising general solution to the critical dilemma between equivariance and non-linear expressiveness in deep learning methodologies. We apply our theory and method to the electronic-structure Hamiltonian prediction tasks, demonstrating state-of-the-art performance across six benchmark databases.
keywords:
Deep Learning Theory, Neural Networks, SO(3)-equivariance, Science for AI, AI for Science
1 Introduction
With the advantages in computational complexity and generalization capabilities, deep learning paradigm has vigorously driven advancements in physics research [1]. For example, in predicting electronic-structure Hamiltonians, traditional Density Functional Theory (DFT) methods [2, 3] suffer from computational complexities of $\mathcal{O}(N^{3})$. In contrast, deep learning approaches have significantly reduced this complexity to $\mathcal{O}(N)$ [4, 5, 6, 7], accelerating processes by thousands or even millions of times. This reduction in computational demand not only unlocks possibilities for analyzing extremely large atomic systems but also paves the way for efficient material simulation, discovery, and design that were previously unimaginable.
However, deep learning methods still face substantial challenges when processing physical systems. To align with fundamental physical laws, these methods must strictly adhere to symmetry principles. For example, physical quantities such as force fields and electronic-structure Hamiltonians must be equivariant under 3D rotational transformations, i.e., elements from the SO(3) group. Traditional deep learning methods [8, 9] typically approximated SO(3)-equivariance through rotational data augmentation. However, this approach does not provide the strict assurance of SO(3)-equivariance required for physical research. Besides, the calculation of physical quantities calls for ultra-high numerical accuracy. For example, it is typically required that the errors on Hamiltonian calculation being restricted within the sub-millielectron volts (sub-meV) scale, to ensure accurate derivation of downstream physical quantities and related material properties. This necessitates that the neural networks possess strong capabilities to express the complex non-linear mappings from atomic structures to the regression targets. However, existing deep learning methods are unable to simultaneously ensure strict SO(3)-equivariance and high numerical accuracy. The root cause of this problem lies in the conflict between SO(3)-equivariance and non-linear expressiveness: specifically, directly applying non-linear activation functions on SO(3)-equivariant features (with degree $l\geq 1$) leads to the loss of equivariance, while bypassing non-linear mappings severely restricts the network’s expressive capabilities and thereby lowering down the achievable accuracy. Although several efforts have been made to address this challenge—introducing gated activation mechanisms [7, 10, 11, 12], employing discrete spherical harmonic bases [13, 14, 15], and developing hybrid regression approaches [16]—they still fall short, to varying degrees, of fully reconciling strict adherence to SO(3)-equivariance with the requirements for high-precision generalization capability. As analyzed in previous literature [13], this issue is commonly found in machine learning tasks that demand both strict equivariance and fine-grained generalization performance.
To address this challenge, we make theoretical and methodological explorations on unifying strict SO(3)-equivariance with strong non-linear expressiveness within the realm of deep representation learning. We are inspired by the insight that invariant quantities in transformation often reflect the mathematical nature of physical laws and can induce other quantities with equivariant properties. For example, in special relativity [17], the spacetime interval between events is an invariant under Lorentz transformations. This invariance is foundational in formulating physical laws that involve covariant quantities. Furthermore, from the perspective of deep representation learning, the attribute of preserving invariance under non-linear operations is a significant advantage, due to its compatibility with non-linear expressive capabilities. Built upon these insights, we theoretically solve the equivariance-expressiveness dilemma by intensively exploring and making use of the intrinsic relationships between SO(3)-invariant and SO(3)-equivariant quantities and representations: we first dedicate efforts to learning high-quality SO(3)-invariant features with ample non-linear expressiveness, and subsequently, we derive SO(3)-equivariant non-linear representations from these SO(3)-invariant ones. Specifically:
First, we construct quantities that are theoretically SO(3)-invariant, namely $tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})$, where $\mathbf{Q}$ are the SO(3)-equivariant regression targets, $tr(\cdot)$ signifies the trace operation, and $\dagger$ denotes the conjugate transpose operation. A significant advantage of these SO(3)-invariant quantities lies in the fact that they are directly derived from the SO(3)-equivariant target labels and can serve as unique supervision labels for the effective learning of informative SO(3)-invariant features that capture the intrinsic symmetry properties of the mathematical structure of $\mathbf{Q}$ without requiring additional labeling resources.
Second, we propose a gradient-based mechanism to induce SO(3)-equivariant representations as well as the regression target $\mathbf{Q}$, from high-quality SO(3)-invariant features learned under the supervision of $tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})$. This mechanism can incorporate non-linear expressive capabilities into SO(3)-equivariant representations while preserving their equivariant properties, as we prove.
We develop our theory into a unified SO(3)-equivariant non-linear representation learning method, and apply it to the computation of electronic structure Hamiltonians, which are central to quantum physics and inherently complex, presenting substantial challenges for regression models. Our method significantly enhances the state-of-the-art performance on six benchmark Hamiltonian databases, demonstrating that our approach not only maintains strict SO(3)-equivariance but also unleashes powerful non-linear expressiveness, thereby boosting the prediction accuracy of Hamiltonians, providing an effective deep learning tool for efficient and accurate electronic structure calculations of materials.
Our theory and method for bridging symmetry and expressiveness of deep learning are general and can be easily applied on predicting other physical quantities such as forces, which also obey SO(3)-equivariance. Furthermore, it can also be extended to other application areas beyond physics. For instance, in some computer vision tasks, such as autonomous driving [18], where there is a high demand for robustness against coordinate transformations of cameras, the mainstream approach has been to achieve approximate SO(3)-equivariance through data augmentation. However, this does not guarantee absolute safety. Our work holds great potential to construct deep models with high generalization performance under the premise of strict SO(3)-equivariance, potentially revolutionizing techniques in such tasks.
2 The Equivariance-Expressiveness Dilemma in Existing Deep Learning Methods
In this section, we introduce existing efforts to harmonize the deep learning paradigm with SO(3)-equivariance, analyzing the dilemma between strict SO(3)-equivariance and non-linear expressive capabilities.
2.1 Approaches with Strict SO(3)-equivariance but Weak Expressiveness
Most existing methods have focused on achieving SO(3)-equivariance at the neural mechanism level by applying group theory-guaranteed SO(3)-equivariant operators [19, 20]. These operators, including linear scaling, element-wise sum, direct products, direct sums, and Clebsch-Gordan decomposition, have been used to build graph neural network architectures for tasks in 3D point cloud analysis [21], molecular dynamic simulation and property prediction [11, 10], as well as Hamiltonian prediction [22, 4, 7, 12]. However, these operators could not provide sufficient expressiveness to fit complex non-linear mappings. The majority of these operators are linear operations, and the only one (direct product) that introduces some non-linearity can only approximate a simple squaring function, and cannot be extended to more complex non-linear forms due to the huge computational burden and difficulty in convergence. Even worse, directly applying highly-expressive non-linear activation functions, such as $Sigmoid$, $Softmax$, and $SiLU$, to SO(3)-equivariant features may result in the loss of equivariance. This dilemma leads to an unsatisfactory trade-off between strict equivariance and non-linear expressiveness, ultimately reducing the achievable performance of these regression models. To alleviate this issue, these methods introduced a gated activation mechanism that processes SO(3)-invariant features ($l=0$) using non-linear activations and employs them as linear gating coefficients for SO(3)-equivariant features ($l\geq 1$), aiming to enhance their expressive power while maintaining strict equivariance. However, from the perspective of SO(3)-equivariant features, this mechanism is a combination of linear operations and cannot fundamentally increase their non-linear expressive capabilities.
2.2 Approaches with Remarkable Expressiveness but Approximate SO(3)-equivariance
Li et al. [6] proposed a local coordinate strategy to achieve SO(3)-equivariance. This strategy relied on the transformation relationship between local and global coordinates to achieve equivariance, without imposing restrictions on the neural network, allowing the use of network modules that possess non-linear expressiveness. However, this strategy, solely depending on coordinate transformations, only works under holistic rigid rotations of the atomic system. Without exploration of a neural mechanism for equivariance, when the local coordinates experience non-rigid perturbations such as thermal movements or bilayer twists, this strategy offers no guarantees on maintaining symmetry properties; Zitnick et al. [13] decomposed features into weighted sums of spherical harmonic basis functions, each corresponding to a discrete point on a unit sphere. Such discrete decomposition enables non-linear operations on the coefficients to enhance expressiveness. In subsequent developments [14, 15, 23], this approach has demonstrated a remarkable capacity to fit complex functions. However, it degenerates from continuous to discrete rotational equivariance, no longer maintaining strict equivariance to continuous rotational operations; Yin et al. [16] proposed a hybrid framework consisting of both group theory-guaranteed SO(3)-equivariant mechanisms and non-linear mechanisms to regress Hamiltonians. In this framework, the non-linear mechanisms showed remarkable capability at learning SO(3)-equivariance from the data with the help of the theoretically SO(3)-equivariant mechanisms, and released powerful non-linear expressive capabilities to achieve more numerical accuracy. However, the equivariance achieved through data-driven methods does not have strict theoretical guarantees. In summary, although these methods with approximate SO(3)-equivariance achieve high numerical accuracy due to their extensive use of non-linear mechanisms, they are best suited for the applications that primarily require numerical approximations but not for tasks that demand theoretical depth in physics. In some applications, the demands for symmetry are extremely high. Even minor deviations from perfect symmetry can result in incorrect physical results.
Combining the analysis in this section, existing deep learning methods have compromised either strict SO(3)-equivariance or non-linear expressiveness when processing physical systems. In this paper, we develop a novel framework that theoretically combines strict SO(3)-equivariance with the non-linear expressiveness of neural networks. This solution resolves the equivariance-expressiveness dilemma by deeply exploring the intrinsic relationships between SO(3)-invariant and SO(3)-equivariant quantities.
3 Problem Formalization
Physical quantities that are SO(3)-equivariant typically possess their respective equivariant formulas under both matrix-formed direct-product states and vector-formed direct-sum states, which are mathematically equivalent to each other.
Without loss of generality, let ${\bf{Q}}^{l_{p}\otimes l_{q}}$ denote an SO(3)-equivariant quantity in direct-product state formed by $l_{p}\otimes l_{q}$, i.e., the direct product between degrees $l_{p}$ and $l_{q}$. It obeys the following SO(3)-equivariant law:
[…]
where $\dagger$ denotes the conjugate transpose operation. ${\bf{R}}\in\mathbb{R}^{3\times 3}$ is the rotational matrix and ${\bf{D}}^{l_{p}}({\bf{R}})\in\mathbb{R}^{(2l_{p}+1)\times(2l_{p}+1)}$, ${\bf{D}}^{l_{q}}({\bf{R}})\in\mathbb{R}^{(2l_{q}+1)\times(2l_{q}+1)}$ are the Wigner-D matrices of degrees $l_{p}$ and $l_{q}$, respectively, ${\bf{Q}}({\bf{R}})^{l_{p}\otimes l_{q}}\in\mathbb{R}^{(2l_{p}+1)\times(2l_{q}+1)}$ denotes the transformed results of ${\bf{Q}}^{l_{p}\otimes l_{q}}\in\mathbb{R}^{(2l_{p}+1)\times(2l_{q}+1)}$ through the rotational operation by ${\bf{R}}$.
${\bf{Q}}^{l_{p}\otimes l_{q}}$ in the direct-product state can be further decomposed into a series of direct-sum state components with vector forms, i.e., ${\bf{q}}^{l}(|l_{p}-l_{q}|\leq l\leq l_{p}+l_{q})$, which follows SO(3)-equivariant law mathematically equivalent to Eq. (1) but with a simpler form:
${\bf{q}}({\bf{R}})^{l}={\bf{D}}^{l}({\bf{R}})\cdot{\bf{q}}^{l},|l_{p}-l_{q}|\leq l\leq l_{p}+l_{q}$ (2)
where ${\bf{q}}^{l}\in\mathbb{R}^{2l+1}$ and ${\bf{q}}({\bf{R}})^{l}\in\mathbb{R}^{2l+1}$ respective denote the components with degree $l$ before and after the rotational operation by ${\bf{R}}$.
Eq. (1) and Eq. (2) are mathematical equivalent to each other, meaning that fulfilling one under the corresponding state naturally satisfies the other. Due to the simplicity and ease of handling of the direct-sum state form, existing deep learning methods [6, 7, 16] typically first predict $\mathbf{q}^{l}(|l_{p}-l_{q}|\leq l\leq l_{p}+l_{q})$, then convert it to $\mathbf{Q}^{l_{p}\otimes l_{q}}$ through Clebsch-Gordan matrices. In the direct-sum state, to obey the equivariant law of Eq. (2) for the regression target, the hidden representations of neural networks must also satisfy the same form of equivariance:
${\bf{f}}({\bf{R}})^{(k)l}={\bf{D}}^{l}({\bf{R}})\cdot{\bf{f}}^{(k)l}$ (3)
where ${\bf{f}}^{(k)l}\in\mathbb{R}^{2l+1}$ and ${\bf{f}}({\bf{R}})^{(k)l}\in\mathbb{R}^{2l+1}$ respectively denote one channel of features with degree $l$ before and after the rotational operation by ${\bf{R}}$, at the $k$ th hidden layer.
Due to the intrinsic complexity and non-linearity of physic quantities, neural networks on regressing these quantities are supposed to equip with non-linear mappings denoted as $g_{nonlin}(\cdot)$ to fully capture the intrinsic patterns of the physical quantities, which is crucial for precise and generalizable prediction performance. Meanwhile, $g_{nonlin}(\cdot)$ must also preserve SO(3)-equivariance, which is expressed as:
${\bf{f}}({\bf{R}})^{(k+1)l}={\bf{D}}^{l}({\bf{R}})\cdot{\bf{f}}^{(k+1)l},{\bf{f}}^{(k+1)l}=g_{nonlin}({\bf{f}}^{(k)l})$ (4)
However, directly implementing $g_{nonlin}(\cdot)$ as neural network module with non-linear mechanisms, such as $Sigmoid$, $Softmax$ and $Silu$, may result in the destruction of strict equivariance. How to make $g_{nonlin}(\cdot)$ both theoretically SO(3)-equivariant and capable of non-linear expressiveness, is the core problem this paper aims at solving.
4 Theory
This section establishes a rigorous theoretical foundation for integrating SO(3)-equivariance with non-linear expressiveness in the deep learning paradigm. First, we propose a theoretical construct to derive SO(3)-invariant quantities from the SO(3)-equivariant targets, aiming at facilitating the learning of high-quality SO(3)-invariant features. Second, we propose a gradient-based mechanism encoding features maintaining theoretically SO(3)-equivariance while possessing the non-linear expressive capability induced from SO(3)-invariant features to capture the complex non-linear patterns inherent in physical quantities.
Theorem 1.
The quantity $\mathbf{T}=tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})$ is SO(3)-invariant, where $\mathbf{Q}$ is the simplified representation (without superscripts) of $\mathbf{Q}^{l_{p}\otimes l_{q}}$ defined in Section 3, and $\dagger$ denotes the conjugate transpose operation, $tr(\cdot)$ is the trace operation.
Proof.
Under an SO(3) rotation represented by the rotational matrix $\mathbf{R}$, $\mathbf{Q}=\mathbf{Q}^{l_{p}\otimes l_{q}}$ is transformed as $\mathbf{Q}(\mathbf{R})$:
$\mathbf{Q}(\mathbf{R})=\mathbf{D}^{l_{p}}(\mathbf{R})\cdot\mathbf{Q}\cdot\mathbf{D}^{l_{q}}(\mathbf{R})^{\dagger},$
where $\mathbf{D}^{l_{p}}(\mathbf{R})$ and $\mathbf{D}^{l_{q}}(\mathbf{R})$ are the Wigner-D matrices for the respective orbital degrees of atoms $i$ and $j$, corresponding to the rotation $\mathbf{R}$.
The conjugate transpose of the transformed quantity is:
$\mathbf{Q}(\mathbf{R})^{\dagger}=\mathbf{D}^{l_{q}}(\mathbf{R})\cdot\mathbf{Q}^{\dagger}\cdot\mathbf{D}^{l_{p}}(\mathbf{R})^{\dagger}.$
Thus, the product $\mathbf{Q}(\mathbf{R})\cdot\mathbf{Q}(\mathbf{R})^{\dagger}$ is:
$\mathbf{Q}(\mathbf{R})\cdot\mathbf{Q}(\mathbf{R})^{\dagger}=(\mathbf{D}^{l_{p}}(\mathbf{R})\cdot\mathbf{Q}\cdot\mathbf{D}^{l_{q}}(\mathbf{R})^{\dagger})\cdot(\mathbf{D}^{l_{q}}(\mathbf{R})\cdot\mathbf{Q}^{\dagger}\cdot\mathbf{D}^{l_{p}}(\mathbf{R})^{\dagger}).$
Using the cyclic property of the trace, which states that the trace of a product of matrices remains unchanged under cyclic permutations (i.e., $tr(ABC)=tr(BCA)=tr(CAB)$), and combining the properties that $\mathbf{D}^{l_{p}}(\mathbf{R})\cdot\mathbf{D}^{l_{p}}(\mathbf{R})^{\dagger}=\mathbf{I}$ and $\mathbf{D}^{l_{q}}(\mathbf{R})\cdot\mathbf{D}^{l_{q}}(\mathbf{R})^{\dagger}=\mathbf{I}$, we can rearrange the terms inside the trace as follows:
$\mathbf{T}(\mathbf{R})=tr(\mathbf{Q}(\mathbf{R})\cdot\mathbf{Q}(\mathbf{R})^{\dagger})=tr((\mathbf{D}^{l_{p}}(\mathbf{R})\cdot\mathbf{Q}\cdot\mathbf{D}^{l_{q}}(\mathbf{R})^{\dagger})\cdot(\mathbf{D}^{l_{q}}(\mathbf{R})\cdot\mathbf{Q}^{\dagger}\cdot\mathbf{D}^{l_{p}}(\mathbf{R})^{\dagger}))$
$=tr(\mathbf{D}^{l_{p}}(\mathbf{R})\cdot\mathbf{Q}\cdot\mathbf{Q}^{\dagger}\cdot\mathbf{D}^{l_{p}}(\mathbf{R})^{\dagger})=tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger}\cdot\mathbf{D}^{l_{p}}(\mathbf{R})^{\dagger}\cdot\mathbf{D}^{l_{p}}(\mathbf{R}))=tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})=\mathbf{T}.$
Therefore, $\mathbf{T}=tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})$ is invariant under SO(3) transformations, its SO(3)-invariance is proved. ∎
Theorem 2.
The non-linear neural mapping defined as the following is SO(3)-equivariant:
$\mathbf{v}=g_{nonlin}(\mathbf{f})=\frac{\partial(s_{nonlin}(\text{CGDecomp}(\mathbf{f}\otimes\mathbf{f},0)))}{\partial\mathbf{f}}$ (5)
where $\mathbf{f}$ is an input SO(3)-equivariant feature with degree $l$ in the direct-sum state, $\otimes$ denotes the direct product operation, $\text{CGDecomp}(\cdot,0)$ refers to performing a Clebsch-Gordan decomposition of the tensor and returning the scalar component of degree 0, and $s_{nonlin}(\cdot)$ represents non-linear neural mappings, $\mathbf{v}$ is the features encoded by $g_{nonlin}(\cdot)$
Proof.
The input feature $\mathbf{f}$ is SO(3)-equivariant, meaning that under an SO(3) rotation represented by $\mathbf{R}$, it transforms as follows:
$\mathbf{f}(\mathbf{R})=\mathbf{D}^{l}(\mathbf{R})\cdot\mathbf{f}$
where $\mathbf{D}^{l}(\mathbf{R})$ is the Wigner-D matrix corresponding to degree $l$.
First, according to group theory, $u=\text{CGDecomp}(\mathbf{f}\otimes\mathbf{f},0)$ is an SO(3)-invariant scalar because the degree-zero component from the Clebsch-Gordan decomposition is invariant under rotations. Since applying a non-linear operation to a SO(3)-invariant quantity does not change its invariance, $z=s_{\text{nonlin}}(u)$ is also SO(3)-invariant, independent to the specific form of $s_{\text{nonlin}}(\cdot)$. It formally holds that:
$z=z(\mathbf{R})$ (6)
Next, we apply the chain rule in Jacobian form. Considering $\mathbf{f}(\mathbf{R})$ is a column vector, to facilitate the application of the chain rule in vector form, we first transpose it into a row vector $\mathbf{f}(\mathbf{R})^{T}$, then differentiate:
$\frac{\partial z(\mathbf{R})}{\partial\mathbf{f}^{T}(\mathbf{R})}=\frac{\partial z}{\partial\mathbf{f}^{T}(\mathbf{R})}=\frac{\partial z}{\partial\mathbf{f}^{T}}\frac{\partial\mathbf{f}^{T}}{\partial\mathbf{f}(\mathbf{R})^{T}}=\frac{\partial z}{\partial\mathbf{f}^{T}}\cdot\mathbf{D}^{l}(\mathbf{R})^{-1}=\frac{\partial z}{\partial\mathbf{f}^{T}}\cdot\mathbf{D}^{l}(\mathbf{R})^{T}$ (7)
Here we utilize the property that $\mathbf{D}^{l}(\mathbf{R})^{-1}=\mathbf{D}^{l}(\mathbf{R})^{T}$. Since the representations of neural networks are generally real numbers, the Wigner-D matrix is also real unitary.
Finally, we transpose the result back to a column vector:
$\mathbf{v}(\mathbf{R})=g_{nonlin}(\mathbf{f}(\mathbf{R}))=(\frac{\partial z(\mathbf{R})}{\partial\mathbf{f}^{T}(\mathbf{R})})^{T}=(\frac{\partial z}{\partial\mathbf{f}^{T}}\cdot\mathbf{D}^{l}(\mathbf{R})^{T})^{T}=\mathbf{D}^{l}(\mathbf{R})\cdot\frac{\partial z}{\partial\mathbf{f}}=\mathbf{D}^{l}(\mathbf{R})\cdot\mathbf{v}$ (8)
This proves that $g_{\text{nonlin}}(\cdot)$ is a SO(3)-equivariant non-linear operator: when applying its non-linearity to a SO(3)-equivariant feature $\mathbf{f}$, the output feature $\mathbf{v}$ remains SO(3)-equivariant. ∎
Remark 1.
In Theorem 1, the term $tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})$ is equal to the square of the Frobenius norm of the matrix $\mathbf{Q}$, i.e., $tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})=\|\mathbf{Q}\|_{F}^{2}=\sum_{m,n}|\mathbf{Q}_{mn}|^{2}$ The invariance of Frobenius norm under unitary transformations reflects the profound connections between modulus and invariance in physics.
Remark 2.
In Theorem 2, $g_{\text{nonlin}}(\cdot)$ can handle equivariant features of any degree and the returned feature $\mathbf{v}$ has the same degree $l$ as the input feature $\mathbf{f}$. $\mathbf{v}$ is endowed with non-linear expressiveness with the effect of $s_{nonlin}(\cdot)$, which can be any differentiable non-linear network module, including non-linear activation functions and normalization operations. The gradient operation $\frac{\partial z}{\partial\mathbf{f}}$ in $g_{\text{nonlin}}(\cdot)$ is used during the forward inference process to promote feature encoding, expanding the usage and significance of gradients within the neural network paradigm. It is worth noting that, although the internal representations of neural network modules including $g_{\text{nonlin}}(\cdot)$ are in the real field, they are still able to predict targets in complex field with equivariance. Literature [7] has provided mechanisms for converting the network outputs in the real field into regression targets with real and imaginary parts.
Remark 3.
$z$ in Equation (6) is the bridge connecting Theorem 1 and Theorem 2 in implementing a SO(3)-equivariant non-linear representation learning approach. Since $z$ is an SO(3)-invariant feature, it can be learned by $s_{\text{nonlin}}(\cdot)$, i.e., the neural mechanisms with non-linear expressive capability, through supervised learning incorporated with the SO(3)-invariant label $tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})$. Following this, by taking the gradient of $z$ with respect to $\mathbf{f}$, the non-linear expressiveness of $z$ is transferred to the equivariant feature $\mathbf{v}$, while maintaining strict SO(3)-equivariance as proved above. Subsequently, merging $\mathbf{v}$ and $\mathbf{f}$, and applying $g_{\text{nonlin}}(\cdot)$ in a stacked manner can yield rich SO(3)-equivariant non-linear representations.
Figure 1: Methodological framework for learning SO(3)-equivariant representations with non-linear expressiveness to regress complex SO(3)-equivariant targets.
Based on the new theory, we present a general method for learning non-linear representations that are equivariant to SO(3), as shown in Figure 1. The core of our method can be abstractly referred to as TraceGrad. At the label level, it incorporates an SO(3)-invariant trace quantity introduced in Theorem 1 as a supervisory signal for learning SO(3)-invariant features; at the representation level, it induces non-linear SO(3)-equivariant features from non-linear SO(3)-invariant features through gradient-based operator introduced in Theorem 2. For details of the method, please refer to Section 8.
5 Application to the Hamiltonian Prediction Tasks
The tasks of predicting electronic-structure Hamiltonian are chosen as the application study of our theoretical and methodological framework, on two motivations. On one hand, the Hamiltonian occupies a central position in quantum mechanics. Form the Hamiltonian, key physical quantities such as charge density, band structure, and Berry phase can be calculated, bringing deeply understanding of the electrical, optical, magnetic, and transport properties of atomic systems. On the other hand, the intrinsic symmetry with respect to the SO(3) group as well as the high-dimensional complexity (as shown in Figure 2) of the Hamiltonian provide a challenging touchstone of whether a deep learning method possesses both SO(3)-equivariance and strong expressive power.
Figure 2: Illustration of the high-dimensional complexity of the Hamiltonian. In an atomic system, a Hamiltonian matrix is defined between each pair of atoms within each other’s cutoff radius, constituting a large cluster of Hamiltonian matrices for all atom pairs. Besides, each Hamiltonian matrix contains multiple basic blocks, with the $\mathbf{Q}^{l_{p}\otimes l_{q}}$ defined in Section 3 corresponding to a basic block $\mathbf{H}^{l_{p}\otimes l_{q}}$ here. Thus, for a Hamiltonian prediction model, the space of the regression target is typically huge.
Experiments are conducted on six benchmark databases of atomic systems that have Hamiltonian annotations, all of which are established and released by previous literature [6, 7]. These databases include:
• Monolayer Graphene (MG). This structure consists of a single layer of carbon atoms in a hexagonal lattice. It is known for its exceptional thermal conductivity, electrical properties, and mechanical strength. The potential applications of monolayer graphene extend across electronics, energy storage devices, and composite materials due to its unique properties. In the benchmark of MG, the sizes of the training set, validation set, and testing set are 270, 90, and 90 respectively.
• Monolayer MoS2 (MM), which is a single layer of molybdenum disulfide and part of the transition metal dichalcogenides (TMDCs) group. This kind of structures is a semiconductor with a direct bandgap, making it suitable for optoelectronics, photovoltaics, and semiconducting components. It also shows promise in applications requiring high sensitivity to chemicals, making it valuable for sensors. In the benchmark of MM, the sizes of the training set, validation set, and testing set are 300, 100, and 100 respectively.
• Bilayer Graphene (BG), which consists of two layers of graphene stacked together. This kind of structures exhibits unique electronic properties such as tunable bandgap and enhanced conductivity when subjected to an electric field. Bilayer Graphene is being explored for use in advanced electronic devices, such as fast transistors and efficient batteries. In the benchmark database of BG, the sizes of the training set and validation set are 180 and 60 respectively, all of them are non-twisted samples; in the testing set, the numbers of non-twisted and twisted samples are 60 and 9 respectively.
• Bilayer Bismuthene (BB), which consists of two layers of bismuth atoms arranged in a honeycomb-like structure. This kind of materials shows potential for topological insulating properties and has significant implications for quantum computing and spintronic devices due to its strong spin-orbit coupling. In the benchmark of BB, the sizes of the training set and validation set are 231 and 113 respectively, all of them are non-twisted samples; in the testing set, the numbers of non-twisted and twisted samples are 113 and 4 respectively.
• Bilayer Bi2Te3 (BT), which is a compound of bismuth and tellurium, known for its thermoelectric properties. This kind of structures is particularly valuable in thermoelectric devices for power generation and refrigeration technologies where waste heat can be converted into electrical energy. In the benchmark of BT, the sizes of the training set and validation set are 204 and 38 respectively, all of them are non-twisted samples; in the testing set, the numbers of non-twisted and twisted samples are 12 and 2 respectively.
• Bilayer Bi2Se3 (BS), another bismuth-tellurium compound, similar to Bi2Te3 but with selenium replacing tellurium. It also exhibits excellent thermoelectric properties and potential for topological insulator applications. This makes it suitable for energy harvesting applications and advanced electronic devices based on topological insulator technology. In the benchmark of BS, the sizes of the training set and validation set are 231 and 113 respectively, all of them are non-twisted samples; in the testing set, the numbers of non-twisted and twisted samples are 113 and 2 respectively.
These databases are both diverse and representative. They encompass both strong chemical bonds between single-layer atoms and weaker van der Waals interactions between bilayers, and also cover varying degrees of Spin-Orbit Coupling (SOC) effects: structures like BT, BB, and BS exhibit strong SOC effects, while others have weak SOC. These atomic structures are highly valuable in the field of information and energy science and technology. Yet, accurately predicting their Hamiltonians poses significant challenges due to structural deformations from thermal movements and inter-layer twists, as depicted in Fig. 3, where the twisted structures have garnered considerable attention as a research focus because of their potential to exhibit novel electrical properties and quantum topological behaviors [24, 25]. During twisting, the relative rotation between atoms and their coordinate system introduces SO(3)-equivariant effects, while changes in orientation between two layers alter the vdW interactions. These dynamic conditions challenge both the equivariance and expressive capabilities of the regression model. In our experiments, predicting the Hamiltonian of twisted samples are even challenging, as we strictly adhere to the official division of training, validation, and testing sets as defined upon the dataset’s release [^1] https://github.com/Xiaoxun-Gong/DeepH-E3. In this division, twisted samples are only included in the testing set and are absent from the training set. This setup forces the regression model to solely rely on its generalization ability and the inherent SO(3)-equivariance to extrapolate the twisting conditions.
In summary, these benchmarks present scientific challenges for a regression model, pushing the boundaries of modeling capabilities in predicting complex quantum behaviors under dynamic conditions.
Figure 3: Illustration of challenging samples in the experimental databases, showing structural distortions of atomic systems due to thermal movements and twists between the two layers, which require a regression model to possess both strict SO(3)-equivariance and strong expressive capacity.
6 Results and Analysis
To validate the effectiveness of the proposed method, we compare results from the four setups:
• Baseline model (DeepHE3 [7]): We select DeepHE3 as our baseline model since it represents the state-of-the-art (SOTA) among strictly SO(3)-equivariant approaches for predicting electronic-structure Hamiltonians. This choice facilitates further breakthroughs based on the SOTA foundation. DeepHE3 is trained using the mean squared error (MSE) loss of the Hamiltonian. In the subsequent experimental setups, DeepHE3 will serve as a base encoder, providing basic SO(3)-equivariant features that are handled by our method as more expressive representations.
• DeepHE3+Trace: This experimental setup, an ablation term, only implements half part of our method. Specifically, it extends the architecture of DeepHE3 by adding our SO(3)-invariant encoding and decoding branches and using the trace quantity $\mathbf{T}=\text{tr}(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})=\text{tr}(\mathbf{H}^{l_{p}\otimes l_{q}}\cdot(\mathbf{H}^{l_{p}\otimes l_{q}})^{\dagger})$ to train them. As for ablation study, this setup does not include mechanisms for encoding non-linear SO(3)-equivariant features via the gradient-induced operator; instead, it directly uses the features outputted by DeepHE3 for Hamiltonian regression.
• DeepHE3+Grad: This setup is also an ablation term and implements the other half part of our method in contrast to the previous ablation term. Specifically, it incorporates our SO(3)-invariant encoder branch as well as the gradient-induced operator to yield SO(3)-equivariant non-linear features based on the representations provided by the base encoders of DeepHE3 for Hamiltonian prediction. As for ablation study, this setup continues to use the single-task training pipeline of DeepHE3, predicting only the Hamiltonian without joint supervision training through the trace label.
• DeepHE3+TraceGrad. This setup corresponds to a complete implementation of our framework. Extending beyond the baseline architecture and pipeline of DeepHE3, at the label level, we introduce the trace quantity to guide the learning of SO(3)-invariant features; Meanwhile, at the representation level, we leverage non-linear SO(3)-invariant features induced by the gradient operator to yield non-linear SO(3)-equivariant features for Hamiltonian prediction.
The performance of interest for the electronic-structure Hamiltonian prediction task are SO(3)-equivariance and numerical accuracy. Our method has already been theoretically proven to strictly obey SO(3)-equivariance in Section 4; therefore, we just focus on evaluating numerical accuracy here. Four metrics designed by Yin et al. [16] are adopted, i.e., $MAE_{all}$, which calculates the Mean Absolute Error (MAE) averaged across all elements of the Hamiltonian matrices from all testing samples in the current database; $MAE^{cha}_{sample}$, which calculates the MAE on the most challenging testing samples where the baseline model (DeepHE3) performs the worst in the current database; $MAE_{block}$, which calculates the MAE for each basic block of the Hamiltonian matrix, averaged across all testing samples in the current database; $MAE^{cha}_{block}$, which corresponds to the MAE on the most challenging Hamiltonian block where the baseline model shows the poorest performance (with the largest $MAE_{block}$) in the current dataset. These metrics comprehensively reflect the accuracy performance, considering not only the average accuracy but also the accuracy on difficult samples and challenging blocks of the Hamiltonian matrices.
We collect results from $10$ independently repeated experiments on each database and list them in Tables 1 and 2, including the values of $MAE_{all}$, $MAE^{cha}_{sample}$, and $MAE^{cha}_{block}$, for monolayer and bilayer structures, respectively. Considering the large volume of statistics on $MAE_{block}$ for all blocks, we just display the complete results in Figures 4 and 5 with the form of bar charts for DeepHE3 and DeepHE3+TraceGrad. The results for DeepHE3 listed in these tables and figures are directly taken from the reproduction of Yin et al. [16], a paper which is in fact previous work by our own team. Since the experimental conditions for DeepHE3 are exactly the same between the previous and current work of ours, we directly list the results from the previous paper here. As pointed out in [16], the reproduction was based on DeepHE3’s open-source project11footnotemark: 1, which is newer than its original publication, and the performance was slightly better than the original one. As we use a fixed random seed ($42$) for all of the experiments, we find that the standard deviation of each metric is very small (all less than $0.007$ meV) across the $10$ repeated experiments. Since the fluctuation is negligible, in the following tables and figures, we just present the mean values of $MAE_{all}$, $MAE^{cha}_{sample}$, and $MAE^{cha}_{block}$ across the repeated experiments, without listing the standard deviations of each entry to save space.
Table 1: Comparison on the $MAE_{all}$, $MAE^{cha}_{sample}$ and $MAE^{cha}_{block}$ metrics (/meV) for the four experimental setups on the databases of monolayer structures.
Methods $MG$ $MM$
$MAE_{all}$ $MAE^{cha}_{sample}$ $MAE^{cha}_{block}$ $MAE_{all}$ $MAE^{cha}_{sample}$ $MAE^{cha}_{block}$
DeepHE3 (Baseline) 0.251 0.357 0.362 0.406 0.574 1.103
DeepHE3+Trace 0.230 0.344 0.348 0.378 0.537 1.091
DeepHE3+Grad 0.185 0.269 0.258 0.308 0.453 0.924
DeepHE3+TraceGrad 0.175 0.257 0.228 0.285 0.412 0.808
Table 2: Comparison on the $MAE_{all}$, $MAE^{cha}_{sample}$ and $MAE^{cha}_{block}$ metrics (/meV) for the four experimental setups on the databases of bilayer structures, where the superscripts $nt$ and $t$ respectively denote the non-twisted and twisted subsets.
Methods $BG^{nt}$ $BG^{t}$
$MAE_{all}$ $MAE^{cha}_{sample}$ $MAE^{cha}_{block}$ $MAE_{all}$ $MAE^{cha}_{sample}$ $MAE^{cha}_{block}$
DeepHE3 (Baseline) 0.389 0.453 0.644 0.264 0.429 0.609
DeepHE3+Trace 0.362 0.417 0.593 0.251 0.401 0.480
DeepHE3+Grad 0.320 0.356 0.511 0.222 0.389 0.446
DeepHE3+TraceGrad 0.291 0.323 0.430 0.198 0.372 0.406
Methods $BB^{nt}$ $BB^{t}$
$MAE_{all}$ $MAE^{cha}_{sample}$ $MAE^{cha}_{block}$ $MAE_{all}$ $MAE^{cha}_{sample}$ $MAE^{cha}_{block}$
DeepHE3 (Baseline) 0.274 0.304 1.042 0.468 0.602 2.399
DeepHE3+Trace 0.259 0.285 0.928 0.429 0.570 1.782
DeepHE3+Grad 0.243 0.272 0.824 0.406 0.542 1.431
DeepHE3+TraceGrad 0.226 0.256 0.740 0.384 0.503 1.284
Methods $BT^{nt}$ $BT^{t}$
$MAE_{all}$ $MAE^{cha}_{sample}$ $MAE^{cha}_{block}$ $MAE_{all}$ $MAE^{cha}_{sample}$ $MAE^{cha}_{block}$
DeepHE3 (Baseline) 0.447 0.480 1.387 0.831 0.850 4.572
DeepHE3+Trace 0.406 0.462 1.239 0.784 0.812 4.520
DeepHE3+Grad 0.342 0.365 0.750 0.742 0.786 4.463
DeepHE3+TraceGrad 0.295 0.312 0.718 0.735 0.755 4.418
Methods $BS^{nt}$ $BS^{t}$
$MAE_{all}$ $MAE^{cha}_{sample}$ $MAE^{cha}_{block}$ $MAE_{all}$ $MAE^{cha}_{sample}$ $MAE^{cha}_{block}$
DeepHE3 (Baseline) 0.397 0.424 0.867 0.370 0.390 0.875
DeepHE3+Trace 0.382 0.397 0.843 0.351 0.367 0.838
DeepHE3+Grad 0.343 0.365 0.696 0.324 0.339 0.746
DeepHE3+TraceGrad 0.300 0.332 0.644 0.291 0.302 0.674
Figure 4: $MAE_{\text{block}}$ on each basic block of the Hamiltonian matrices of monolayer structures.
Combining the results from Table 1, Table 2, Figure 4 and Figure 5, we find that the proposed TraceGrad method significantly enhances the accuracy performance of the baseline method DeepHE3, both on average and for challenging samples or blocks, both on the non-twisted samples and the twisted samples marking strong difficulty for the expressive power and generalization ability of regression models as analyzed in Section 5. Specifically, TraceGrad lowers down the $MAE_{all}$ of DeepHE3 with a relative ratio of $30.28\%$, $29.80\%$, $25.19\%$, $25.00\%$, $17.52\%$, $17.95\%$, $34.00\%$, $11.55\%$, $24.43\%$, and $21.35\%$, on MG, MM, BGnt, BGt, BBnt, BBt, BTnt, BTt, BSnt, and BSt, respectively; and lowering down $MAE_{cha}$ by $28.01\%$, $28.22\%$, $28.70\%$, $13.29\%$, $15.79\%$, $16.45\%$, $35.00\%$, $11.18\%$, $21.70\%$, and $22.56\%$, on the respective datasets; and achieving better results in the vast majority of basic blocks. Particularly, for the blocks where DeepHE3 performs the worst, TraceGrad reduces the $MAE$ ($MAE^{cha}_{block}$) by $37.01\%$, $26.75\%$, $33.23\%$, $33.33\%$, $28.98\%$, $46.48\%$, $48.23\%$, $3.37\%$, $25.72\%$, and $22.97\%$, on the respective datasets, from which the $MAE^{cha}_{block}$ in MM, BGnt and BTnt are lowered down to within $1meV$; the $MAE^{cha}_{block}$ in BBt is decreased by nearly half to approach $1meV$. This indicates that the proposed representation learning framework, while maintaining strict SO(3)-equivariance, significantly enhances the expressiveness of equivariant representations, thereby showcasing immense advantages in capturing the complexity of physical systems and refreshing the SOTA accuracy on the six benchmark databases.
Figure 5: $MAE_{\text{block}}$ on each basic block of the Hamiltonian matrix in non-twisted (marked by superscripts $nt$) and twisted (marked by superscripts $t$) bilayer structures.
From the results of ablation terms, we can obtain a more fine-grained experimental analysis. By comparing among the results of DeepHE3, DeepHE3+Trace, DeepHE3+Grad, and DeepHE3+TraceGrad, we can conclude that the two core mechanisms of our method, i.e., the SO(3)-invariant trace supervision mechanism (Trace) at the label level as well as the gradient-based induction mechanism (Grad) at the representation layer, can contribute to the performance individually. Moreover, their combination provides even better performance. This is because, on one hand, with the gradient-based induction mechanism as a bridge, the SO(3)-invariant features learned from the trace label can be transformed into the SO(3)-equivariant representations during inference; on the other hand, with trace label, the SO(3)-invariant network branch has a strong supervisory signal, enabling it to learn the intrinsic symmetry and complexity of the regression targets, which enhances the quality of SO(3)-invariant features and ultimately benefits the encoding of SO(3)-equivariant features. The value of such complementarity has been fully demonstrated in the experimental results.
The main focus of this work is to bridge rigorous SO(3)-equivariance with the non-linear expressive capabilities of neural networks. While approximately SO(3)-equivariant techniques, which limit the depth of theoretical exploration, are not the focus of this study, they remain of interest in a few numerical computation applications where precision is highlighted more over strict equivariance. Considering this, in the appendix, we explore the synergy of our methods with the Transformer-based approximately SO(3)-equivariant technique [16], and the experimental results also demonstrate a significant accuracy improvement brought by our approach, further supporting the generality of our methodology.
7 Future Work
In future research, various extensions are conceivable across theoretical, methodological, and application fields. Theoretically, we can generalize SO(3) group symmetry to a broader range and more complex groups. This will enable deep learning frameworks to accommodate richer mathematical structures. From a methodological perspective, this study employs the DeepHE3 framework for comparative purposes. However, it is feasible to adopt a more sophisticated network, such as the Transformer-based SO(3)-equivariant non-linear representation learning framework, which utilizes the general TraceGrad methodology. The current methodology extends beyond predicting electronic Hamiltonians and is capable of generally predicting multiple physical quantities and properties that require equivariance. Furthermore, the methodology is not confined to physical research alone and may find applications in areas such as robotics, autonomous vehicles, and motion tracking systems. It handles 3D point cloud data and other vision information to achieve safer and more reliable prediction results with SO(3)-symmetry. Looking forward to future endeavors, the foundational and pivotal role of this work is clear.
8 Method
8.1 Encoding and Decoding Framework
As shown in Figure 1, our encoding framework corresponds to a total of $K$ encoding modules, which are sequentially connected to form a cascading stack. For the $k$-th module ($1\leq k\leq K$), it first introduces an SO(3)-equivariant base encoder, e.g., the encoder from DeepHE3 [7] which is composed of recently developed equivariant operators [19, 20] like linear scaling, direct products, direct sums, Clebsch-Gordan decomposition, gated activation, equivariant normalization, and etc., to encode the physical system’s initial representations, such as spherical harmonics [26], or representations passed from the previous neural layers, as equivariant features $\textbf{f}^{(k)}$ in the current hidden layer. Next, we construct the feature $\textbf{v}^{(k)}=g_{\text{nonlin}}(\textbf{f}^{(k)})$, to achieve sufficient non-linear expressiveness while maintaining SO(3)-equivariance, where $g_{\text{nonlin}}(\cdot)$ is the function defined in Equation (5). As for $s_{nonlin}(\cdot)$, i.e., the built-in non-linear mechanism within $g_{\text{nonlin}}(\cdot)$, although it can be implemented as any differentiable non-linear neural module, here for simplicity, we implement it as a fully connected feed-forward structure containing the SiLU activation function and layer normalization. To facilitate learning and optimization, we connect the input and output of $g_{\text{nonlin}}(\cdot)$ in a residual manner [27], serving as the output of the current encoding module, i.e., $\textbf{o}^{(k)}=\textbf{f}^{(k)}+\textbf{v}^{(k)}$. $\textbf{o}^{(k)}$ is then fed into the next encoder module to yield $\textbf{f}^{(k+1)}$, $\textbf{o}^{(k+1)}$, …., until the last encoder module, which is followed by the decoder modules.
For an expressive feature encoder handling a complex physical system, $\textbf{f}^{(k)}$ is usually not a feature of single degree but a direct-sum concatenation of series of components $\textbf{f}^{(k)l_{1}}$, $\textbf{f}^{(k)l_{2}}$, …, $\textbf{f}^{(k)l_{k}}$ at multiple degrees, i.e., $l_{1}$, $l_{2}$, …, $l_{k}$, where some components share the same degree while others differ. In this case, it becomes necessary to extend the decomposition operator, i.e., $\text{CGDecomp}(\cdot)$ in Equation (5), to accommodate various components of degrees. Moreover, in the context of neural networks, introducing learnable parameters $W$ may improve the model capacity. Based on these considerations, when constructing the encoding module, the $\text{CGDecomp}(\cdot)$ operation in Equation (5) can be expanded as $\text{CGDecomp}_{ext}(\cdot)$:
$u^{(k)}=\text{CGDecomp}_{\text{ext}}(\textbf{f}^{(k)}\otimes\textbf{f}^{(k)},0,W)=\sum_{1\leq i,j\leq k,l_{i}=l_{j}}W_{ij}\cdot\text{CGDecomp}(\textbf{f}^{(k)l_{i}}\otimes\textbf{f}^{(k)l_{j}},0)$ (9)
where $W$ represents learnable parameters. This expansion does not change the equivariance or invariance of the features. It is obvious that, the output ($u^{(k)}$) of $\text{CGDecomp}_{\text{ext}}(\cdot)$ is also SO(3)-invariant, just like the original $\text{CGDecomp}(\cdot)$. Thus, even with such an expansion, the features $z^{(k)}=s_{\text{nonlin}}(u^{(k)})$ remain SO(3)-invariant and $\textbf{v}^{(k)}=\frac{\partial z^{(k)}}{\partial\mathbf{f}^{(k)}}$ remain SO(3)-equivariant.
The outputted SO(3)-equivariant features, i.e., $\textbf{o}^{(K)}$, as well as the outputted SO(3)-invariant features, i.e., $\textbf{z}^{(K)}$, from the final ($K$ th) layer of the encoding modules, are fed into the corresponding decoding modules to yield the predictions of $\mathbf{Q}$ and $\mathbf{T}=tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})$, respectively. Since complex non-linear transformations have already been incorporated in the encoding framework, the decoders can be simple: the SO(3)-equivariant decoder is implemented using an SO(3)-equivariant linear mapping [20], and the SO(3)-invariant decoder is implemented as a shallow feed-forward module.
8.2 Training and Inference
The training loss function is shown as the following:
$\begin{split}\min_{\theta}loss&=loss_{Q}+\mu(loss_{Q},loss_{T})\cdot loss_{T},\\ loss_{Q}&=MSE(\widehat{\mathbf{Q}},\mathbf{Q}^{*}),\\ loss_{T}&=MSE(\widehat{\mathbf{T}},\mathbf{T}^{*})\end{split}$ (10)
where $\theta$ denote all of the learnable parameters of our framework, $\widehat{\mathbf{Q}}$, $\widehat{\mathbf{T}}$ and $\mathbf{Q}^{*}$, $\mathbf{T}^{*}$ respectively denote the predictions and labels of $\mathbf{Q}$ and $\mathbf{T}$. In order that the both of the SO(3)-equivariant and SO(3)-invariant branches can be stably converged, we introduce $\mu(loss_{Q},loss_{T})$, i.e., a function returning a coefficient to regularize the relative scale of the two loss terms:
$\mu(loss_{Q},loss_{T})=\lambda\cdot No\_Grad(\frac{loss_{Q}}{loss_{T}})$ (11)
where $No\_Grad(\cdot)$ denotes gradient discarding when calculating such coefficient, as this coefficient is only used for adjusting the relative scale of the two loss terms and should not itself be a source of training gradients, otherwise it would counteract $loss_{T}$ in Equation (10). $\lambda$ here is a hyperparameter, whose value is determined as $0.3$ based on the performance on the validation set. The introduction of $\mu(loss_{Q},loss_{T})$ effectively prevents the numerical disparity between the two loss terms to stabilize the training for both of them.
In our framework, the SO(3)-invariant non-linear features $z^{(k)}(1\leq k\leq K)$ learned with the help of $\mathbf{T}$ contribute to the learning of SO(3)-equivariant representations $\textbf{v}^{(k)}(1\leq k\leq K)$ and the prediction of $\mathbf{Q}$. To achieve this goal, in the training phase, both the SO(3)-invariant feature encoder and the SO(3)-invariant decoder, used for predicting $\mathbf{T}$ from $z^{(k)}$, need to be trained. While in the inference phase, since $\mathbf{T}$ is not the final regression target, the SO(3)-invariant decoder can be removed, yet the SO(3)-invariant encoder still should be retained, used to induce SO(3)-equivariant features.
9 Acknowledgements
This work was supported by the National Natural Science Foundation of China (Grant Number 12134012) and the Strategic Priority Research Program of Chinese Academy of Sciences (Grant Number XDB0500201).
Appendix A Empirical Study on Combining our Method with Approximately SO(3)-equivariance Framework
Table 3: Comparison on the $MAE_{all}$, $MAE^{cha}_{sample}$ and $MAE^{cha}_{block}$ metrics for approximately SO(3)-equivariant methods on the monolayer MoS2 ($MM$) database. The results of the compared methods are copied from their papers, where the empty items are due to the data not being provided in the original text.
Methods $MM$
$MAE_{all}$ $MAE^{cha}_{sample}$ $MAE^{cha}_{block}$
DeepH-2 [23] 0.21 - -
HarmoSE [16] 0.233 0.293 0.406
HarmoSE+TraceGrad 0.178 0.228 0.296
Figure 6: Comparison on the $MAE_{block}$ metric for approximately SO(3)-equivariant methods on the monolayer MoS2 ($MM$) database.
In this part, taking the MoS2 dataset as a case study, we evaluate the performance of combining our trace supervision and gradient induction method (TraceGrad) with the approximately equivariant approach HarmoSE [16]. We here take HarmoSE as the backbone encoder, and yields features by TraceGrad to enrich its representations. The experimental results in Table 3 and Figure 6 demonstrate that TraceGrad significantly enhances the accuracy of HarmoSE, surpassing DeepH-2 and achieving new SOTA results. Both DeepH-2 and HarmoSE sacrificed strict SO(3)-equivariance to fully release the expressive capabilities of graph Transformers, aiming for the ultimate in prediction accuracy. Despite this, TraceGrad still manages to significantly exceed their accuracy, further confirming the superiority and generality of our method in learning expressive representations of physical systems.

### v3

@@
 Abstract
-We present both a theoretical and a methodological framework that addresses a critical challenge in applying deep learning to physical systems: the reconciliation of non-linear expressiveness with SO(3)-equivariance in predictions of SO(3)-equivariant quantities, such as the electronic-structure Hamiltonian.
+We present both a theoretical and a methodological framework that addresses a critical challenge in applying deep learning to physical systems: the reconciliation of non-linear expressiveness with SO(3)-equivariance in predictions of SO(3)-equivariant quantities.
 Inspired by covariant theory in physics, we address this problem by exploring the mathematical relationships between SO(3)-invariant and SO(3)-equivariant quantities and their representations.
@@
 This mechanism can incorporate non-linear expressive capabilities into SO(3)-equivariant representations, while theoretically preserving their equivariant properties as we prove.
-Our approach offers a promising general solution to the critical dilemma between equivariance and non-linear expressiveness in deep learning methodologies.
-We apply our theory and method to the electronic-structure Hamiltonian prediction tasks, demonstrating state-of-the-art performance across six benchmark databases.
+We apply our theory and method to the electronic-structure Hamiltonian prediction tasks, experimental results on eight benchmark databases covering multiple types of elements and challenging scenarios show dramatic breakthroughs on the state-of-the-art prediction accuracy, with improvements of up to 40% in predicting Hamiltonians and up to 76% in predicting downstream physical quantities such as occupied orbital energy.
+Our approach goes beyond handling physical systems and offers a promising general solution to the critical dilemma between equivariance and non-linear expressiveness for the deep learning paradigm.
 keywords:
-Deep Learning Theory, Neural Networks, SO(3)-equivariance, Science for AI, AI for Science
+Deep Learning, Neural Networks, SO(3)-equivariance, Atomic Modeling, Hamiltonian Prediction
 1 Introduction
 With the advantages in computational complexity and generalization capabilities, deep learning paradigm has vigorously driven advancements in physics research [1].
-For example, in predicting electronic-structure Hamiltonians, traditional Density Functional Theory (DFT) methods [2, 3] suffer from computational complexities of $\mathcal{O}(N^{3})$.
-In contrast, deep learning approaches have significantly reduced this complexity to $\mathcal{O}(N)$ [4, 5, 6, 7], accelerating processes by thousands or even millions of times.
+For example, in predicting electronic-structure Hamiltonians, traditional Density Functional Theory (DFT) methods [2, 3] suffer from computational complexities of $\mathcal{O}(N^{3})$, where $N$ is the number of atoms in a system.
+In contrast, deep learning approaches have significantly reduced this complexity to $\mathcal{O}(N)$ [4, 5, 6, 7, 8], accelerating processes by thousands or even millions of times.
 This reduction in computational demand not only unlocks possibilities for analyzing extremely large atomic systems but also paves the way for efficient material simulation, discovery, and design that were previously unimaginable.
@@
 However, this approach does not provide the strict assurance of SO(3)-equivariance required for physical research.
-Besides, the calculation of physical quantities calls for ultra-high numerical accuracy.
-For example, it is typically required that the errors on Hamiltonian calculation being restricted within the sub-millielectron volts (sub-meV) scale, to ensure accurate derivation of downstream physical quantities and related material properties.
-This necessitates that the neural networks possess strong capabilities to express the complex non-linear mappings from atomic structures to the regression targets.
+Besides, the calculation of physical quantities calls for high numerical accuracy, necessitating that the neural networks possess strong capabilities to express the complex non-linear mappings from atomic structures to the regression targets.
 However, existing deep learning methods are unable to simultaneously ensure strict SO(3)-equivariance and high numerical accuracy.
-The root cause of this problem lies in the conflict between SO(3)-equivariance and non-linear expressiveness: specifically, directly applying non-linear activation functions on SO(3)-equivariant features (with degree $l\geq 1$) leads to the loss of equivariance, while bypassing non-linear mappings severely restricts the network’s expressive capabilities and thereby lowering down the achievable accuracy.
-Although several efforts have been made to address this challenge—introducing gated activation mechanisms [7, 10, 11, 12], employing discrete spherical harmonic bases [13, 14, 15], and developing hybrid regression approaches [16]—they still fall short, to varying degrees, of fully reconciling strict adherence to SO(3)-equivariance with the requirements for high-precision generalization capability.
-As analyzed in previous literature [13], this issue is commonly found in machine learning tasks that demand both strict equivariance and fine-grained generalization performance.
-To address this challenge, we make theoretical and methodological explorations on unifying strict SO(3)-equivariance with strong non-linear expressiveness within the realm of deep representation learning.
+The root cause of this problem lies in the conflict between SO(3)-equivariance and non-linear expressiveness: specifically, directly applying non-linear activation functions on SO(3)-equivariant features (with degree $l\geq 1$) may lead to the loss of equivariance, while bypassing non-linear mappings severely restricts the network’s expressive capabilities and thereby lowering down the achievable accuracy.
+As pointed out in previous literature [11], this issue is commonly found in machine learning tasks that demand both strict equivariance and fine-grained generalization performance.
+Although several efforts have been made to alleviate this issue by introducing gated activation mechanisms [8, 7, 12, 13, 14], developing local coordinate strategy [6], and employing discrete spherical channels [11, 15, 16], they still, to varying degrees, compromised either non-linear expressiveness [7, 8, 12, 13, 14] or strict SO(3)-equivariance [11, 15, 16] when processing physical systems.
+Please refer to Supplementary Information A for an in-depth analysis.
+To address the equivariance-expressiveness dilemma, we make theoretical and methodological explorations on unifying strict SO(3)-equivariance with strong non-linear expressiveness within the realm of deep representation learning.
 We are inspired by the insight that invariant quantities in transformation often reflect the mathematical nature of physical laws and can induce other quantities with equivariant properties.
@@
 This invariance is foundational in formulating physical laws that involve covariant quantities.
-Furthermore, from the perspective of deep representation learning, the attribute of preserving invariance under non-linear operations is a significant advantage, due to its compatibility with non-linear expressive capabilities.
+Furthermore, from the perspective of deep representation learning, the attribute that invariance is preserved under non-linear operations is a significant advantage, given its compatibility with non-linear expressive capabilities.
 Built upon these insights, we theoretically solve the equivariance-expressiveness dilemma by intensively exploring and making use of the intrinsic relationships between SO(3)-invariant and SO(3)-equivariant quantities and representations: we first dedicate efforts to learning high-quality SO(3)-invariant features with ample non-linear expressiveness, and subsequently, we derive SO(3)-equivariant non-linear representations from these SO(3)-invariant ones.
@@
 A significant advantage of these SO(3)-invariant quantities lies in the fact that they are directly derived from the SO(3)-equivariant target labels and can serve as unique supervision labels for the effective learning of informative SO(3)-invariant features that capture the intrinsic symmetry properties of the mathematical structure of $\mathbf{Q}$ without requiring additional labeling resources.
-Second, we propose a gradient-based mechanism to induce SO(3)-equivariant representations as well as the regression target $\mathbf{Q}$, from high-quality SO(3)-invariant features learned under the supervision of $tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})$.
+Second, we propose a gradient-based mechanism to induce SO(3)-equivariant representations for predicting the regression target $\mathbf{Q}$, from high-quality non-linear SO(3)-invariant features learned under the supervision of $tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})$.
 This mechanism can incorporate non-linear expressive capabilities into SO(3)-equivariant representations while preserving their equivariant properties, as we prove.
-We develop our theory into a unified SO(3)-equivariant non-linear representation learning method, and apply it to the computation of electronic structure Hamiltonians, which are central to quantum physics and inherently complex, presenting substantial challenges for regression models.
-Our method significantly enhances the state-of-the-art performance on six benchmark Hamiltonian databases, demonstrating that our approach not only maintains strict SO(3)-equivariance but also unleashes powerful non-linear expressiveness, thereby boosting the prediction accuracy of Hamiltonians, providing an effective deep learning tool for efficient and accurate electronic structure calculations of materials.
-Our theory and method for bridging symmetry and expressiveness of deep learning are general and can be easily applied on predicting other physical quantities such as forces, which also obey SO(3)-equivariance.
-Furthermore, it can also be extended to other application areas beyond physics.
-For instance, in some computer vision tasks, such as autonomous driving [18], where there is a high demand for robustness against coordinate transformations of cameras, the mainstream approach has been to achieve approximate SO(3)-equivariance through data augmentation.
-However, this does not guarantee absolute safety.
-Our work holds great potential to construct deep models with high generalization performance under the premise of strict SO(3)-equivariance, potentially revolutionizing techniques in such tasks.
-2 The Equivariance-Expressiveness Dilemma in Existing Deep Learning Methods
-In this section, we introduce existing efforts to harmonize the deep learning paradigm with SO(3)-equivariance, analyzing the dilemma between strict SO(3)-equivariance and non-linear expressive capabilities.
-2.1 Approaches with Strict SO(3)-equivariance but Weak Expressiveness
-Most existing methods have focused on achieving SO(3)-equivariance at the neural mechanism level by applying group theory-guaranteed SO(3)-equivariant operators [19, 20].
-These operators, including linear scaling, element-wise sum, direct products, direct sums, and Clebsch-Gordan decomposition, have been used to build graph neural network architectures for tasks in 3D point cloud analysis [21], molecular dynamic simulation and property prediction [11, 10], as well as Hamiltonian prediction [22, 4, 7, 12].
-However, these operators could not provide sufficient expressiveness to fit complex non-linear mappings.
-The majority of these operators are linear operations, and the only one (direct product) that introduces some non-linearity can only approximate a simple squaring function, and cannot be extended to more complex non-linear forms due to the huge computational burden and difficulty in convergence.
-Even worse, directly applying highly-expressive non-linear activation functions, such as $Sigmoid$, $Softmax$, and $SiLU$, to SO(3)-equivariant features may result in the loss of equivariance.
-This dilemma leads to an unsatisfactory trade-off between strict equivariance and non-linear expressiveness, ultimately reducing the achievable performance of these regression models.
-To alleviate this issue, these methods introduced a gated activation mechanism that processes SO(3)-invariant features ($l=0$) using non-linear activations and employs them as linear gating coefficients for SO(3)-equivariant features ($l\geq 1$), aiming to enhance their expressive power while maintaining strict equivariance.
-However, from the perspective of SO(3)-equivariant features, this mechanism is a combination of linear operations and cannot fundamentally increase their non-linear expressive capabilities.
-2.2 Approaches with Remarkable Expressiveness but Approximate SO(3)-equivariance
-Li et al.
-[6] proposed a local coordinate strategy to achieve SO(3)-equivariance.
-This strategy relied on the transformation relationship between local and global coordinates to achieve equivariance, without imposing restrictions on the neural network, allowing the use of network modules that possess non-linear expressiveness.
-However, this strategy, solely depending on coordinate transformations, only works under holistic rigid rotations of the atomic system.
-Without exploration of a neural mechanism for equivariance, when the local coordinates experience non-rigid perturbations such as thermal movements or bilayer twists, this strategy offers no guarantees on maintaining symmetry properties; Zitnick et al.
-[13] decomposed features into weighted sums of spherical harmonic basis functions, each corresponding to a discrete point on a unit sphere.
-Such discrete decomposition enables non-linear operations on the coefficients to enhance expressiveness.
-In subsequent developments [14, 15, 23], this approach has demonstrated a remarkable capacity to fit complex functions.
-However, it degenerates from continuous to discrete rotational equivariance, no longer maintaining strict equivariance to continuous rotational operations; Yin et al.
-[16] proposed a hybrid framework consisting of both group theory-guaranteed SO(3)-equivariant mechanisms and non-linear mechanisms to regress Hamiltonians.
-In this framework, the non-linear mechanisms showed remarkable capability at learning SO(3)-equivariance from the data with the help of the theoretically SO(3)-equivariant mechanisms, and released powerful non-linear expressive capabilities to achieve more numerical accuracy.
-However, the equivariance achieved through data-driven methods does not have strict theoretical guarantees.
-In summary, although these methods with approximate SO(3)-equivariance achieve high numerical accuracy due to their extensive use of non-linear mechanisms, they are best suited for the applications that primarily require numerical approximations but not for tasks that demand theoretical depth in physics.
-In some applications, the demands for symmetry are extremely high.
-Even minor deviations from perfect symmetry can result in incorrect physical results.
-Combining the analysis in this section, existing deep learning methods have compromised either strict SO(3)-equivariance or non-linear expressiveness when processing physical systems.
-In this paper, we develop a novel framework that theoretically combines strict SO(3)-equivariance with the non-linear expressiveness of neural networks.
-This solution resolves the equivariance-expressiveness dilemma by deeply exploring the intrinsic relationships between SO(3)-invariant and SO(3)-equivariant quantities.
+We develop our theory into a unified SO(3)-equivariant non-linear representation learning method, and apply it to the computation of electronic-structure Hamiltonian, which occupies a central position in quantum mechanics to induce key physical quantities such as electronic wavefuction, orbital energy, and band structure, bringing deeply understanding of the electrical, optical, magnetic, and transport properties of atomic systems, while posing significant difficulties for machine learning techniques due to its intrinsic symmetry with respect to the SO(3) group as well as the high-dimensional complexity, as shown in Fig.
+1.
+Our method dramatically enhances the state-of-the-art performance on eight benchmark Hamiltonian databases from the renowned DeepH and QH9 benchmark series [6, 8, 19], showing excellent generality to various challenging scenarios such as thermal motions, bilayer twists, scale variations, and new trajectories, with an accuracy improvement of Hamiltonians of up to 40%.
+Moreover, our approach substantially enhances the prediction accuracy of downstream physical quantities of Hamiltonians including occupied orbital energies and electronic wavefunctions, with an accuracy improvement by a maximum of 76% and 17%, respectively.
+Our leading performance comprehensively demonstrate that our method, while satisfying SO(3)-equivariance, possesses excellent expressive power and generalization performance, providing an effective deep learning tool for efficient and accurate electronic-structure calculations of atomic systems.
+Figure 1: Illustration of the high-dimensional complexity of the Hamiltonian.
+In an atomic system, a Hamiltonian matrix is defined between each pair of atoms within each other’s cutoff radius, constituting a large cluster of Hamiltonian matrices for all atom pairs.
+Besides, each Hamiltonian matrix contains multiple basic blocks, with the $\mathbf{Q}=\mathbf{Q}^{l_{p}\otimes l_{q}}$ defined in Section 2 corresponding to a basic block $\mathbf{H}^{l_{p}\otimes l_{q}}$ here.
+Thus, for a Hamiltonian prediction model, the space of the regression target is typically huge.
 2 Problem Formalization
@@
 ${\bf{f}}({\bf{R}})^{(k+1)l}={\bf{D}}^{l}({\bf{R}})\cdot{\bf{f}}^{(k+1)l},{\bf{f}}^{(k+1)l}=g_{nonlin}({\bf{f}}^{(k)l})$ (4)
-However, directly implementing $g_{nonlin}(\cdot)$ as neural network module with non-linear mechanisms, such as $Sigmoid$, $Softmax$ and $Silu$, may result in the destruction of strict equivariance.
+However, directly implementing $g_{nonlin}(\cdot)$ as neural network module with non-linear activation functions, such as $Sigmoid$, $Softmax$ and $Silu$, may result in the destruction of strict equivariance.
 How to make $g_{nonlin}(\cdot)$ both theoretically SO(3)-equivariant and capable of non-linear expressiveness, is the core problem this paper aims at solving.
@@
 First, we propose a theoretical construct to derive SO(3)-invariant quantities from the SO(3)-equivariant targets, aiming at facilitating the learning of high-quality SO(3)-invariant features.
-Second, we propose a gradient-based mechanism encoding features maintaining theoretically SO(3)-equivariance while possessing the non-linear expressive capability induced from SO(3)-invariant features to capture the complex non-linear patterns inherent in physical quantities.
+Second, we propose a gradient-based mechanism encoding features maintaining theoretically SO(3)-equivariance while possessing the non-linear expressive capability induced from SO(3)-invariant non-linear features to capture the complex non-linear patterns inherent in physical quantities.
 Theorem 1.
 The quantity $\mathbf{T}=tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})$ is SO(3)-invariant, where $\mathbf{Q}$ is the simplified representation (without superscripts) of $\mathbf{Q}^{l_{p}\otimes l_{q}}$ defined in Section 2, and $\dagger$ denotes the conjugate transpose operation, $tr(\cdot)$ is the trace operation.
+Theorem 2.
+The non-linear neural mapping defined as the following is SO(3)-equivariant, satisfying the requirements specified in Eq. (4):
+$\mathbf{v}=g_{nonlin}(\mathbf{f})=\frac{\partial z}{\partial\mathbf{f}}\quad\text{s.t.}\quad z=s_{nonlin}(u),u=\text{CGDecomp}(\mathbf{f}\otimes\mathbf{f},0)$ (5)
+where $\mathbf{f}$ is an input SO(3)-equivariant feature with degree $l$ in the direct-sum state, $\otimes$ denotes the direct-product operation of tensors, $\text{CGDecomp}(\cdot,0)$ refers to performing a Clebsch-Gordan decomposition of the tensor and returning the scalar component of degree 0, $s_{nonlin}(\cdot)$ represents non-linear neural modules, and $\mathbf{v}$ is the outputted feature encoded by $g_{nonlin}(\cdot)$.
+Proofs of the two theorems are included in Supplementary Information B.
+Remark 1.
+In Theorem 1, the term $tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})$ is equal to the square of the Frobenius norm of the matrix $\mathbf{Q}$, i.e., $tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})=\|\mathbf{Q}\|_{F}^{2}=\sum_{m,n}|\mathbf{Q}_{mn}|^{2}$.
+The invariance of Frobenius norm under unitary transformations reflects the profound connections between modulus and invariance in physics.
+Remark 2.
+In Theorem 2, $g_{\text{nonlin}}(\cdot)$ can handle equivariant features of any degree and the returned feature $\mathbf{v}$ has the same degree $l$ as the input feature $\mathbf{f}$.
+$\mathbf{v}$ is endowed with non-linear expressiveness with the effect of $s_{nonlin}(\cdot)$, which can be any differentiable non-linear network module, including non-linear activation function as a part.
+The gradient operation $\frac{\partial z}{\partial\mathbf{f}}$ in $g_{\text{nonlin}}(\cdot)$ is used during the forward inference process to promote feature encoding, expanding the usage and significance of gradients within the deep learning paradigm.
+Taking $z$ in Eq.
+(5) as the bridge connecting Theorem 1 and Theorem 2, we present a general method for learning non-linear representations that are equivariant to SO(3), as shown in Fig.
+2.
+The core of our method can be abstractly referred to as TraceGrad.
+At the label level, it incorporates the SO(3)-invariant trace quantity $tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})$ introduced in Theorem 1 as a supervisory signal for learning SO(3)-invariant non-linear feature $z$.
+Given the attribute that invariance is preserved under non-linear operations, $z$ is encoded by $s_{nonlin}(\cdot)$ to obtain non-linear expressive capabilities.
+At the representation level, by taking the gradient of $z$ with respect to $\mathbf{f}$, the non-linear expressiveness of $z$ is transferred to the equivariant feature $\mathbf{v}$, while maintaining strict SO(3)-equivariance as we prove.
+Subsequently, merging $\mathbf{v}$ and $\mathbf{f}$, and applying $g_{\text{nonlin}}(\cdot)$ in a stacked manner can yield rich SO(3)-equivariant non-linear representations to predict the SO(3)-equivariant regression target.
+For details of the method, please refer to Section 6.
+Figure 2: Methodological framework for learning SO(3)-equivariant representations with non-linear expressiveness to regress complex SO(3)-equivariant targets.
+4 Results and Analyses
+We apply our theory and method to the electronic-structure Hamiltonian prediction task, and collect results on eight benchmark databases, i.e., Monolayer Graphene ($MG$), Monolayer MoS2 ($MM$), Bilayer Graphene ($BG$), Bilayer Bismuthene ($BB$), Bilayer Bi2Te3 ($BT$), Bilayer Bi2Se3 ($BS$), QH9-Stable (QS), and QH9-Dynamic (QD).
+The first six databases, consisting of crystal systems with elements like C, Mo, S, Bi, Te and Se, are from the DeepH benchmark series [6, 8]; while the last two databases, composed of molecular systems of C, H, O, N, F, are from the QH9 benchmark series [19] .
+These databases present diverse and complex challenges to the generality of a regression model.
+Regarding $MG$, $MM$, and QD, as their samples are prepared from a 300K environment, the thermal motions lead to complex non-rigid deformations, increasing the difficulty of Hamiltonian prediction.
+For $BG$, $BB$, $BT$, and $BS$, their twisted samples with an interplay of SO(3)-equivariant effects and van der Waals (vdW) force variations bring significant generalization challenges, which are further exacerbated by the absence of any twisted samples in the training sets.
+Besides, $BB$, $BT$, and $BS$ exhibit strong spin-orbit coupling (SOC) effects, which further increase the complexity of Hamiltonian modeling.
+For QS and QD, as we select the most difficult subset split strategies provided in their paper, namely ’ood’ and ’mol’, respectively, yielding differences in atom numbers and trajectories between training and testing samples, challenges in scale and trajectory extrapolation become prominent.
+Detailed information of these databases, including their statistics and visualizations, can be found in the Supplementary Information C.
+We use a comprehensive set of metrics to deeply evaluate the performance of electronic-structure Hamiltonian prediction models.
+On the databases from the DeepH benchmark series [6, 8], we adpot the metrics of Mean Absolute Error (MAE) between predicted and ground truth Hamiltonians introduced by Yin et al.
+[17], including $MAE^{H}_{all}$ for measuring average MAE of all samples and matrix elements, $MAE^{H}_{cha\_s}$ for measuring the MAE of challenging samples where the baseline model performs the worst, $MAE^{H}_{block}$ for measuring the MAE of different basic blocks in the Hamiltonian matrix, and $MAE^{H}_{cha\_b}$ for measuring the MAE on the most challenging Hamiltonian block where the baseline model shows the poorest performance (with the largest $MAE^{H}_{block}$).
+These metrics comprehensively reflect the accuracy performance, covering not only the average accuracy but also the accuracy on difficult samples and challenging blocks of the Hamiltonian matrices.
+On the two databases from the QH9 benchmark series, we adopt the metrics introduced by their original paper [19], including MAE of Hamiltonian matrices, which are further subdivided into $MAE^{H}_{all}$ for measuring average MAE, $MAE^{H}_{diag}$ for measuring MAE of Hamiltonian matrix formed by an atom with itself, and $MAE^{H}_{non\_diag}$ for measuring MAE of Hamiltonian matrices formed by different atoms; as well as the MAE ($MAE^{\epsilon}$) of occupied orbital energies $\epsilon$ induced by the predicted Hamiltonians and compared to the ground truth ones, and the cosine similarity ($Sim(\psi)$) between the electronic wavefunctions $\psi$ induced by the predicted and ground truth Hamiltonians.
+$\epsilon$ and $\psi$ are crucial downstream physical quantities for determining multiple properties of the atomic systems as well as their dynamics, highly reflecting the application values of the Hamiltonian regression model.
+We compare experimental results from two setups: the first one is the experimental results of the baseline SO(3)-equivariant regression model [8, 7] for Hamiltonian prediction, and the second one is the experimental results of extending the architecture and pipeline of the baseline model through the proposed TraceGrad method, which incorporates non-linear expressiveness into the SO(3)-equivariant features of the baseline model with the gradient operations of SO(3)-invariant non-linear features learned by the trace targets.
+We choose DeepH-E3 [8] as the baseline model for databases from the DeepH benchmark series [6, 8]; and we choose QHNet [7] as the baseline model for databases from the QH9 benchmark series [19].
+The results of DeepH-E3 to be compared are taken from the reproduction of Yin et al.
+[17].
+The reproduction, based on DeepH-E3’s newest open-source project [^1] https://github.com/Xiaoxun-Gong/DeepH-E3, showed slightly better results than that in the original publication, and was evaluated more comprehensively across multiple metrics.
+The results of QHNet to be compared are taken from their original paper [19], and for the unification of MAE units, we convert the units of MAE from $10^{-6}$ Hartree ($E_{h}$) in the original paper to meV222 $27211.41E_{h}=$ 27211.4 meV.
+The results of our methods are the average from 10 independent repeated experiments.
+To ensure reproducibility, we used a fixed random seed for all random processes.
+As a result, the standard deviation of the Hamiltonian prediction MAE across repeated experiments does not exceed $0.007$ meV, which is negligible.
+We list the results of DeepH-E3 and DeepH-E3+TraceGrad in Tables 1 and 2 for monolayer structure databases ($MG$, $MM$) and bilayer structure databases ($BG$, $BB$, $BT$, $BS$) from the DeepH benchmark series, reporting the values of $MAE^{H}_{all}$, $MAE^{H}_{cha\_s}$, and $MAE^{H}_{cha\_b}$.
+Regarding the metric of $MAE^{H}_{block}$ for every Hamiltonian block, due to its large data volume, we just present its values for representative datasets here in Fig.
+3, and attach its values for all databases from the DeepH series in Supplementary Information E.
+From these results, we could find that the proposed TraceGrad method dramatically enhances the accuracy performance of the baseline method DeepH-E3, both on average and for challenging samples or blocks, both on the non-twisted samples and the twisted samples.
+Specifically, on the corresponding datasets, TraceGrad lowers down the $MAE^{H}_{all}$ and $MAE^{H}_{cha}$ of DeepH-E3 with relative ratios of up to $34\%$ and $35\%$, respectively, and achieving better results in the vast majority of basic blocks.
+Particularly, for the blocks where DeepH-E3 perform the worst, TraceGrad reduces the $MAE$ ($MAE^{H}_{cha\_b}$) by a maximum of $48\%$.
+The leading performance on the $MG$ and $MM$ databases prepared at 300K demonstrates the robustness of our method against thermal motion.
+The high accuracy on the $BB$, $BT$, and $BS$ databases, which have strong SOC effects, indicates our method’s strong capability to model such effects.
+The excellent performance on the $BG^{t}$, $BB^{t}$, $BT^{t}$, and $BS^{t}$ subsets showcases the method’s superior generalization to twisted structures, which are not present in the training data.
+The outstanding performance on such samples highlights the good potential for studying twist-related phenomena, given interlayer twists are hot research topics because they may bring new electrical and transport properties [20, 21, 22].
+Additionally, the $BG^{t}$ and $BB^{t}$ testing subsets contain significantly larger atomic systems compared to the training set (see Supplementary Information C for statistics of their sizes), yet our method still excels on these subsets for all metrics, demonstrating its good scalability on the sizes of atomic systems it handles.
+We list the results of QHNet and QHNet+TraceGrad in Tables 3 and 4, for the $QS$ and $QD$ databases, respectively, reporting the values of $MAE^{H}_{all}$, $MAE^{H}_{diag}$, $MAE^{H}_{non\_diag}$, $MAE^{\epsilon}$, and $Sim(\psi)$.
+From these results, we can observe that the proposed TraceGrad method significantly enhances the accuracy performance of the baseline method QHNet across all of these metrics on both of the two databases.
+Specifically, TraceGrad lowers down the $MAE^{H}_{all}$, $MAE^{H}_{non\_diag}$, $MAE^{H}_{diag}$, $MAE^{\epsilon}$ and $Sim(\psi)$ of QHNet with relative ratios of up to 40%, 39%, 40%, 76%, and 17%, on the corresponding databases.
+The significant improvements on the $QS$ database, partitioned using the ’ood’ split strategy [19] without scales overlapping among the training, validation, and testing sets, once again demonstrate the method’s strong generalization capabilities across different scales.
+Meanwhile, the improvements on the $QD$ database under the ’mol’ strategy [19] partitioning the training, validation, and testing sets with samples from totally different thermal motion trajectories, highlight the method’s robustness in generalizing to new thermal motion sequences.
+Furthermore, the significant promotion on the prediction accuracy of $\epsilon$, i.e., occupied orbital energies which are crucial for determining electronic properties such as optical characteristics and conductivity in atomic systems, and $\psi$, i.e., the electronic wavefunctions essential for understanding electron distribution and interactions, underscore good application potential of our method for tasks like material design and quantum computing.
+The experimental results above indicate that the proposed representation learning framework, while maintaining strict SO(3)-equivariance, significantly enhances the expressiveness of equivariant representations, thus excellently capturing the complexity of physical systems and substantially breakthroughs the SOTA accuracy on the eight benchmark databases.
+The consistent improvements brought by our framework to different baseline models (DeepH-E3 or QHNet) and in various challenging scenarios, such as thermal motions, bilayer twists, scale variations, and new trajectories, fully demonstrate the universality and significance of our proposed theory and method.
+Table 1: Experimental results measured by the $MAE^{H}_{all}$, $MAE^{H}_{cha\_s}$, and $MAE^{H}_{cha\_b}$ metrics (meV) on the Monolayer Graphene ($MG$) and Monolayer MoS2 ($MM$) databases.
+$\downarrow$ means lower values of the metrics correspond to better accuracy.
+Methods $MG$ $MM$
+$MAE$ ()($\downarrow$)( ↓ )
+$MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$ $MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$
+DeepH-E3 (Baseline) 0.251 0.357 0.362 0.406 0.574 1.103
+DeepH-E3+TraceGrad 0.175 0.257 0.228 0.285 0.412 0.808
+Table 2: Experimental results measured by the $MAE^{H}_{all}$, $MAE^{H}_{cha\_s}$ and $MAE^{H}_{cha\_b}$ metrics (meV) on the Bilayer Graphene ($BG$), Bilayer Bismuthene ($BB$), Bilayer Bi2Te3 ($BT$) and Bilayer Bi2Se3 ($BS$) databases, where the superscripts $nt$ and $t$ respectively denote the non-twisted and twisted subsets.
+$\downarrow$ means lower values of the metrics correspond to better accuracy.
+Methods $BG^{nt}$ $BG^{t}$
+$MAE$ ()($\downarrow$)( ↓ )
+$MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$ $MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$
+DeepH-E3 (Baseline) 0.389 0.453 0.644 0.264 0.429 0.609
+DeepH-E3+TraceGrad 0.291 0.323 0.430 0.198 0.372 0.406
+Methods $BB^{nt}$ $BB^{t}$
+$MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$ $MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$
+DeepH-E3 (Baseline) 0.274 0.304 1.042 0.468 0.602 2.399
+DeepH-E3+TraceGrad 0.226 0.256 0.740 0.384 0.503 1.284
+Methods $BT^{nt}$ $BT^{t}$
+$MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$ $MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$
+DeepH-E3 (Baseline) 0.447 0.480 1.387 0.831 0.850 4.572
+DeepH-E3+TraceGrad 0.295 0.312 0.718 0.735 0.755 4.418
+Methods $BS^{nt}$ $BS^{t}$
+$MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$ $MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$
+DeepH-E3 (Baseline) 0.397 0.424 0.867 0.370 0.390 0.875
+DeepH-E3+TraceGrad 0.300 0.332 0.644 0.291 0.302 0.674
+Figure 3: Visualization of $MAE^{H}_{block}$ on each basic block of the Hamiltonian matrices for representative datasets, where the superscripts $nt$ and $t$ respectively denote the non-twisted and twisted subsets.
+Lower value of $MAE^{H}_{block}$ corresponds to better accuracy.
+The $MAE^{H}_{block}$ for other datasets from DeepH series can be found in Supplementary Information E.
+Table 3: Experimental results measured by the $MAE^{H}_{all}$, $MAE^{H}_{diag}$, $MAE^{H}_{non\_diag}$, $MAE^{\epsilon}$, and $Sim(\psi)$ metrics on the QH9-Stable (QS) database using the ’ood’ split strategy [19].
+$\downarrow$ means lower values of the metrics correspond to better accuracy, while $\uparrow$ means higher value of the metric corresponds to better performance.
+The units of MAE metrics are meV, while $Sim(\psi)$ is the cosine similarity which is dimensionless.
+Methods MAE ()($\downarrow$)( ↓ ) $Sim(\psi)$ ()($\uparrow$)( ↑ )
+$MAE^{H}_{all}$ $MAE^{H}_{diag}$ $MAE^{H}_{non\_diag}$ $MAE^{\epsilon}$
+QHNet 1.962 3.040 1.902 17.528 0.937
+QHNet+TraceGrad 1.191 2.125 1.139 8.579 0.948
+Table 4: Experimental results measured by the $MAE^{H}_{all}$, $MAE^{H}_{diag}$, $MAE^{H}_{non\_diag}$, $MAE^{\epsilon}$, and $Sim(\psi)$ metrics on the QH9-Dynamic (QD) database using the ’mol’ split strategy [19].
+$\downarrow$ means lower values of the metrics correspond to better accuracy, while $\uparrow$ means higher value of the metric corresponds to better performance.
+The units of MAE metrics are meV, while $Sim(\psi)$ is the cosine similarity which is dimensionless.
+Methods MAE ()($\downarrow$)( ↓ ) $Sim(\psi)$ ()($\uparrow$)( ↑ )
+$MAE^{H}_{all}$ $MAE^{H}_{diag}$ $MAE^{H}_{non\_diag}$ $MAE^{\epsilon}$
+QHNet 4.733 11.347 4.182 264.483 0.792
+QHNet+TraceGrad 2.819 6.844 2.497 63.375 0.927
+We present experimental results of ablation study in Supplementary Information F.
+These fine-grained experiments assess the individual and combined contributions of the two core mechanisms of the proposed TraceGrad method on the six databases from the DeepH series.
+These mechanisms are the SO(3)-invariant trace supervision mechanism (Trace) at the label level and the gradient-based induction mechanism (Grad) at the representation layer.
+Results indicate that each mechanism can contribute individually to the performance.
+Moreover, their combination provides even better performance.
+This is because, on one hand, with the gradient-based induction mechanism as a bridge, the SO(3)-invariant features learned from the trace label can be transformed into the SO(3)-equivariant representations during inference; on the other hand, with trace label, the SO(3)-invariant network branch has a strong supervisory signal, enabling it to learn the intrinsic symmetry and complexity of the regression targets, enhancing the quality of SO(3)-invariant features and ultimately benefits the encoding of SO(3)-equivariant features.
+The value of such complementarity has been fully demonstrated in the experimental results.
+5 Future Work
+In future research, various extensions are conceivable across theoretical, methodological, and application fields:
+First, from a theoretical perspective, we can generalize SO(3) group symmetry to a broader range and more complex groups.
+This will enable deep learning frameworks to accommodate richer mathematical structures.
+Second, from a methodological perspective, we plan to combine our framework with advanced Transformer [23] and KAN [24] architectures to continuously enhance the expressive capabilities of equivariant features.
+Third, from the application perspective, the current methodology extends beyond predicting electronic-structure Hamiltonians and is capable of generally predicting multiple physical quantities and properties that require equivariance.
+Furthermore, the methodology is not confined to physical research alone and may find applications in areas such as robotics, autonomous vehicles, and motion tracking systems.
+For instance, in the vision tasks of autonomous vehicles [25], where there is a high demand for robustness against coordinate transformations of cameras, the mainstream approach has been to achieve approximate SO(3)-equivariance through data augmentation.
+However, this does not guarantee absolute reliability and safety.
+Our work holds great potential to construct deep models with high generalization performance under the premise of strict SO(3)-equivariance, potentially revolutionizing techniques in such tasks.
+Looking forward to future endeavors, the foundational and pivotal role of this work is clear.
+6 Method
+6.1 Encoding and Decoding Framework
+As shown in Fig.
+2, our encoding framework corresponds to a total of $K$ encoding modules, which are sequentially connected to form a cascading stack.
+For the $k$-th module ($1\leq k\leq K$), it first introduces an SO(3)-equivariant base encoder, e.g., the encoder from DeepH-E3 [8] or QHNet [7] which is composed of recently developed equivariant operators [26, 27] like linear scaling, direct products, direct sums, Clebsch-Gordan decomposition, gated activation, equivariant normalization, and etc., to encode the physical system’s initial representations, such as spherical harmonics [28], or representations passed from the previous neural layers, as equivariant feature $\textbf{f}^{(k)}$ in the current hidden layer.
+Next, we construct the feature $\textbf{v}^{(k)}=g_{\text{nonlin}}(\textbf{f}^{(k)})$, to achieve sufficient non-linear expressiveness while maintaining SO(3)-equivariance, where $g_{\text{nonlin}}(\cdot)$ is the function defined in Eq.
+(5).
+$s_{nonlin}(\cdot)$, i.e., $g_{\text{nonlin}}(\cdot)$’s built-in non-linear mechanism, which can be any differentiable non-linear neural module, is implemented here as a feed-forward module with a hidden layer size of $1024$, using Layer Normalization for stabilization and regularization, and employing SiLU as the non-linear mapping.
+To facilitate learning and optimization, we connect the input and output of $g_{\text{nonlin}}(\cdot)$ in a residual manner [29], serving as the output of the current encoding module, i.e., $\textbf{o}^{(k)}=\textbf{f}^{(k)}+\textbf{v}^{(k)}$.
+$\textbf{o}^{(k)}$ is then fed into the next encoder module to yield $\textbf{f}^{(k+1)}$, $\textbf{o}^{(k+1)}$, …., until the last encoder module, which is followed by the decoder modules.
+For an expressive feature encoder handling a complex physical system, $\textbf{f}^{(k)}$ is usually not a feature of single degree but a direct-sum concatenation of series of components $\textbf{f}^{(k)l_{1}}$, $\textbf{f}^{(k)l_{2}}$, …, $\textbf{f}^{(k)l_{k}}$ at multiple degrees, i.e., $l_{1}$, $l_{2}$, …, $l_{k}$, where some components share the same degree while others differ.
+In this case, it becomes necessary to extend the decomposition operator, i.e., $\text{CGDecomp}(\cdot)$ in Eq.
+(5), to accommodate various components of degrees.
+Moreover, in the context of neural networks, introducing learnable parameters $W$ may improve the model capacity.
+Based on these considerations, when constructing the encoding module, the $\text{CGDecomp}(\cdot)$ operation in Eq.
+(5) can be expanded as $\text{CGDecomp}_{ext}(\cdot)$:
+$u^{(k)}=\text{CGDecomp}_{\text{ext}}(\textbf{f}^{(k)}\otimes\textbf{f}^{(k)},0,W)=\sum_{1\leq i,j\leq k,l_{i}=l_{j}}W_{ij}\cdot\text{CGDecomp}(\textbf{f}^{(k)l_{i}}\otimes\textbf{f}^{(k)l_{j}},0)$ (6)
+where $W$ represents learnable parameters.
+This expansion does not change the equivariance or invariance of the features.
+It is obvious that, the output ($u^{(k)}$) of $\text{CGDecomp}_{\text{ext}}(\cdot)$ is also SO(3)-invariant, just like the original $\text{CGDecomp}(\cdot)$.
+Thus, even with such an expansion, the feature $z^{(k)}=s_{\text{nonlin}}(u^{(k)})$ remains SO(3)-invariant and $\textbf{v}^{(k)}=\frac{\partial z^{(k)}}{\partial\mathbf{f}^{(k)}}$ remain SO(3)-equivariant.
+[…]
+Since complex non-linear transformations have already been incorporated in the encoding framework, the decoders can be simple: the SO(3)-equivariant decoder is implemented using an SO(3)-equivariant linear mapping [27], and the SO(3)-invariant decoder is implemented as a shallow feed-forward module.
+6.2 Training and Inference
+The training loss function is shown as the following:
+$\begin{split}\min_{\theta}loss&=loss_{Q}+\mu(loss_{Q},loss_{T})\cdot loss_{T},\\ loss_{Q}&=MSE(\widehat{\mathbf{Q}},\mathbf{Q}^{*}),\\ loss_{T}&=MSE(\widehat{\mathbf{T}},\mathbf{T}^{*})\end{split}$ (7)
+where $\theta$ denotes all of the learnable parameters of our framework, $\widehat{\mathbf{Q}}$, $\widehat{\mathbf{T}}$ and $\mathbf{Q}^{*}$, $\mathbf{T}^{*}$ respectively denote the predictions and labels of $\mathbf{Q}$ and $\mathbf{T}$.
+In order that the both of the SO(3)-equivariant and SO(3)-invariant branches can be stably converged, we introduce $\mu(loss_{Q},loss_{T})$, i.e., a function returning a coefficient to regularize the relative scale of the two loss terms:
+$\mu(loss_{Q},loss_{T})=\lambda\cdot No\_Grad(\frac{loss_{Q}}{loss_{T}})$ (8)
+where $No\_Grad(\cdot)$ denotes gradient discarding when calculating such coefficient, as this coefficient is only used for adjusting the relative scale of the two loss terms and should not itself be a source of training gradients, otherwise it would counteract $loss_{T}$ in Eq.
+(7).
+The introduction of $\mu(loss_{Q},loss_{T})$ effectively prevents the numerical disparity between the two loss terms to stabilize the training for both of them.
+All of the encoding and decoding modules are trained jointly by Eq.
+(7), and the training configurations are included in Supplementary Information D.
+In our framework, the SO(3)-invariant non-linear features $z^{(k)}(1\leq k\leq K)$ learned with the help of $\mathbf{T}$ contribute to the learning of SO(3)-equivariant representations $\textbf{v}^{(k)}(1\leq k\leq K)$ and the prediction of $\mathbf{Q}$.
+To achieve this goal, in the training phase, both the SO(3)-invariant feature encoder and the SO(3)-invariant decoder, used for predicting $\mathbf{T}$ from $z^{(k)}$, need to be trained.
+While in the inference phase, since $\mathbf{T}$ is not the final regression target, the SO(3)-invariant decoder can be removed, yet the SO(3)-invariant encoder still should be retained, used to induce SO(3)-equivariant features.
+6.3 Time Complexity
+Typically, one of the key advantage of deep learning algorithms for electronic-structure calculation lies in their low computational complexity which can asymptotically scale as $O(N)$, where $N$ is the number of atoms of a system.
+This is a significant efficiency advantage compared to the $O(N^{3})$ computational complexity of traditional DFT method, enabling efficient and low-cost prediction of electronic structures for large-scale atomic systems.
+Our method improves the accuracy of Hamiltonian predictions within the deep learning paradigm while maintaining this advantage of asymptotic complexity.
+The baseline models we select, whether DeepH-E3 or QHNet, are efficient models with an asymptotic computational complexity of $O(N)$ as $N$ increases.
+The neural module $g_{nonlin}(\cdot)$ we introduce efficiently operates on each individual feature f from the baseline models to obtain the corresponding non-linear SO(3)-equivariant feature v, rather than relying on complex interactions between multiple features to derive new ones.
+This ensures that the computational complexity remains asymptotically $O(N)$, with good potential to efficiently handle large atomic systems.
+7 Acknowledgements
+This work was supported by the National Natural Science Foundation of China (Grant Number 12134012) and the Strategic Priority Research Program of Chinese Academy of Sciences (Grant Number XDB0500201).
+Appendix A The Equivariance-Expressiveness Dilemma in Existing Deep Learning Methods
+In this section, we introduce existing efforts to harmonize the deep learning paradigm with SO(3)-equivariance, analyzing the dilemma between strict SO(3)-equivariance and non-linear expressive capabilities.
+This dilemma leads to an unsatisfactory trade-off between strict equivariance and non-linear expressiveness, ultimately limiting the achievable performance and scopes of application for these regression models.
+A.1 Approaches with Strict SO(3)-equivariance but Weak Expressiveness
+Most existing methods have focused on achieving SO(3)-equivariance at the neural mechanism level by applying group theory-guaranteed SO(3)-equivariant operators [26, 27].
+These operators, including linear scaling, element-wise sum, direct products, direct sums, and Clebsch-Gordan decomposition, have been used to build graph neural network architectures for tasks in 3D point cloud analysis [30], molecular dynamic simulation and property prediction [13, 12], as well as Hamiltonian prediction [31, 4, 8, 14].
+However, these operators could not provide sufficient expressiveness to fit complex non-linear mappings.
+The majority of these operators are linear operations, and the only one (direct product) that introduces some non-linearity can only approximate a simple squaring function, and cannot be extended to more complex non-linear forms due to the huge computational burden and difficulty in convergence.
+Even worse, directly applying highly-expressive non-linear activation functions, such as $Sigmoid$, $Softmax$, and $SiLU$, to SO(3)-equivariant features may result in the loss of equivariance.
+To alleviate this issue, these methods introduced a gated activation mechanism that processes SO(3)-invariant features ($l=0$) using non-linear activations and employs them as linear gating coefficients for SO(3)-equivariant features ($l\geq 1$), aiming to enhance their expressive power while maintaining strict equivariance.
+However, from the perspective of SO(3)-equivariant features, this mechanism is a combination of linear operations and cannot fundamentally increase their non-linear expressive capabilities.
+A.2 Approaches with Remarkable Expressiveness but Approximate SO(3)-equivariance
+Li et al.
+[6] proposed a local coordinate strategy to achieve SO(3)-equivariance.
+This strategy relied on the transformation relationship between local and global coordinates to achieve equivariance, without imposing restrictions on the neural network, allowing the use of network modules that possess non-linear expressiveness.
+However, this strategy, solely depending on coordinate transformations, only works under holistic rigid rotations of the atomic system.
+Without exploration of a neural mechanism for equivariance, when the local coordinates experience non-rigid perturbations such as thermal movements or bilayer twists, this strategy offers no guarantees on maintaining symmetry properties; Zitnick et al.
+[11] decomposed features into weighted sums of spherical harmonic basis functions, each corresponding to a discrete point on a unit sphere, through inner-product operations.
+Such discrete decomposition enables non-linear operations on the coefficients to enhance expressiveness.
+In subsequent developments [15, 16, 32], this approach has demonstrated a remarkable capacity to fit complex functions.
+However, the inner-product operations with discrete basis make such approach degenerate from continuous to discrete rotational equivariance, no longer maintaining strict equivariance to continuous rotational operations, as pointed out by Zhang et al.
+[1].
+Additionally, due to the high complexity of its discrete sampling strategies, tensor orders ($m$) corresponding to large degrees ($l$) are often partially pruned in architectures such as Equiformerv2 [16], further sacrificing the equivariance of features.
+In summary, although these methods with approximate SO(3)-equivariance achieve high numerical accuracy due to their extensive use of non-linear mechanisms, they are best suited for the applications that primarily require numerical approximations but not for tasks that demand theoretical depth in physics.
+In some applications, the demands for symmetry are extremely high.
+Even minor deviations from perfect symmetry can result in incorrect physical results.
+Combining these analyses, existing deep learning methods have compromised either strict SO(3)-equivariance or non-linear expressiveness when processing physical systems.
+In this paper, we develop a novel framework that theoretically combines strict SO(3)-equivariance with the non-linear expressiveness of neural networks.
+This solution resolves the equivariance-expressiveness dilemma by deeply exploring the intrinsic relationships between SO(3)-invariant and SO(3)-equivariant quantities and their representations.
+Appendix B Proof of Theory
+B.1 Proof of Theorem 1
 Proof.
@@
 $\mathbf{Q}(\mathbf{R})=\mathbf{D}^{l_{p}}(\mathbf{R})\cdot\mathbf{Q}\cdot\mathbf{D}^{l_{q}}(\mathbf{R})^{\dagger},$
-where $\mathbf{D}^{l_{p}}(\mathbf{R})$ and $\mathbf{D}^{l_{q}}(\mathbf{R})$ are the Wigner-D matrices for the respective orbital degrees of atoms $i$ and $j$, corresponding to the rotation $\mathbf{R}$.
+where $\mathbf{D}^{l_{p}}(\mathbf{R})$ and $\mathbf{D}^{l_{q}}(\mathbf{R})$ are the Wigner-D matrices for the degrees of $l_{p}$ and $l_{q}$, respectively, corresponding to the rotation $\mathbf{R}$.
 The conjugate transpose of the transformed quantity is:
@@
 Therefore, $\mathbf{T}=tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})$ is invariant under SO(3) transformations, its SO(3)-invariance is proved. ∎
-Theorem 2.
-The non-linear neural mapping defined as the following is SO(3)-equivariant:
-$\mathbf{v}=g_{nonlin}(\mathbf{f})=\frac{\partial(s_{nonlin}(\text{CGDecomp}(\mathbf{f}\otimes\mathbf{f},0)))}{\partial\mathbf{f}}$ (5)
-where $\mathbf{f}$ is an input SO(3)-equivariant feature with degree $l$ in the direct-sum state, $\otimes$ denotes the direct product operation, $\text{CGDecomp}(\cdot,0)$ refers to performing a Clebsch-Gordan decomposition of the tensor and returning the scalar component of degree 0, and $s_{nonlin}(\cdot)$ represents non-linear neural mappings, $\mathbf{v}$ is the features encoded by $g_{nonlin}(\cdot)$
+B.2 Proof of Theorem 2
 Proof.
@@
 First, according to group theory, $u=\text{CGDecomp}(\mathbf{f}\otimes\mathbf{f},0)$ is an SO(3)-invariant scalar because the degree-zero component from the Clebsch-Gordan decomposition is invariant under rotations.
-Since applying a non-linear operation to a SO(3)-invariant quantity does not change its invariance, $z=s_{\text{nonlin}}(u)$ is also SO(3)-invariant, independent to the specific form of $s_{\text{nonlin}}(\cdot)$.
+Since applying a non-linear operation to an SO(3)-invariant quantity does not change its invariance, $z=s_{\text{nonlin}}(u)$ is also SO(3)-invariant, independent to the specific form of $s_{\text{nonlin}}(\cdot)$.
 It formally holds that:
@@
 Here we utilize the property that $\mathbf{D}^{l}(\mathbf{R})^{-1}=\mathbf{D}^{l}(\mathbf{R})^{T}$.
-Since the representations of neural networks are generally real numbers, the Wigner-D matrix is also real unitary.
+Since the representations of neural networks are generally real numbers, the Wigner-D matrix is also real unitary  [^3] In Theorem 1 and Theorem 2, the Wigner-D matrices are in the complex and real fields, respectively, because the output space may be complex while the internal representations of neural networks are typically real.
+Although the internal representations of neural network modules including $g_{\text{nonlin}}(\cdot)$ are in the real field, they are still able to predict targets in complex field with equivariance.
+Literature [8] has provided mechanisms for converting the network outputs in the real field into regression targets with real and imaginary parts..
 Finally, we transpose the result back to a column vector:
 $\mathbf{v}(\mathbf{R})=g_{nonlin}(\mathbf{f}(\mathbf{R}))=(\frac{\partial z(\mathbf{R})}{\partial\mathbf{f}^{T}(\mathbf{R})})^{T}=(\frac{\partial z}{\partial\mathbf{f}^{T}}\cdot\mathbf{D}^{l}(\mathbf{R})^{T})^{T}=\mathbf{D}^{l}(\mathbf{R})\cdot\frac{\partial z}{\partial\mathbf{f}}=\mathbf{D}^{l}(\mathbf{R})\cdot\mathbf{v}$ (11)
-This proves that $g_{\text{nonlin}}(\cdot)$ is a SO(3)-equivariant non-linear operator: when applying its non-linearity to a SO(3)-equivariant feature $\mathbf{f}$, the output feature $\mathbf{v}$ remains SO(3)-equivariant.
+This proves that $g_{\text{nonlin}}(\cdot)$ is an SO(3)-equivariant non-linear operator: when applying its non-linearity to a SO(3)-equivariant feature $\mathbf{f}$, the output feature $\mathbf{v}$ remains SO(3)-equivariant.
 ∎
-Remark 1.
-In Theorem 1, the term $tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})$ is equal to the square of the Frobenius norm of the matrix $\mathbf{Q}$, i.e., $tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})=\|\mathbf{Q}\|_{F}^{2}=\sum_{m,n}|\mathbf{Q}_{mn}|^{2}$ The invariance of Frobenius norm under unitary transformations reflects the profound connections between modulus and invariance in physics.
-Remark 2.
-In Theorem 2, $g_{\text{nonlin}}(\cdot)$ can handle equivariant features of any degree and the returned feature $\mathbf{v}$ has the same degree $l$ as the input feature $\mathbf{f}$.
-$\mathbf{v}$ is endowed with non-linear expressiveness with the effect of $s_{nonlin}(\cdot)$, which can be any differentiable non-linear network module, including non-linear activation functions and normalization operations.
-The gradient operation $\frac{\partial z}{\partial\mathbf{f}}$ in $g_{\text{nonlin}}(\cdot)$ is used during the forward inference process to promote feature encoding, expanding the usage and significance of gradients within the neural network paradigm.
-It is worth noting that, although the internal representations of neural network modules including $g_{\text{nonlin}}(\cdot)$ are in the real field, they are still able to predict targets in complex field with equivariance.
-Literature [7] has provided mechanisms for converting the network outputs in the real field into regression targets with real and imaginary parts.
-Remark 3.
-$z$ in Equation (6) is the bridge connecting Theorem 1 and Theorem 2 in implementing a SO(3)-equivariant non-linear representation learning approach.
-Since $z$ is an SO(3)-invariant feature, it can be learned by $s_{\text{nonlin}}(\cdot)$, i.e., the neural mechanisms with non-linear expressive capability, through supervised learning incorporated with the SO(3)-invariant label $tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})$.
-Following this, by taking the gradient of $z$ with respect to $\mathbf{f}$, the non-linear expressiveness of $z$ is transferred to the equivariant feature $\mathbf{v}$, while maintaining strict SO(3)-equivariance as proved above.
-Subsequently, merging $\mathbf{v}$ and $\mathbf{f}$, and applying $g_{\text{nonlin}}(\cdot)$ in a stacked manner can yield rich SO(3)-equivariant non-linear representations.
-Figure 1: Methodological framework for learning SO(3)-equivariant representations with non-linear expressiveness to regress complex SO(3)-equivariant targets.
-Based on the new theory, we present a general method for learning non-linear representations that are equivariant to SO(3), as shown in Figure 1.
-The core of our method can be abstractly referred to as TraceGrad.
-At the label level, it incorporates an SO(3)-invariant trace quantity introduced in Theorem 1 as a supervisory signal for learning SO(3)-invariant features; at the representation level, it induces non-linear SO(3)-equivariant features from non-linear SO(3)-invariant features through gradient-based operator introduced in Theorem 2.
-For details of the method, please refer to Section 8.
-5 Application to the Hamiltonian Prediction Tasks
-The tasks of predicting electronic-structure Hamiltonian are chosen as the application study of our theoretical and methodological framework, on two motivations.
-On one hand, the Hamiltonian occupies a central position in quantum mechanics.
-Form the Hamiltonian, key physical quantities such as charge density, band structure, and Berry phase can be calculated, bringing deeply understanding of the electrical, optical, magnetic, and transport properties of atomic systems.
-On the other hand, the intrinsic symmetry with respect to the SO(3) group as well as the high-dimensional complexity (as shown in Figure 2) of the Hamiltonian provide a challenging touchstone of whether a deep learning method possesses both SO(3)-equivariance and strong expressive power.
-Figure 2: Illustration of the high-dimensional complexity of the Hamiltonian.
-In an atomic system, a Hamiltonian matrix is defined between each pair of atoms within each other’s cutoff radius, constituting a large cluster of Hamiltonian matrices for all atom pairs.
-Besides, each Hamiltonian matrix contains multiple basic blocks, with the $\mathbf{Q}^{l_{p}\otimes l_{q}}$ defined in Section 3 corresponding to a basic block $\mathbf{H}^{l_{p}\otimes l_{q}}$ here.
-Thus, for a Hamiltonian prediction model, the space of the regression target is typically huge.
-Experiments are conducted on six benchmark databases of atomic systems that have Hamiltonian annotations, all of which are established and released by previous literature [6, 7].
-These databases include:
-• Monolayer Graphene (MG).
-This structure consists of a single layer of carbon atoms in a hexagonal lattice.
-It is known for its exceptional thermal conductivity, electrical properties, and mechanical strength.
-The potential applications of monolayer graphene extend across electronics, energy storage devices, and composite materials due to its unique properties.
-In the benchmark of MG, the sizes of the training set, validation set, and testing set are 270, 90, and 90 respectively.
-• Monolayer MoS2 (MM), which is a single layer of molybdenum disulfide and part of the transition metal dichalcogenides (TMDCs) group.
-This kind of structures is a semiconductor with a direct bandgap, making it suitable for optoelectronics, photovoltaics, and semiconducting components.
-It also shows promise in applications requiring high sensitivity to chemicals, making it valuable for sensors.
-In the benchmark of MM, the sizes of the training set, validation set, and testing set are 300, 100, and 100 respectively.
-• Bilayer Graphene (BG), which consists of two layers of graphene stacked together.
-This kind of structures exhibits unique electronic properties such as tunable bandgap and enhanced conductivity when subjected to an electric field.
-Bilayer Graphene is being explored for use in advanced electronic devices, such as fast transistors and efficient batteries.
-In the benchmark database of BG, the sizes of the training set and validation set are 180 and 60 respectively, all of them are non-twisted samples; in the testing set, the numbers of non-twisted and twisted samples are 60 and 9 respectively.
-• Bilayer Bismuthene (BB), which consists of two layers of bismuth atoms arranged in a honeycomb-like structure.
-This kind of materials shows potential for topological insulating properties and has significant implications for quantum computing and spintronic devices due to its strong spin-orbit coupling.
-In the benchmark of BB, the sizes of the training set and validation set are 231 and 113 respectively, all of them are non-twisted samples; in the testing set, the numbers of non-twisted and twisted samples are 113 and 4 respectively.
-• Bilayer Bi2Te3 (BT), which is a compound of bismuth and tellurium, known for its thermoelectric properties.
-This kind of structures is particularly valuable in thermoelectric devices for power generation and refrigeration technologies where waste heat can be converted into electrical energy.
-In the benchmark of BT, the sizes of the training set and validation set are 204 and 38 respectively, all of them are non-twisted samples; in the testing set, the numbers of non-twisted and twisted samples are 12 and 2 respectively.
-• Bilayer Bi2Se3 (BS), another bismuth-tellurium compound, similar to Bi2Te3 but with selenium replacing tellurium.
-It also exhibits excellent thermoelectric properties and potential for topological insulator applications.
-This makes it suitable for energy harvesting applications and advanced electronic devices based on topological insulator technology.
-In the benchmark of BS, the sizes of the training set and validation set are 231 and 113 respectively, all of them are non-twisted samples; in the testing set, the numbers of non-twisted and twisted samples are 113 and 2 respectively.
-These databases are both diverse and representative.
-They encompass both strong chemical bonds between single-layer atoms and weaker van der Waals interactions between bilayers, and also cover varying degrees of Spin-Orbit Coupling (SOC) effects: structures like BT, BB, and BS exhibit strong SOC effects, while others have weak SOC.
-These atomic structures are highly valuable in the field of information and energy science and technology.
-Yet, accurately predicting their Hamiltonians poses significant challenges due to structural deformations from thermal movements and inter-layer twists, as depicted in Fig.
-3, where the twisted structures have garnered considerable attention as a research focus because of their potential to exhibit novel electrical properties and quantum topological behaviors [24, 25].
-During twisting, the relative rotation between atoms and their coordinate system introduces SO(3)-equivariant effects, while changes in orientation between two layers alter the vdW interactions.
-These dynamic conditions challenge both the equivariance and expressive capabilities of the regression model.
-In our experiments, predicting the Hamiltonian of twisted samples are even challenging, as we strictly adhere to the official division of training, validation, and testing sets as defined upon the dataset’s release [^1] https://github.com/Xiaoxun-Gong/DeepH-E3.
-In this division, twisted samples are only included in the testing set and are absent from the training set.
-This setup forces the regression model to solely rely on its generalization ability and the inherent SO(3)-equivariance to extrapolate the twisting conditions.
-In summary, these benchmarks present scientific challenges for a regression model, pushing the boundaries of modeling capabilities in predicting complex quantum behaviors under dynamic conditions.
-Figure 3: Illustration of challenging samples in the experimental databases, showing structural distortions of atomic systems due to thermal movements and twists between the two layers, which require a regression model to possess both strict SO(3)-equivariance and strong expressive capacity.
-6 Results and Analysis
-To validate the effectiveness of the proposed method, we compare results from the four setups:
-• Baseline model (DeepHE3 [7]): We select DeepHE3 as our baseline model since it represents the state-of-the-art (SOTA) among strictly SO(3)-equivariant approaches for predicting electronic-structure Hamiltonians.
-This choice facilitates further breakthroughs based on the SOTA foundation.
-DeepHE3 is trained using the mean squared error (MSE) loss of the Hamiltonian.
-In the subsequent experimental setups, DeepHE3 will serve as a base encoder, providing basic SO(3)-equivariant features that are handled by our method as more expressive representations.
-• DeepHE3+Trace: This experimental setup, an ablation term, only implements half part of our method.
-Specifically, it extends the architecture of DeepHE3 by adding our SO(3)-invariant encoding and decoding branches and using the trace quantity $\mathbf{T}=\text{tr}(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})=\text{tr}(\mathbf{H}^{l_{p}\otimes l_{q}}\cdot(\mathbf{H}^{l_{p}\otimes l_{q}})^{\dagger})$ to train them.
-As for ablation study, this setup does not include mechanisms for encoding non-linear SO(3)-equivariant features via the gradient-induced operator; instead, it directly uses the features outputted by DeepHE3 for Hamiltonian regression.
-• DeepHE3+Grad: This setup is also an ablation term and implements the other half part of our method in contrast to the previous ablation term.
-Specifically, it incorporates our SO(3)-invariant encoder branch as well as the gradient-induced operator to yield SO(3)-equivariant non-linear features based on the representations provided by the base encoders of DeepHE3 for Hamiltonian prediction.
-As for ablation study, this setup continues to use the single-task training pipeline of DeepHE3, predicting only the Hamiltonian without joint supervision training through the trace label.
-• DeepHE3+TraceGrad.
-This setup corresponds to a complete implementation of our framework.
-Extending beyond the baseline architecture and pipeline of DeepHE3, at the label level, we introduce the trace quantity to guide the learning of SO(3)-invariant features; Meanwhile, at the representation level, we leverage non-linear SO(3)-invariant features induced by the gradient operator to yield non-linear SO(3)-equivariant features for Hamiltonian prediction.
-The performance of interest for the electronic-structure Hamiltonian prediction task are SO(3)-equivariance and numerical accuracy.
-Our method has already been theoretically proven to strictly obey SO(3)-equivariance in Section 4; therefore, we just focus on evaluating numerical accuracy here.
-Four metrics designed by Yin et al.
-[16] are adopted, i.e., $MAE_{all}$, which calculates the Mean Absolute Error (MAE) averaged across all elements of the Hamiltonian matrices from all testing samples in the current database; $MAE^{cha}_{sample}$, which calculates the MAE on the most challenging testing samples where the baseline model (DeepHE3) performs the worst in the current database; $MAE_{block}$, which calculates the MAE for each basic block of the Hamiltonian matrix, averaged across all testing samples in the current database; $MAE^{cha}_{block}$, which corresponds to the MAE on the most challenging Hamiltonian block where the baseline model shows the poorest performance (with the largest $MAE_{block}$) in the current dataset.
-These metrics comprehensively reflect the accuracy performance, considering not only the average accuracy but also the accuracy on difficult samples and challenging blocks of the Hamiltonian matrices.
-We collect results from $10$ independently repeated experiments on each database and list them in Tables 1 and 2, including the values of $MAE_{all}$, $MAE^{cha}_{sample}$, and $MAE^{cha}_{block}$, for monolayer and bilayer structures, respectively.
-Considering the large volume of statistics on $MAE_{block}$ for all blocks, we just display the complete results in Figures 4 and 5 with the form of bar charts for DeepHE3 and DeepHE3+TraceGrad.
-The results for DeepHE3 listed in these tables and figures are directly taken from the reproduction of Yin et al.
-[16], a paper which is in fact previous work by our own team.
-Since the experimental conditions for DeepHE3 are exactly the same between the previous and current work of ours, we directly list the results from the previous paper here.
-As pointed out in [16], the reproduction was based on DeepHE3’s open-source project11footnotemark: 1, which is newer than its original publication, and the performance was slightly better than the original one.
-As we use a fixed random seed ($42$) for all of the experiments, we find that the standard deviation of each metric is very small (all less than $0.007$ meV) across the $10$ repeated experiments.
-Since the fluctuation is negligible, in the following tables and figures, we just present the mean values of $MAE_{all}$, $MAE^{cha}_{sample}$, and $MAE^{cha}_{block}$ across the repeated experiments, without listing the standard deviations of each entry to save space.
-Table 1: Comparison on the $MAE_{all}$, $MAE^{cha}_{sample}$ and $MAE^{cha}_{block}$ metrics (/meV) for the four experimental setups on the databases of monolayer structures.
+Appendix C Information of Experimental Databases
+In this part, we provide detailed information about the experimental databases, including the statistical information of the six databases from the DeepH series [6, 8], and the two databases from the QH9 series [19], listed in Table 5 and Table 6, respectively.
+Additionally, we visualize two types of challenging testing samples: samples with non-rigid deformation from thermal motions, as well as the bilayer samples with interlayer twists, which are shown in Fig.
+4 and Fig.
+5, respectively.
+Table 5: Statistic information of the six benchmark databases, i.e., Monolayer Graphene ($MG$), Monolayer MoS2 ($MM$), Bilayer Graphene ($BG$), Bilayer Bismuthene ($BB$), Bilayer Bi2Te3 ($BT$), Bilayer Bi2Se3 ($BS$), from the DeepH benchmark series [6, 8].
+SOC: effects of Spin-Orbit Coupling.
+$m$: number of samples in the current dataset; $a_{max}$: maximum number of atoms for a sample in the current dataset.
+$a_{min}$: minimum number of atoms for a sample in the current dataset.
+$nt$: non-twisted samples.
+$t$: twisted samples.
+Statistic Types MG MM BG BB BT BS
+Elements C Mo, S C Bi Bi, Te Bi, Se
+SOC weak weak weak strong strong strong
+Training ($nt$) $m$ 270 300 180 231 204 231
+$a_{max}$ 72 75 64 36 90 90
+$a_{min}$ 72 75 64 36 90 90
+Validation ($nt$) $m$ 90 100 60 113 38 113
+$a_{max}$ 72 75 64 36 90 90
+$a_{min}$ 72 75 64 36 90 90
+Testing ($nt$) $m$ 90 100 60 113 12 113
+$a_{max}$ 72 75 64 36 90 90
+$a_{min}$ 72 75 64 36 90 90
+Testing ($t$) $m$ 9 4 2 2
+$a_{max}$ 1084 244 130 190
+$a_{min}$ 28 28 70 70
+Table 6: Basic statistics of the two benchmark databases, QH9-Stable ($QS$) and QH9-Dynamic ($QD$), from the QH9 benchmark series [19].
+For the $QS$ database, the ’ood’ strategy from the official settings is used to split the training, validation, and test sets, ensuring that the atom number of samples do not overlap across the three subsets.
+For the $QD$ database, the ’mol’ strategy from the official settings is employed to split the training, validation, and test sets, ensuring that there are no thermal motion samples from the same temporal trajectory across the three subsets.
+These split strategies aim to assess the regression model’s extrapolation capability with respect to the number of atoms as well as the temporal trajectories.
+$m$: number of samples in the current dataset.
+$a_{max}$: maximum number of atoms for a sample in the current dataset.
+$a_{min}$: minimum number of atoms for a sample in the current dataset.
+Statistic Types QS QD
+Elements C, H, O, N, F C, H, O, N, F
+Training $m$ 104,001 79,900
+$a_{max}$ 20 19
+$a_{min}$ 3 10
+Validation $m$ 17,495 9,900
+$a_{max}$ 22 19
+$a_{min}$ 21 10
+Testing $m$ 9,335 10,100
+$a_{max}$ 29 19
+$a_{min}$ 23 10
+Figure 4: Visualization of testing samples exhibiting non-rigid deformations due to thermal motions
+Figure 5: Visualization of testing samples with interlayer twists.
+Appendix D Training Configurations
+$\lambda$ in the training loss function is set as $0.3$ according to parameter selection on the validation sets.
+The initial learning rates for experiments on the $MG$, $MM$, $BG$, $BB$, $BT$ and $BS$ databases are set as $0.003$, $0.005$, $0.003$, $0.005$, $0.004$, and $0.005$, respectively, the training batch size for them is set as $1$ and the optimizer is selected as Adam; the initial learning rate for experiments on the $QS$ and $QD$ databases is set as $0.0005$, the training batch size is set as $16$, and the optimizer is selected as AdamW.
+Appendix E Visualization of Block-level MAE Statistics
+The MAE metrics (denoted as $MAE^{H}_{block}$) for the two setups, i.e., DeepH-E3 and DeepH-E3+TraceGrad, across different blocks of the Hamiltonian matrix for six databases from the DeepH benchmark series are illustrated in Fig.
+6 and 7.
+Fig.
+6 presents the results for monolayer structures, while Fig.
+7 focuses on bilayer structures.
+From these figures, it can be observed that our method, TraceGrad, brings significant accuracy improvements over the baseline method, i.e., DeepH-E3, on most of the blocks of the Hamiltonian matrix, especially on those blocks where DeepH-E3 performs poorly.
+Figure 6: Visualization of $MAE^{H}_{block}$ on each basic block of the Hamiltonian matrices for the $MG$ and $MM$ databases.
+Figure 7: Visualization of $MAE^{H}_{block}$ on each basic block of the Hamiltonian matrices for the non-twisted (marked with superscripts $nt$) and twisted (marked with superscripts $t$) structures of $BG$, $BB$, $BT$ and $BS$.
+Appendix F Ablation Study
+We conduct fine-grained ablation study on the six databases from DeepH benchmark series, comparing results from the four setups:
+• DeepH-E3 [8]: the baseline model.
+• DeepH-E3+Trace: this experimental setup, an ablation term, only implements half part of our method.
+Specifically, it extends the architecture of DeepH-E3 by adding our SO(3)-invariant encoding and decoding branches and using the trace quantity $\mathbf{T}=\text{tr}(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})=\text{tr}(\mathbf{H}^{l_{p}\otimes l_{q}}\cdot(\mathbf{H}^{l_{p}\otimes l_{q}})^{\dagger})$ to train them.
+As for ablation study, this setup does not include the gradient-based mechanism for encoding SO(3)-equivariant non-linear features; instead, it directly uses the SO(3)-equivariant features outputted by DeepH-E3 for Hamiltonian regression.
+• DeepH-E3+Grad: this setup is also an ablation term and implements the other half part of our method in contrast to the previous ablation term.
+Specifically, it incorporates our SO(3)-invariant encoder branch as well as the gradient-induced operator to yield SO(3)-equivariant non-linear features based on the representations provided by the baseline encoders of DeepH-E3 for Hamiltonian prediction.
+As for ablation study, this setup continues to use the single-task training pipeline of DeepH-E3, predicting only the Hamiltonian without joint supervision training through the trace label.
+• DeepH-E3+TraceGrad: this is a complete implementation of our framework extending beyond of the architecture and training pipeline of DeepH-E3, at the label level, we introduce the trace quantity to guide the learning of SO(3)-invariant features; Meanwhile, at the representation level, we leverage the gradient operator to yield SO(3)-equivariant non-linear features for Hamiltonian prediction.
+Table 7: Ablation study MAE results (meV) on the $MG$ and $MM$ databases. $\downarrow$ means lower values of the metrics correspond to better accuracy.
 Methods $MG$ $MM$
-$MAE_{all}$ $MAE^{cha}_{sample}$ $MAE^{cha}_{block}$ $MAE_{all}$ $MAE^{cha}_{sample}$ $MAE^{cha}_{block}$
-DeepHE3 (Baseline) 0.251 0.357 0.362 0.406 0.574 1.103
-DeepHE3+Trace 0.230 0.344 0.348 0.378 0.537 1.091
-DeepHE3+Grad 0.185 0.269 0.258 0.308 0.453 0.924
-DeepHE3+TraceGrad 0.175 0.257 0.228 0.285 0.412 0.808
-Table 2: Comparison on the $MAE_{all}$, $MAE^{cha}_{sample}$ and $MAE^{cha}_{block}$ metrics (/meV) for the four experimental setups on the databases of bilayer structures, where the superscripts $nt$ and $t$ respectively denote the non-twisted and twisted subsets.
+$MAE$ ()($\downarrow$)( ↓ )
+$MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$ $MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$
+DeepH-E3 (Baseline) 0.251 0.357 0.362 0.406 0.574 1.103
+DeepH-E3+Trace 0.230 0.344 0.348 0.378 0.537 1.091
+DeepH-E3+Grad 0.185 0.269 0.258 0.308 0.453 0.924
+DeepH-E3+TraceGrad 0.175 0.257 0.228 0.285 0.412 0.808
+Table 8: Ablation study MAE results (meV) on $BG$, $BB$, $BT$and $BS$.
+The superscripts $nt$ and $t$ respectively denote the non-twisted and twisted subsets.
+$\downarrow$ means lower values of the metrics correspond to better accuracy.
 Methods $BG^{nt}$ $BG^{t}$
-$MAE_{all}$ $MAE^{cha}_{sample}$ $MAE^{cha}_{block}$ $MAE_{all}$ $MAE^{cha}_{sample}$ $MAE^{cha}_{block}$
-DeepHE3 (Baseline) 0.389 0.453 0.644 0.264 0.429 0.609
-DeepHE3+Trace 0.362 0.417 0.593 0.251 0.401 0.480
-DeepHE3+Grad 0.320 0.356 0.511 0.222 0.389 0.446
-DeepHE3+TraceGrad 0.291 0.323 0.430 0.198 0.372 0.406
+$MAE$ ()($\downarrow$)( ↓ )
+$MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$ $MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$
+DeepH-E3 (Baseline) 0.389 0.453 0.644 0.264 0.429 0.609
+DeepH-E3+Trace 0.362 0.417 0.593 0.251 0.401 0.480
+DeepH-E3+Grad 0.320 0.356 0.511 0.222 0.389 0.446
+DeepH-E3+TraceGrad 0.291 0.323 0.430 0.198 0.372 0.406
 Methods $BB^{nt}$ $BB^{t}$
-$MAE_{all}$ $MAE^{cha}_{sample}$ $MAE^{cha}_{block}$ $MAE_{all}$ $MAE^{cha}_{sample}$ $MAE^{cha}_{block}$
-DeepHE3 (Baseline) 0.274 0.304 1.042 0.468 0.602 2.399
-DeepHE3+Trace 0.259 0.285 0.928 0.429 0.570 1.782
-DeepHE3+Grad 0.243 0.272 0.824 0.406 0.542 1.431
-DeepHE3+TraceGrad 0.226 0.256 0.740 0.384 0.503 1.284
+$MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$ $MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$
+DeepH-E3 (Baseline) 0.274 0.304 1.042 0.468 0.602 2.399
+DeepH-E3+Trace 0.259 0.285 0.928 0.429 0.570 1.782
+DeepH-E3+Grad 0.243 0.272 0.824 0.406 0.542 1.431
+DeepH-E3+TraceGrad 0.226 0.256 0.740 0.384 0.503 1.284
 Methods $BT^{nt}$ $BT^{t}$
-$MAE_{all}$ $MAE^{cha}_{sample}$ $MAE^{cha}_{block}$ $MAE_{all}$ $MAE^{cha}_{sample}$ $MAE^{cha}_{block}$
-DeepHE3 (Baseline) 0.447 0.480 1.387 0.831 0.850 4.572
-DeepHE3+Trace 0.406 0.462 1.239 0.784 0.812 4.520
-DeepHE3+Grad 0.342 0.365 0.750 0.742 0.786 4.463
-DeepHE3+TraceGrad 0.295 0.312 0.718 0.735 0.755 4.418
+$MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$ $MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$
+DeepH-E3 (Baseline) 0.447 0.480 1.387 0.831 0.850 4.572
+DeepH-E3+Trace 0.406 0.462 1.239 0.784 0.812 4.520
+DeepH-E3+Grad 0.342 0.365 0.750 0.742 0.786 4.463
+DeepH-E3+TraceGrad 0.295 0.312 0.718 0.735 0.755 4.418
 Methods $BS^{nt}$ $BS^{t}$
-$MAE_{all}$ $MAE^{cha}_{sample}$ $MAE^{cha}_{block}$ $MAE_{all}$ $MAE^{cha}_{sample}$ $MAE^{cha}_{block}$
-DeepHE3 (Baseline) 0.397 0.424 0.867 0.370 0.390 0.875
-DeepHE3+Trace 0.382 0.397 0.843 0.351 0.367 0.838
-DeepHE3+Grad 0.343 0.365 0.696 0.324 0.339 0.746
-DeepHE3+TraceGrad 0.300 0.332 0.644 0.291 0.302 0.674
-Figure 4: $MAE_{\text{block}}$ on each basic block of the Hamiltonian matrices of monolayer structures.
-Combining the results from Table 1, Table 2, Figure 4 and Figure 5, we find that the proposed TraceGrad method significantly enhances the accuracy performance of the baseline method DeepHE3, both on average and for challenging samples or blocks, both on the non-twisted samples and the twisted samples marking strong difficulty for the expressive power and generalization ability of regression models as analyzed in Section 5.
-Specifically, TraceGrad lowers down the $MAE_{all}$ of DeepHE3 with a relative ratio of $30.28\%$, $29.80\%$, $25.19\%$, $25.00\%$, $17.52\%$, $17.95\%$, $34.00\%$, $11.55\%$, $24.43\%$, and $21.35\%$, on MG, MM, BGnt, BGt, BBnt, BBt, BTnt, BTt, BSnt, and BSt, respectively; and lowering down $MAE_{cha}$ by $28.01\%$, $28.22\%$, $28.70\%$, $13.29\%$, $15.79\%$, $16.45\%$, $35.00\%$, $11.18\%$, $21.70\%$, and $22.56\%$, on the respective datasets; and achieving better results in the vast majority of basic blocks.
-Particularly, for the blocks where DeepHE3 performs the worst, TraceGrad reduces the $MAE$ ($MAE^{cha}_{block}$) by $37.01\%$, $26.75\%$, $33.23\%$, $33.33\%$, $28.98\%$, $46.48\%$, $48.23\%$, $3.37\%$, $25.72\%$, and $22.97\%$, on the respective datasets, from which the $MAE^{cha}_{block}$ in MM, BGnt and BTnt are lowered down to within $1meV$; the $MAE^{cha}_{block}$ in BBt is decreased by nearly half to approach $1meV$.
-This indicates that the proposed representation learning framework, while maintaining strict SO(3)-equivariance, significantly enhances the expressiveness of equivariant representations, thereby showcasing immense advantages in capturing the complexity of physical systems and refreshing the SOTA accuracy on the six benchmark databases.
-Figure 5: $MAE_{\text{block}}$ on each basic block of the Hamiltonian matrix in non-twisted (marked by superscripts $nt$) and twisted (marked by superscripts $t$) bilayer structures.
+$MAE_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$ $MAE_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$
+DeepH-E3 (Baseline) 0.397 0.424 0.867 0.370 0.390 0.875
+DeepH-E3+Trace 0.382 0.397 0.843 0.351 0.367 0.838
+DeepH-E3+Grad 0.343 0.365 0.696 0.324 0.339 0.746
+DeepH-E3+TraceGrad 0.300 0.332 0.644 0.291 0.302 0.674
+Experimental results of the four setups are listed in Table 7 and 8.
 From the results of ablation terms, we can obtain a more fine-grained experimental analysis.
-By comparing among the results of DeepHE3, DeepHE3+Trace, DeepHE3+Grad, and DeepHE3+TraceGrad, we can conclude that the two core mechanisms of our method, i.e., the SO(3)-invariant trace supervision mechanism (Trace) at the label level as well as the gradient-based induction mechanism (Grad) at the representation layer, can contribute to the performance individually.
+By comparing among the results of DeepH-E3, DeepH-E3+Trace, DeepH-E3+Grad, and DeepH-E3+TraceGrad, we can conclude that the two core mechanisms of our method, i.e., the SO(3)-invariant trace supervision mechanism (Trace) at the label level as well as the gradient-based induction mechanism (Grad) at the representation layer, can contribute to the performance individually.
 Moreover, their combination provides even better performance.
-This is because, on one hand, with the gradient-based induction mechanism as a bridge, the SO(3)-invariant features learned from the trace label can be transformed into the SO(3)-equivariant representations during inference; on the other hand, with trace label, the SO(3)-invariant network branch has a strong supervisory signal, enabling it to learn the intrinsic symmetry and complexity of the regression targets, which enhances the quality of SO(3)-invariant features and ultimately benefits the encoding of SO(3)-equivariant features.
-The value of such complementarity has been fully demonstrated in the experimental results.
-The main focus of this work is to bridge rigorous SO(3)-equivariance with the non-linear expressive capabilities of neural networks.
-While approximately SO(3)-equivariant techniques, which limit the depth of theoretical exploration, are not the focus of this study, they remain of interest in a few numerical computation applications where precision is highlighted more over strict equivariance.
-Considering this, in the appendix, we explore the synergy of our methods with the Transformer-based approximately SO(3)-equivariant technique [16], and the experimental results also demonstrate a significant accuracy improvement brought by our approach, further supporting the generality of our methodology.
-7 Future Work
-In future research, various extensions are conceivable across theoretical, methodological, and application fields.
-Theoretically, we can generalize SO(3) group symmetry to a broader range and more complex groups.
-This will enable deep learning frameworks to accommodate richer mathematical structures.
-From a methodological perspective, this study employs the DeepHE3 framework for comparative purposes.
-However, it is feasible to adopt a more sophisticated network, such as the Transformer-based SO(3)-equivariant non-linear representation learning framework, which utilizes the general TraceGrad methodology.
-The current methodology extends beyond predicting electronic Hamiltonians and is capable of generally predicting multiple physical quantities and properties that require equivariance.
-Furthermore, the methodology is not confined to physical research alone and may find applications in areas such as robotics, autonomous vehicles, and motion tracking systems.
-It handles 3D point cloud data and other vision information to achieve safer and more reliable prediction results with SO(3)-symmetry.
-Looking forward to future endeavors, the foundational and pivotal role of this work is clear.
-8 Method
-8.1 Encoding and Decoding Framework
-As shown in Figure 1, our encoding framework corresponds to a total of $K$ encoding modules, which are sequentially connected to form a cascading stack.
-For the $k$-th module ($1\leq k\leq K$), it first introduces an SO(3)-equivariant base encoder, e.g., the encoder from DeepHE3 [7] which is composed of recently developed equivariant operators [19, 20] like linear scaling, direct products, direct sums, Clebsch-Gordan decomposition, gated activation, equivariant normalization, and etc., to encode the physical system’s initial representations, such as spherical harmonics [26], or representations passed from the previous neural layers, as equivariant features $\textbf{f}^{(k)}$ in the current hidden layer.
-Next, we construct the feature $\textbf{v}^{(k)}=g_{\text{nonlin}}(\textbf{f}^{(k)})$, to achieve sufficient non-linear expressiveness while maintaining SO(3)-equivariance, where $g_{\text{nonlin}}(\cdot)$ is the function defined in Equation (5).
-As for $s_{nonlin}(\cdot)$, i.e., the built-in non-linear mechanism within $g_{\text{nonlin}}(\cdot)$, although it can be implemented as any differentiable non-linear neural module, here for simplicity, we implement it as a fully connected feed-forward structure containing the SiLU activation function and layer normalization.
-To facilitate learning and optimization, we connect the input and output of $g_{\text{nonlin}}(\cdot)$ in a residual manner [27], serving as the output of the current encoding module, i.e., $\textbf{o}^{(k)}=\textbf{f}^{(k)}+\textbf{v}^{(k)}$.
-$\textbf{o}^{(k)}$ is then fed into the next encoder module to yield $\textbf{f}^{(k+1)}$, $\textbf{o}^{(k+1)}$, …., until the last encoder module, which is followed by the decoder modules.
-For an expressive feature encoder handling a complex physical system, $\textbf{f}^{(k)}$ is usually not a feature of single degree but a direct-sum concatenation of series of components $\textbf{f}^{(k)l_{1}}$, $\textbf{f}^{(k)l_{2}}$, …, $\textbf{f}^{(k)l_{k}}$ at multiple degrees, i.e., $l_{1}$, $l_{2}$, …, $l_{k}$, where some components share the same degree while others differ.
-In this case, it becomes necessary to extend the decomposition operator, i.e., $\text{CGDecomp}(\cdot)$ in Equation (5), to accommodate various components of degrees.
-Moreover, in the context of neural networks, introducing learnable parameters $W$ may improve the model capacity.
-Based on these considerations, when constructing the encoding module, the $\text{CGDecomp}(\cdot)$ operation in Equation (5) can be expanded as $\text{CGDecomp}_{ext}(\cdot)$:
-$u^{(k)}=\text{CGDecomp}_{\text{ext}}(\textbf{f}^{(k)}\otimes\textbf{f}^{(k)},0,W)=\sum_{1\leq i,j\leq k,l_{i}=l_{j}}W_{ij}\cdot\text{CGDecomp}(\textbf{f}^{(k)l_{i}}\otimes\textbf{f}^{(k)l_{j}},0)$ (9)
-where $W$ represents learnable parameters.
-This expansion does not change the equivariance or invariance of the features.
-It is obvious that, the output ($u^{(k)}$) of $\text{CGDecomp}_{\text{ext}}(\cdot)$ is also SO(3)-invariant, just like the original $\text{CGDecomp}(\cdot)$.
-Thus, even with such an expansion, the features $z^{(k)}=s_{\text{nonlin}}(u^{(k)})$ remain SO(3)-invariant and $\textbf{v}^{(k)}=\frac{\partial z^{(k)}}{\partial\mathbf{f}^{(k)}}$ remain SO(3)-equivariant.
-The outputted SO(3)-equivariant features, i.e., $\textbf{o}^{(K)}$, as well as the outputted SO(3)-invariant features, i.e., $\textbf{z}^{(K)}$, from the final ($K$ th) layer of the encoding modules, are fed into the corresponding decoding modules to yield the predictions of $\mathbf{Q}$ and $\mathbf{T}=tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})$, respectively.
-Since complex non-linear transformations have already been incorporated in the encoding framework, the decoders can be simple: the SO(3)-equivariant decoder is implemented using an SO(3)-equivariant linear mapping [20], and the SO(3)-invariant decoder is implemented as a shallow feed-forward module.
-8.2 Training and Inference
-The training loss function is shown as the following:
-$\begin{split}\min_{\theta}loss&=loss_{Q}+\mu(loss_{Q},loss_{T})\cdot loss_{T},\\ loss_{Q}&=MSE(\widehat{\mathbf{Q}},\mathbf{Q}^{*}),\\ loss_{T}&=MSE(\widehat{\mathbf{T}},\mathbf{T}^{*})\end{split}$ (10)
-where $\theta$ denote all of the learnable parameters of our framework, $\widehat{\mathbf{Q}}$, $\widehat{\mathbf{T}}$ and $\mathbf{Q}^{*}$, $\mathbf{T}^{*}$ respectively denote the predictions and labels of $\mathbf{Q}$ and $\mathbf{T}$.
-In order that the both of the SO(3)-equivariant and SO(3)-invariant branches can be stably converged, we introduce $\mu(loss_{Q},loss_{T})$, i.e., a function returning a coefficient to regularize the relative scale of the two loss terms:
-$\mu(loss_{Q},loss_{T})=\lambda\cdot No\_Grad(\frac{loss_{Q}}{loss_{T}})$ (11)
-where $No\_Grad(\cdot)$ denotes gradient discarding when calculating such coefficient, as this coefficient is only used for adjusting the relative scale of the two loss terms and should not itself be a source of training gradients, otherwise it would counteract $loss_{T}$ in Equation (10).
-$\lambda$ here is a hyperparameter, whose value is determined as $0.3$ based on the performance on the validation set.
-The introduction of $\mu(loss_{Q},loss_{T})$ effectively prevents the numerical disparity between the two loss terms to stabilize the training for both of them.
-In our framework, the SO(3)-invariant non-linear features $z^{(k)}(1\leq k\leq K)$ learned with the help of $\mathbf{T}$ contribute to the learning of SO(3)-equivariant representations $\textbf{v}^{(k)}(1\leq k\leq K)$ and the prediction of $\mathbf{Q}$.
-To achieve this goal, in the training phase, both the SO(3)-invariant feature encoder and the SO(3)-invariant decoder, used for predicting $\mathbf{T}$ from $z^{(k)}$, need to be trained.
-While in the inference phase, since $\mathbf{T}$ is not the final regression target, the SO(3)-invariant decoder can be removed, yet the SO(3)-invariant encoder still should be retained, used to induce SO(3)-equivariant features.
-9 Acknowledgements
-This work was supported by the National Natural Science Foundation of China (Grant Number 12134012) and the Strategic Priority Research Program of Chinese Academy of Sciences (Grant Number XDB0500201).
-Appendix A Empirical Study on Combining our Method with Approximately SO(3)-equivariance Framework
-Table 3: Comparison on the $MAE_{all}$, $MAE^{cha}_{sample}$ and $MAE^{cha}_{block}$ metrics for approximately SO(3)-equivariant methods on the monolayer MoS2 ($MM$) database.
-The results of the compared methods are copied from their papers, where the empty items are due to the data not being provided in the original text.
-Methods $MM$
-$MAE_{all}$ $MAE^{cha}_{sample}$ $MAE^{cha}_{block}$
-DeepH-2 [23] 0.21 - -
-HarmoSE [16] 0.233 0.293 0.406
-HarmoSE+TraceGrad 0.178 0.228 0.296
-Figure 6: Comparison on the $MAE_{block}$ metric for approximately SO(3)-equivariant methods on the monolayer MoS2 ($MM$) database.
-In this part, taking the MoS2 dataset as a case study, we evaluate the performance of combining our trace supervision and gradient induction method (TraceGrad) with the approximately equivariant approach HarmoSE [16].
-We here take HarmoSE as the backbone encoder, and yields features by TraceGrad to enrich its representations.
-The experimental results in Table 3 and Figure 6 demonstrate that TraceGrad significantly enhances the accuracy of HarmoSE, surpassing DeepH-2 and achieving new SOTA results.
-Both DeepH-2 and HarmoSE sacrificed strict SO(3)-equivariance to fully release the expressive capabilities of graph Transformers, aiming for the ultimate in prediction accuracy.
-Despite this, TraceGrad still manages to significantly exceed their accuracy, further confirming the superiority and generality of our method in learning expressive representations of physical systems.

### v4

@@
 Abstract
-We present both a theoretical and a methodological framework that addresses a critical challenge in applying deep learning to physical systems: the reconciliation of non-linear expressiveness with SO(3)-equivariance in predictions of SO(3)-equivariant quantities.
-Inspired by covariant theory in physics, we address this problem by exploring the mathematical relationships between SO(3)-invariant and SO(3)-equivariant quantities and their representations.
+We propose both a theoretical and a methodological framework to address a critical challenge in applying deep learning to physical systems: the reconciliation of non-linear expressiveness with SO(3)-equivariance in predictions of SO(3)-equivariant quantities.
+Inspired by covariant theory in physics, we present a solution by exploring the mathematical relationships between SO(3)-invariant and SO(3)-equivariant quantities and their representations.
 We first construct theoretical SO(3)-invariant quantities derived from the SO(3)-equivariant regression targets, and use these invariant quantities as supervisory labels to guide the learning of high-quality SO(3)-invariant features.
 Given that SO(3)-invariance is preserved under non-linear operations, the encoding process for invariant features can extensively utilize non-linear mappings, thereby fully capturing the non-linear patterns inherent in physical systems.
-Building on this foundation, we propose a gradient-based mechanism to induce SO(3)-equivariant encodings of various degrees from the learned SO(3)-invariant features.
-This mechanism can incorporate non-linear expressive capabilities into SO(3)-equivariant representations, while theoretically preserving their equivariant properties as we prove.
-We apply our theory and method to the electronic-structure Hamiltonian prediction tasks, experimental results on eight benchmark databases covering multiple types of elements and challenging scenarios show dramatic breakthroughs on the state-of-the-art prediction accuracy, with improvements of up to 40% in predicting Hamiltonians and up to 76% in predicting downstream physical quantities such as occupied orbital energy.
-Our approach goes beyond handling physical systems and offers a promising general solution to the critical dilemma between equivariance and non-linear expressiveness for the deep learning paradigm.
-keywords:
-Deep Learning, Neural Networks, SO(3)-equivariance, Atomic Modeling, Hamiltonian Prediction
+Building on this, we propose a gradient-based mechanism to induce SO(3)-equivariant encodings of various degrees from the learned SO(3)-invariant features.
+This mechanism can incorporate non-linear expressive capabilities into SO(3)-equivariant representations, while theoretically preserving their equivariant properties as we prove, establishing a strong foundation for regressing complex SO(3)-equivariant targets.
+We apply our theory and method to the electronic-structure Hamiltonian prediction tasks, experimental results on eight benchmark databases covering multiple types of systems and challenging scenarios show substantial improvements on the state-of-the-art prediction accuracy of deep learning paradigm.
+Our method boosts Hamiltonian prediction accuracy by up to 40% and enhances downstream physical quantities, such as occupied orbital energy, by a maximum of 76%.
+Our method also significantly promotes the acceleration ratios for the convergence of traditional Density Functional Theory (DFT) methods.
 1 Introduction
-With the advantages in computational complexity and generalization capabilities, deep learning paradigm has vigorously driven advancements in physics research [1].
-For example, in predicting electronic-structure Hamiltonians, traditional Density Functional Theory (DFT) methods [2, 3] suffer from computational complexities of $\mathcal{O}(N^{3})$, where $N$ is the number of atoms in a system.
-In contrast, deep learning approaches have significantly reduced this complexity to $\mathcal{O}(N)$ [4, 5, 6, 7, 8], accelerating processes by thousands or even millions of times.
-This reduction in computational demand not only unlocks possibilities for analyzing extremely large atomic systems but also paves the way for efficient material simulation, discovery, and design that were previously unimaginable.
+With the advantages in computational complexity and generalization capabilities, deep learning paradigm has vigorously driven advancements in physics research (Zhang et al., 2023).
+For example, in predicting electronic-structure Hamiltonians, traditional Density Functional Theory (DFT) methods (Hohenberg & Kohn, 1964; Kohn & Sham, 1965) suffer from computational complexities of $\mathcal{O}(N^{3})$, where $N$ is the number of atoms in a system.
+In contrast, deep learning approaches have significantly reduced this complexity to $\mathcal{O}(N)$ (Unke et al., 2021; Gu et al., 2022; Li et al., 2022; Yu et al., 2023b; Gong et al., 2023), opening up new possibilities for analyzing extremely large atomic systems, enabling efficient materials simulation and design, as well as groundbreaking molecular pharmaceutical research that was previously unimaginable.
 However, deep learning methods still face substantial challenges when processing physical systems.
@@
 For example, physical quantities such as force fields and electronic-structure Hamiltonians must be equivariant under 3D rotational transformations, i.e., elements from the SO(3) group.
-Traditional deep learning methods [9, 10] typically approximated SO(3)-equivariance through rotational data augmentation.
-However, this approach does not provide the strict assurance of SO(3)-equivariance required for physical research.
 Besides, the calculation of physical quantities calls for high numerical accuracy, necessitating that the neural networks possess strong capabilities to express the complex non-linear mappings from atomic structures to the regression targets.
-However, existing deep learning methods are unable to simultaneously ensure strict SO(3)-equivariance and high numerical accuracy.
+However, it is challenging for deep learning methods to simultaneously ensure strict SO(3)-equivariance and high numerical accuracy on modeling physical systems.
 The root cause of this problem lies in the conflict between SO(3)-equivariance and non-linear expressiveness: specifically, directly applying non-linear activation functions on SO(3)-equivariant features (with degree $l\geq 1$) may lead to the loss of equivariance, while bypassing non-linear mappings severely restricts the network’s expressive capabilities and thereby lowering down the achievable accuracy.
-As pointed out in previous literature [11], this issue is commonly found in machine learning tasks that demand both strict equivariance and fine-grained generalization performance.
-Although several efforts have been made to alleviate this issue by introducing gated activation mechanisms [8, 7, 12, 13, 14], developing local coordinate strategy [6], and employing discrete spherical channels [11, 15, 16], they still, to varying degrees, compromised either non-linear expressiveness [7, 8, 12, 13, 14] or strict SO(3)-equivariance [11, 15, 16] when processing physical systems.
-Please refer to Supplementary Information A for an in-depth analysis.
+This issue is commonly found in physics-oriented machine learning tasks that demand both strict equivariance and fine-grained generalization performance, as analyzed by Zitnick et al.
+(2022).
+Several recent efforts have been made to alleviate this issue (Zitnick et al., 2022; Passaro & Zitnick, 2023; Wang et al., 2024c; Yin et al., 2024), however, they often compromised strict SO(3)-equivariance.
 To address the equivariance-expressiveness dilemma, we make theoretical and methodological explorations on unifying strict SO(3)-equivariance with strong non-linear expressiveness within the realm of deep representation learning.
 We are inspired by the insight that invariant quantities in transformation often reflect the mathematical nature of physical laws and can induce other quantities with equivariant properties.
-For example, in special relativity [18], the spacetime interval between events is an invariant under Lorentz transformations.
-This invariance is foundational in formulating physical laws that involve covariant quantities.
-Furthermore, from the perspective of deep representation learning, the attribute that invariance is preserved under non-linear operations is a significant advantage, given its compatibility with non-linear expressive capabilities.
-Built upon these insights, we theoretically solve the equivariance-expressiveness dilemma by intensively exploring and making use of the intrinsic relationships between SO(3)-invariant and SO(3)-equivariant quantities and representations: we first dedicate efforts to learning high-quality SO(3)-invariant features with ample non-linear expressiveness, and subsequently, we derive SO(3)-equivariant non-linear representations from these SO(3)-invariant ones.
+For example, in special relativity (Resnick, 1991), the spacetime interval between events is invariant under Lorentz transformations, and is fundamental in formulating physical laws that involve equivariant quantities.
+For another example, the total electronic energy of an atomic system is SO(3)-invariant, while the derived force fields are SO(3)-equivariant.
+We aim to extend the relationship between invariance and equivariance from specific physical quantities to a more general representation learning framework.
+From the perspective of deep representation learning, the attribute that invariance is preserved under non-linear operations is a significant advantage, given its compatibility with non-linear expressive capabilities.
+Built upon these insights, we propose a solution to the equivariance-expressiveness dilemma by intensively exploring and making use of the intrinsic relationships between SO(3)-invariant and SO(3)-equivariant quantities and representations: we first dedicate efforts to learning high-quality SO(3)-invariant features with ample non-linear expressiveness, and subsequently, we derive SO(3)-equivariant non-linear representations and the target quantities from these SO(3)-invariant ones.
 Specifically:
-First, we construct quantities that are theoretically SO(3)-invariant, namely $tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})$, where $\mathbf{Q}$ are the SO(3)-equivariant regression targets, $tr(\cdot)$ signifies the trace operation, and $\dagger$ denotes the conjugate transpose operation.
+First, we propose a theoretical construct of SO(3)-invariant quantities, namely $tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})$, where $tr(\cdot)$ signifies the trace operation, $\dagger$ denotes the conjugate transpose operation, and $\mathbf{Q}$ denotes the SO(3)-equivariant regression targets we aim to predict.
 A significant advantage of these SO(3)-invariant quantities lies in the fact that they are directly derived from the SO(3)-equivariant target labels and can serve as unique supervision labels for the effective learning of informative SO(3)-invariant features that capture the intrinsic symmetry properties of the mathematical structure of $\mathbf{Q}$ without requiring additional labeling resources.
-Second, we propose a gradient-based mechanism to induce SO(3)-equivariant representations for predicting the regression target $\mathbf{Q}$, from high-quality non-linear SO(3)-invariant features learned under the supervision of $tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})$.
-This mechanism can incorporate non-linear expressive capabilities into SO(3)-equivariant representations while preserving their equivariant properties, as we prove.
-We develop our theory into a unified SO(3)-equivariant non-linear representation learning method, and apply it to the computation of electronic-structure Hamiltonian, which occupies a central position in quantum mechanics to induce key physical quantities such as electronic wavefuction, orbital energy, and band structure, bringing deeply understanding of the electrical, optical, magnetic, and transport properties of atomic systems, while posing significant difficulties for machine learning techniques due to its intrinsic symmetry with respect to the SO(3) group as well as the high-dimensional complexity, as shown in Fig.
-1.
-Our method dramatically enhances the state-of-the-art performance on eight benchmark Hamiltonian databases from the renowned DeepH and QH9 benchmark series [6, 8, 19], showing excellent generality to various challenging scenarios such as thermal motions, bilayer twists, scale variations, and new trajectories, with an accuracy improvement of Hamiltonians of up to 40%.
-Moreover, our approach substantially enhances the prediction accuracy of downstream physical quantities of Hamiltonians including occupied orbital energies and electronic wavefunctions, with an accuracy improvement by a maximum of 76% and 17%, respectively.
-Our leading performance comprehensively demonstrate that our method, while satisfying SO(3)-equivariance, possesses excellent expressive power and generalization performance, providing an effective deep learning tool for efficient and accurate electronic-structure calculations of atomic systems.
-Figure 1: Illustration of the high-dimensional complexity of the Hamiltonian.
-In an atomic system, a Hamiltonian matrix is defined between each pair of atoms within each other’s cutoff radius, constituting a large cluster of Hamiltonian matrices for all atom pairs.
-Besides, each Hamiltonian matrix contains multiple basic blocks, with the $\mathbf{Q}=\mathbf{Q}^{l_{p}\otimes l_{q}}$ defined in Section 2 corresponding to a basic block $\mathbf{H}^{l_{p}\otimes l_{q}}$ here.
-Thus, for a Hamiltonian prediction model, the space of the regression target is typically huge.
+Second, we propose a gradient-based mechanism to induce SO(3)-equivariant representations for predicting the regression target $\mathbf{Q}$ in the inference phase from high-quality non-linear SO(3)-invariant features learned under the supervision of $tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})$ in the training phase.
+Taking SO(3)-invariant features as a bridge, this mechanism can incorporate non-linear expressive capabilities into SO(3)-equivariant representations while preserving their equivariant properties, as we prove, laying a solid foundation for accurately inferring complex SO(3)-equivariant targets.
+We develop our theory into a unified SO(3)-equivariant non-linear representation learning method, and apply it to the computation of electronic-structure Hamiltonian, which occupies a central position in quantum mechanics to induce key physical quantities such as electronic wavefuction, orbital energy, and band structure, bringing deep understanding of the electrical, optical, magnetic, and transport properties of atomic systems, while posing significant difficulties for machine learning techniques due to its intrinsic symmetry with respect to the SO(3) group as well as the high-dimensional complexity, as pointed out by Yin et al.
+(2024).
+Our method significantly improves the state-of-the-art performance in Hamiltonian prediction on eight databases from the well-known DeepH and QH9 benchmark series (Li et al., 2022; Gong et al., 2023; Yu et al., 2023a).
+It demonstrates excellent generalization performance to both crystalline and molecular systems, covering challenging scenarios such as thermal motions, bilayer twists, scale variations, and new trajectories.
+Furthermore, as observed from the experiments on the QH9 benchmark series, our approach also substantially enhances the prediction accuracy of downstream physical quantities of Hamiltonian including occupied orbital energy and electronic wavefunction.
+Moreover, our method also demonstrates superior performance in accelerating the convergence of classical DFT by providing predicted Hamiltonians as initialization matrices.
+Our leading performance comprehensively demonstrates that our method, while satisfying SO(3)-equivariance, possesses excellent expressive power and generalization performance, providing an effective deep learning tool for efficient and accurate electronic-structure calculations of atomic systems.
+2 Related Work
+The SO(3)-equivariant representation learning paradigm (Thomas et al., 2018; Geiger & Smidt, 2022) typically developed group theory-based symmetry operators, such as linear scaling, element-wise sum, direct products, direct sums, Clebsch-Gordan decomposition, and equivariant normalization, to encode equivariant features.
+These operators have been used to construct graph neural network architectures for tasks in 3D point cloud analysis (Fuchs et al., 2020), molecular property prediction and dynamic simulation (Musaelian et al., 2023; Liao & Smidt, 2023), as well as Hamiltonian prediction (Schütt et al., 2019; Unke et al., 2021; Gong et al., 2023; Zhong et al., 2023).
+However, as non-linear activation functions may result in the loss of equivariance, they are restricted when applied to SO(3)-equivariant features with degree $l$ greater than one.
+This restriction severely limits the network’s capability to model complex non-linear mappings.
+To alleviate this issue, methods like DeepH-E3 (Gong et al., 2023), QHNet (Yu et al., 2023b) and Equiformer (Liao & Smidt, 2023) introduced a gated activation mechanism that feeds SO(3)-invariant features ($l=0$) into non-linear activation functions and uses these features as linear gating coefficients for SO(3)-equivariant features ($l\geq 1$), aiming to enhance their expressive power while maintaining strict equivariance.
+Nevertheless, multiplying SO(3)-equivariant features with linear coefficients may not fundamentally improve their non-linear expressiveness.
+In order to improve non-linear expressiveness, Zitnick et al.
+(2022) decomposed SO(3)-equivariant features into SO(3)-invariant coefficients of spherical harmonic basis functions.
+These SO(3)-invariant coefficients were processed by non-linear neural networks to enhance expressiveness, with equivariance regained by recombining the updated coefficients with the basis functions.
+In subsequent developments (Passaro & Zitnick, 2023; Liao et al., 2024; Wang et al., 2024b; c), this approach has demonstrated a remarkable capacity to fit complex functions.
+However, as pointed out by existing literature (Zhang et al., 2023), this approach degenerates from continuous to discrete rotational equivariance, losing strict equivariance to continuous rotational transformations due to the decomposition based on inner-product operations with discrete basis functions; Li et al.
+(2022) proposed a local coordinate strategy, projecting rotating global coordinates onto SO(3)-invariant local ones for the non-linear neural network to encode.
+However, this strategy is effective only for global rigid rotations and fails to maintain symmetry under non-rigid perturbations like thermal fluctuations or bilayer twists, as it lacks a neural mechanism to enforce equivariance; Yin et al.
+(2024) proposed a hybrid framework consisting of both group theory-guaranteed SO(3)-equivariant mechanisms and non-linear mechanisms to regress Hamiltonians.
+In this framework, the non-linear mechanisms showed remarkable capability at learning SO(3)-equivariance from the data with the help of the theoretically SO(3)-equivariant mechanisms, and released powerful non-linear expressive capabilities to achieve more numerical accuracy.
+However, the equivariance achieved through data-driven methods does not have strict theoretical guarantee, even with rotational data augmentation.
+In some applications, the demands for symmetry are extremely high.
+Even minor deviations from perfect SO(3)-equivariance can result in incorrect physical results.
+In this paper, we propose a framework that theoretically combines strict SO(3)-equivariance with the non-linear expressiveness of neural networks to resolve the equivariance-expressiveness dilemma.
 3 Problem Formalization
-Physical quantities that are SO(3)-equivariant typically possess their respective equivariant formulas under both matrix-formed direct-product states and vector-formed direct-sum states, which are mathematically equivalent to each other.
-Without loss of generality, let ${\bf{Q}}^{l_{p}\otimes l_{q}}$ denote an SO(3)-equivariant quantity in direct-product state formed by $l_{p}\otimes l_{q}$, i.e., the direct product between degrees $l_{p}$ and $l_{q}$.
+For a background introduction on groups, group representations, equivariance, and invariance, please refer to Appendix A.
+Here, we directly focus on the equivariance of physical quantities under 3D rotational operations that form the SO(3) group.
+Let ${\bf{Q}}^{l_{p}\otimes l_{q}}$ denote an SO(3)-equivariant quantity in direct-product state formed by $l_{p}\otimes l_{q}$, i.e., the direct product between degrees $l_{p}$ and $l_{q}$.
 It obeys the following SO(3)-equivariant law:
-[…]
+${\bf{Q}}({\bf{R}})^{l_{p}\otimes l_{q}}={\bf{D}}^{l_{p}}({\bf{R}})\cdot{\bf{Q}}^{l_{p}\otimes l_{q}}\cdot{({\bf{D}}^{l_{q}}({\bf{R}}))^{\dagger}}$ (1)
 where $\dagger$ denotes the conjugate transpose operation.
-${\bf{R}}\in\mathbb{R}^{3\times 3}$ is the rotational matrix and ${\bf{D}}^{l_{p}}({\bf{R}})\in\mathbb{R}^{(2l_{p}+1)\times(2l_{p}+1)}$, ${\bf{D}}^{l_{q}}({\bf{R}})\in\mathbb{R}^{(2l_{q}+1)\times(2l_{q}+1)}$ are the Wigner-D matrices of degrees $l_{p}$ and $l_{q}$, respectively, ${\bf{Q}}({\bf{R}})^{l_{p}\otimes l_{q}}\in\mathbb{R}^{(2l_{p}+1)\times(2l_{q}+1)}$ denotes the transformed results of ${\bf{Q}}^{l_{p}\otimes l_{q}}\in\mathbb{R}^{(2l_{p}+1)\times(2l_{q}+1)}$ through the rotational operation by ${\bf{R}}$.
-${\bf{Q}}^{l_{p}\otimes l_{q}}$ in the direct-product state can be further decomposed into a series of direct-sum state components with vector forms, i.e., ${\bf{q}}^{l}(|l_{p}-l_{q}|\leq l\leq l_{p}+l_{q})$, which follows SO(3)-equivariant law mathematically equivalent to Eq.
-(1) but with a simpler form:
+${\bf{R}}\in\mathbb{R}^{3\times 3}$ is the rotational matrix, ${\bf{D}}^{l_{p}}({\bf{R}})\in\mathbb{R}^{(2l_{p}+1)\times(2l_{p}+1)}$ and ${\bf{D}}^{l_{q}}({\bf{R}})\in\mathbb{R}^{(2l_{q}+1)\times(2l_{q}+1)}$ are the Wigner-D matrices of degrees $l_{p}$ and $l_{q}$, respectively; ${\bf{Q}}({\bf{R}})^{l_{p}\otimes l_{q}}\in\mathbb{R}^{(2l_{p}+1)\times(2l_{q}+1)}$ denotes the transformed results of ${\bf{Q}}^{l_{p}\otimes l_{q}}\in\mathbb{R}^{(2l_{p}+1)\times(2l_{q}+1)}$ through the rotational operation by ${\bf{R}}$.
+${\bf{Q}}^{l_{p}\otimes l_{q}}$ in the direct-product state can be further decomposed into a series of direct-sum state components, i.e., ${\bf{q}}^{l}(|l_{p}-l_{q}|\leq l\leq l_{p}+l_{q})$, which follows SO(3)-equivariant law mathematically equivalent to Eq.
+1 but with a simpler form:
 ${\bf{q}}({\bf{R}})^{l}={\bf{D}}^{l}({\bf{R}})\cdot{\bf{q}}^{l},|l_{p}-l_{q}|\leq l\leq l_{p}+l_{q}$ (2)
-where ${\bf{q}}^{l}\in\mathbb{R}^{2l+1}$ and ${\bf{q}}({\bf{R}})^{l}\in\mathbb{R}^{2l+1}$ respective denote the components with degree $l$ before and after the rotational operation by ${\bf{R}}$.
-Eq.
-(1) and Eq.
-(2) are mathematical equivalent to each other, meaning that fulfilling one under the corresponding state naturally satisfies the other.
-Due to the simplicity and ease of handling of the direct-sum state form, existing deep learning methods [6, 8, 7] typically first predict $\mathbf{q}^{l}(|l_{p}-l_{q}|\leq l\leq l_{p}+l_{q})$, then convert it to $\mathbf{Q}^{l_{p}\otimes l_{q}}$ through Clebsch-Gordan matrices.
-In the direct-sum state, to obey the equivariant law of Eq.
-(2) for the regression target, the hidden representations of neural networks must also satisfy the same form of equivariance:
+where ${\bf{q}}^{l}\in\mathbb{R}^{2l+1}$ and ${\bf{q}}({\bf{R}})^{l}\in\mathbb{R}^{2l+1}$ respectively denote the components with degree $l$ before and after the rotational operation by ${\bf{R}}$.
+For ease of processing, the internal representations of SO(3)-equivariant neural networks (Gong et al., 2023; Liao & Smidt, 2023) are typically in the direct-sum form.
+To obey the equivariant law of Eq.
+2 for the regression target, these hidden representations must also satisfy the same form of equivariance:
 ${\bf{f}}({\bf{R}})^{(k)l}={\bf{D}}^{l}({\bf{R}})\cdot{\bf{f}}^{(k)l}$ (3)
-where ${\bf{f}}^{(k)l}\in\mathbb{R}^{2l+1}$ and ${\bf{f}}({\bf{R}})^{(k)l}\in\mathbb{R}^{2l+1}$ respectively denote one channel of features with degree $l$ before and after the rotational operation by ${\bf{R}}$, at the $k$ th hidden layer.
-Due to the intrinsic complexity and non-linearity of physic quantities, neural networks on regressing these quantities are supposed to equip with non-linear mappings denoted as $g_{nonlin}(\cdot)$ to fully capture the intrinsic patterns of the physical quantities, which is crucial for precise and generalizable prediction performance.
-Meanwhile, $g_{nonlin}(\cdot)$ must also preserve SO(3)-equivariance, which is expressed as:
-${\bf{f}}({\bf{R}})^{(k+1)l}={\bf{D}}^{l}({\bf{R}})\cdot{\bf{f}}^{(k+1)l},{\bf{f}}^{(k+1)l}=g_{nonlin}({\bf{f}}^{(k)l})$ (4)
+where ${\bf{f}}^{(k)l}\in\mathbb{R}^{2l+1}$ and ${\bf{f}}({\bf{R}})^{(k)l}\in\mathbb{R}^{2l+1}$ respectively denote one channel of hidden features with degree $l$ before and after the rotational operation by ${\bf{R}}$, at the $k$ th hidden layer.
+Due to the intrinsic complexity and non-linearity of physic quantities, neural networks on regressing these quantities are supposed to equip with non-linear mappings to fully capture the intrinsic patterns of the physical quantities, which is crucial for precise and generalizable prediction performance.
+Meanwhile, the non-linear mappings, denoted as $g_{nonlin}(\cdot)$, must also preserve SO(3)-equivariance, which is expressed as:
+${\bf{f}}({\bf{R}})^{(k+1)l}={\bf{D}}^{l}({\bf{R}})\cdot{\bf{f}}^{(k+1)l},\,\,\,\,\text{subject to}\,\,{\bf{f}}^{(k+1)l}=g_{nonlin}({\bf{f}}^{(k)l})$ (4)
 However, directly implementing $g_{nonlin}(\cdot)$ as neural network module with non-linear activation functions, such as $Sigmoid$, $Softmax$ and $SiLU$, may result in the destruction of strict equivariance.
-How to make $g_{nonlin}(\cdot)$ both theoretically SO(3)-equivariant and capable of non-linear expressiveness, is the core problem this paper aims at solving.
+How to make $g_{nonlin}(\cdot)$ both theoretically SO(3)-equivariant and capable of non-linear expressiveness, and effectively apply it to the prediction of SO(3)-equivariant complex physical quantities, i.e., electronic-structure Hamiltonian in this context, is the core problem this paper aims at solving.
 4 Theory
-This section establishes a rigorous theoretical foundation for integrating SO(3)-equivariance with non-linear expressiveness in the deep learning paradigm.
-First, we propose a theoretical construct to derive SO(3)-invariant quantities from the SO(3)-equivariant targets, aiming at facilitating the learning of high-quality SO(3)-invariant features.
-Second, we propose a gradient-based mechanism encoding features maintaining theoretically SO(3)-equivariance while possessing the non-linear expressive capability induced from SO(3)-invariant non-linear features to capture the complex non-linear patterns inherent in physical quantities.
 Theorem 1.
@@
 Theorem 2.
-The non-linear neural mapping defined as the following is SO(3)-equivariant, satisfying the requirements specified in Eq. (4):
-$\mathbf{v}=g_{nonlin}(\mathbf{f})=\frac{\partial z}{\partial\mathbf{f}}\quad\text{s.t.}\quad z=s_{nonlin}(u),u=\text{CGDecomp}(\mathbf{f}\otimes\mathbf{f},0)$ (5)
-where $\mathbf{f}$ is an input SO(3)-equivariant feature with degree $l$ in the direct-sum state, $\otimes$ denotes the direct-product operation of tensors, $\text{CGDecomp}(\cdot,0)$ refers to performing a Clebsch-Gordan decomposition of the tensor and returning the scalar component of degree 0, $s_{nonlin}(\cdot)$ represents non-linear neural modules, and $\mathbf{v}$ is the outputted feature encoded by $g_{nonlin}(\cdot)$.
-Proofs of the two theorems are included in Supplementary Information B.
-Remark 1.
-In Theorem 1, the term $tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})$ is equal to the square of the Frobenius norm of the matrix $\mathbf{Q}$, i.e., $tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})=\|\mathbf{Q}\|_{F}^{2}=\sum_{m,n}|\mathbf{Q}_{mn}|^{2}$.
-The invariance of Frobenius norm under unitary transformations reflects the profound connections between modulus and invariance in physics.
-Remark 2.
-In Theorem 2, $g_{\text{nonlin}}(\cdot)$ can handle equivariant features of any degree and the returned feature $\mathbf{v}$ has the same degree $l$ as the input feature $\mathbf{f}$.
-$\mathbf{v}$ is endowed with non-linear expressiveness with the effect of $s_{nonlin}(\cdot)$, which can be any differentiable non-linear network module, including non-linear activation function as a part.
-The gradient operation $\frac{\partial z}{\partial\mathbf{f}}$ in $g_{\text{nonlin}}(\cdot)$ is used during the forward inference process to promote feature encoding, expanding the usage and significance of gradients within the deep learning paradigm.
-Taking $z$ in Eq.
-(5) as the bridge connecting Theorem 1 and Theorem 2, we present a general method for learning non-linear representations that are equivariant to SO(3), as shown in Fig.
-2.
+The non-linear neural mapping $g_{nonlin}(\cdot)$ defined as the following is SO(3)-equivariant:
+$\mathbf{v}=g_{nonlin}(\mathbf{f})=\frac{\partial z}{\partial\mathbf{f}},\quad\text{subject to}\quad z=s_{nonlin}(u),u=\text{CGDecomp}(\mathbf{f}\otimes\mathbf{f},0)$ (5)
+where $\mathbf{f}$ is an input SO(3)-equivariant feature with degree $l$ in the direct-sum state, $\otimes$ denotes the direct-product operation of tensors, $\text{CGDecomp}(\cdot,0)$ refers to performing a Clebsch-Gordan decomposition of the tensor and returning the scalar component of degree 0, $s_{nonlin}(\cdot)$ represents arbitrary differentiable non-linear neural modules, and $\mathbf{v}$ is the outputted feature encoded by $g_{nonlin}(\cdot)$.
+The proofs of the two theorems above are presented in Appendix B.
+Figure 1: Methodological framework for learning SO(3)-equivariant representations with non-linear expressiveness to regress complex SO(3)-equivariant targets.
+5 Method
+As shown in Fig.
+1, taking $z$ in Eq.
+5 serving as the bridge between Theorem 1 and Theorem 2, we propose a general method for learning non-linear representations that satisfy the SO(3)-equivariance property outlined in Eq.
+4.
 The core of our method can be abstractly referred to as TraceGrad.
-At the label level, it incorporates the SO(3)-invariant trace quantity $tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})$ introduced in Theorem 1 as a supervisory signal for learning SO(3)-invariant non-linear feature $z$.
+At the label level, it incorporates the SO(3)-invariant trace quantity $tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})$ introduced in Theorem 1 as a supervisory signal for learning $z$, i.e., the SO(3)-invariant internal representations of $g_{\text{nonlin}}(\cdot)$ in Theorem 2.
 Given the attribute that invariance is preserved under non-linear operations, $z$ is encoded by $s_{nonlin}(\cdot)$ to obtain non-linear expressive capabilities.
 At the representation level, by taking the gradient of $z$ with respect to $\mathbf{f}$, the non-linear expressiveness of $z$ is transferred to the equivariant feature $\mathbf{v}$, while maintaining strict SO(3)-equivariance as we prove.
-Subsequently, merging $\mathbf{v}$ and $\mathbf{f}$, and applying $g_{\text{nonlin}}(\cdot)$ in a stacked manner can yield rich SO(3)-equivariant non-linear representations to predict the SO(3)-equivariant regression target.
-For details of the method, please refer to Section 6.
-Figure 2: Methodological framework for learning SO(3)-equivariant representations with non-linear expressiveness to regress complex SO(3)-equivariant targets.
-4 Results and Analyses
-We apply our theory and method to the electronic-structure Hamiltonian prediction task, and collect results on eight benchmark databases, i.e., Monolayer Graphene ($MG$), Monolayer MoS2 ($MM$), Bilayer Graphene ($BG$), Bilayer Bismuthene ($BB$), Bilayer Bi2Te3 ($BT$), Bilayer Bi2Se3 ($BS$), QH9-Stable (QS), and QH9-Dynamic (QD).
-The first six databases, consisting of crystal systems with elements like C, Mo, S, Bi, Te and Se, are from the DeepH benchmark series [6, 8]; while the last two databases, composed of molecular systems of C, H, O, N, F, are from the QH9 benchmark series [19] .
-These databases present diverse and complex challenges to the generality of a regression model.
-Regarding $MG$, $MM$, and QD, as their samples are prepared from a 300K environment, the thermal motions lead to complex non-rigid deformations, increasing the difficulty of Hamiltonian prediction.
-For $BG$, $BB$, $BT$, and $BS$, their twisted samples with an interplay of SO(3)-equivariant effects and van der Waals (vdW) force variations bring significant generalization challenges, which are further exacerbated by the absence of any twisted samples in the training sets.
-Besides, $BB$, $BT$, and $BS$ exhibit strong spin-orbit coupling (SOC) effects, which further increase the complexity of Hamiltonian modeling.
-For QS and QD, as we select the most difficult subset split strategies provided in their paper, namely ’ood’ and ’mol’, respectively, yielding differences in atom numbers and trajectories between training and testing samples, challenges in scale and trajectory extrapolation become prominent.
-Detailed information of these databases, including their statistics and visualizations, can be found in the Supplementary Information C.
-We use a comprehensive set of metrics to deeply evaluate the performance of electronic-structure Hamiltonian prediction models.
-On the databases from the DeepH benchmark series [6, 8], we adpot the metrics of Mean Absolute Error (MAE) between predicted and ground truth Hamiltonians introduced by Yin et al.
-[17], including $MAE^{H}_{all}$ for measuring average MAE of all samples and matrix elements, $MAE^{H}_{cha\_s}$ for measuring the MAE of challenging samples where the baseline model performs the worst, $MAE^{H}_{block}$ for measuring the MAE of different basic blocks in the Hamiltonian matrix, and $MAE^{H}_{cha\_b}$ for measuring the MAE on the most challenging Hamiltonian block where the baseline model shows the poorest performance (with the largest $MAE^{H}_{block}$).
-These metrics comprehensively reflect the accuracy performance, covering not only the average accuracy but also the accuracy on difficult samples and challenging blocks of the Hamiltonian matrices.
-On the two databases from the QH9 benchmark series, we adopt the metrics introduced by their original paper [19], including MAE of Hamiltonian matrices, which are further subdivided into $MAE^{H}_{all}$ for measuring average MAE, $MAE^{H}_{diag}$ for measuring MAE of Hamiltonian matrix formed by an atom with itself, and $MAE^{H}_{non\_diag}$ for measuring MAE of Hamiltonian matrices formed by different atoms; as well as the MAE ($MAE^{\epsilon}$) of occupied orbital energies $\epsilon$ induced by the predicted Hamiltonians and compared to the ground truth ones, and the cosine similarity ($Sim(\psi)$) between the electronic wavefunctions $\psi$ induced by the predicted and ground truth Hamiltonians.
-$\epsilon$ and $\psi$ are crucial downstream physical quantities for determining multiple properties of the atomic systems as well as their dynamics, highly reflecting the application values of the Hamiltonian regression model.
-We compare experimental results from two setups: the first one is the experimental results of the baseline SO(3)-equivariant regression model [8, 7] for Hamiltonian prediction, and the second one is the experimental results of extending the architecture and pipeline of the baseline model through the proposed TraceGrad method, which incorporates non-linear expressiveness into the SO(3)-equivariant features of the baseline model with the gradient operations of SO(3)-invariant non-linear features learned by the trace targets.
-We choose DeepH-E3 [8] as the baseline model for databases from the DeepH benchmark series [6, 8]; and we choose QHNet [7] as the baseline model for databases from the QH9 benchmark series [19].
-The results of DeepH-E3 to be compared are taken from the reproduction of Yin et al.
-[17].
-The reproduction, based on DeepH-E3’s newest open-source project [^1] https://github.com/Xiaoxun-Gong/DeepH-E3, showed slightly better results than that in the original publication, and was evaluated more comprehensively across multiple metrics.
-The results of QHNet to be compared are taken from their original paper [19], and for the unification of MAE units, we convert the units of MAE from $10^{-6}$ Hartree ($E_{h}$) in the original paper to meV222 $27211.41E_{h}=$ 27211.4 meV.
-The results of our methods are the average from 10 independent repeated experiments.
-To ensure reproducibility, we used a fixed random seed for all random processes.
-As a result, the standard deviation of the Hamiltonian prediction MAE across repeated experiments does not exceed $0.007$ meV, which is negligible.
-We list the results of DeepH-E3 and DeepH-E3+TraceGrad in Tables 1 and 2 for monolayer structure databases ($MG$, $MM$) and bilayer structure databases ($BG$, $BB$, $BT$, $BS$) from the DeepH benchmark series, reporting the values of $MAE^{H}_{all}$, $MAE^{H}_{cha\_s}$, and $MAE^{H}_{cha\_b}$.
-Regarding the metric of $MAE^{H}_{block}$ for every Hamiltonian block, due to its large data volume, we just present its values for representative datasets here in Fig.
-3, and attach its values for all databases from the DeepH series in Supplementary Information E.
-From these results, we could find that the proposed TraceGrad method dramatically enhances the accuracy performance of the baseline method DeepH-E3, both on average and for challenging samples or blocks, both on the non-twisted samples and the twisted samples.
-Specifically, on the corresponding datasets, TraceGrad lowers down the $MAE^{H}_{all}$ and $MAE^{H}_{cha}$ of DeepH-E3 with relative ratios of up to $34\%$ and $35\%$, respectively, and achieving better results in the vast majority of basic blocks.
-Particularly, for the blocks where DeepH-E3 perform the worst, TraceGrad reduces the $MAE$ ($MAE^{H}_{cha\_b}$) by a maximum of $48\%$.
-The leading performance on the $MG$ and $MM$ databases prepared at 300K demonstrates the robustness of our method against thermal motion.
-The high accuracy on the $BB$, $BT$, and $BS$ databases, which have strong SOC effects, indicates our method’s strong capability to model such effects.
-The excellent performance on the $BG^{t}$, $BB^{t}$, $BT^{t}$, and $BS^{t}$ subsets showcases the method’s superior generalization to twisted structures, which are not present in the training data.
-The outstanding performance on such samples highlights the good potential for studying twist-related phenomena, given interlayer twists are hot research topics because they may bring new electrical and transport properties [20, 21, 22].
-Additionally, the $BG^{t}$ and $BB^{t}$ testing subsets contain significantly larger atomic systems compared to the training set (see Supplementary Information C for statistics of their sizes), yet our method still excels on these subsets for all metrics, demonstrating its good scalability on the sizes of atomic systems it handles.
-We list the results of QHNet and QHNet+TraceGrad in Tables 3 and 4, for the $QS$ and $QD$ databases, respectively, reporting the values of $MAE^{H}_{all}$, $MAE^{H}_{diag}$, $MAE^{H}_{non\_diag}$, $MAE^{\epsilon}$, and $Sim(\psi)$.
-From these results, we can observe that the proposed TraceGrad method significantly enhances the accuracy performance of the baseline method QHNet across all of these metrics on both of the two databases.
-Specifically, TraceGrad lowers down the $MAE^{H}_{all}$, $MAE^{H}_{non\_diag}$, $MAE^{H}_{diag}$, $MAE^{\epsilon}$ and $Sim(\psi)$ of QHNet with relative ratios of up to 40%, 39%, 40%, 76%, and 17%, on the corresponding databases.
-The significant improvements on the $QS$ database, partitioned using the ’ood’ split strategy [19] without scales overlapping among the training, validation, and testing sets, once again demonstrate the method’s strong generalization capabilities across different scales.
-Meanwhile, the improvements on the $QD$ database under the ’mol’ strategy [19] partitioning the training, validation, and testing sets with samples from totally different thermal motion trajectories, highlight the method’s robustness in generalizing to new thermal motion sequences.
-Furthermore, the significant promotion on the prediction accuracy of $\epsilon$, i.e., occupied orbital energies which are crucial for determining electronic properties such as optical characteristics and conductivity in atomic systems, and $\psi$, i.e., the electronic wavefunctions essential for understanding electron distribution and interactions, underscore good application potential of our method for tasks like material design and quantum computing.
-The experimental results above indicate that the proposed representation learning framework, while maintaining strict SO(3)-equivariance, significantly enhances the expressiveness of equivariant representations, thus excellently capturing the complexity of physical systems and substantially breakthroughs the SOTA accuracy on the eight benchmark databases.
-The consistent improvements brought by our framework to different baseline models (DeepH-E3 or QHNet) and in various challenging scenarios, such as thermal motions, bilayer twists, scale variations, and new trajectories, fully demonstrate the universality and significance of our proposed theory and method.
-Table 1: Experimental results measured by the $MAE^{H}_{all}$, $MAE^{H}_{cha\_s}$, and $MAE^{H}_{cha\_b}$ metrics (meV) on the Monolayer Graphene ($MG$) and Monolayer MoS2 ($MM$) databases.
-$\downarrow$ means lower values of the metrics correspond to better accuracy.
+Subsequently, merging $\mathbf{v}$ and $\mathbf{f}$, and applying $g_{\text{nonlin}}(\cdot)$ in a stacked manner can yield rich SO(3)-equivariant non-linear representations for inferring the SO(3)-equivariant regression target.
+5.1 Encoding and Decoding Framework
+Our encoding framework corresponds to a total of $K$ encoding modules, which are sequentially connected to form a deep encoding framework.
+For the $k$-th module ($1\leq k\leq K$), it first introduces an SO(3)-equivariant backbone encoder, e.g., the encoders from DeepH-E3 (Gong et al., 2023) or QHNet (Yu et al., 2023b) which is composed of recently developed equivariant operators (Thomas et al., 2018; Geiger & Smidt, 2022) like linear scaling, direct products, direct sums, Clebsch-Gordan decomposition, gated activation, equivariant normalization, and etc., to encode the physical system’s initial representations, such as spherical harmonics (Schrodinger, 1926), or representations passed from the previous neural layers, as equivariant feature $\textbf{f}^{(k)}$ in the current hidden layer.
+Next, we construct the feature $\textbf{v}^{(k)}=g_{\text{nonlin}}(\textbf{f}^{(k)})$, to achieve sufficient non-linear expressiveness while maintaining SO(3)-equivariance, where $g_{\text{nonlin}}(\cdot)$, $s_{nonlin}(\cdot)$ are the non-linear functions defined in Eq.
+5.
+The function $s_{nonlin}(\cdot)$ can be implemented as any differentiable non-linear neural network module, such as feed-forward layers with non-linear activation functions like $SiLU$ and normalization operations like $Layernorm$.
+For an expressive representation of a complex physical system, $\textbf{f}^{(k)}$ is usually not a feature of single degree but a direct-sum concatenation of series of components $\{\textbf{f}^{(k)l_{1}},\textbf{f}^{(k)l_{2}},\textbf{f}^{(k)l_{3}},...\}$ at multiple degrees, i.e., $L^{(k)}=\{l_{1},l_{2},l_{3},...\}$, where some components share the same degree while others differ.
+In this case, it becomes necessary to extend the decomposition operator, i.e., $\text{CGDecomp}(\cdot)$ in Eq.
+5, to accommodate various components of degrees.
+Moreover, in the context of neural networks, introducing learnable parameters $W$ and more feature channels $C$ may improve the model capacity.
+Based on these considerations, when constructing the encoding module, the $\text{CGDecomp}(\cdot)$ operation in Eq.
+5 can be expanded as $\text{CGDecomp}_{ext}(\cdot)$, and its outputs can be expanded as $\textbf{u}^{(k)}$, as shown in Eq.
+6:
+$\textbf{u}^{(k)}$ $=[u_{1}^{(k)},...,u_{c}^{(k)},...,u_{C}^{(k)}],$ (6)
+$u_{c}^{(k)}$ $=\text{CGDecomp}_{\text{ext}}(\textbf{f}^{(k)}\otimes\textbf{f}^{(k)},0,W)=\sum_{l_{i},l_{j}\in L^{(k)},l_{i}=l_{j}}W^{c}_{ij}\cdot\text{CGDecomp}(\textbf{f}^{(k)l_{i}}\otimes\textbf{f}^{(k)l_{j}},0)$
+where $W^{c}_{ij}$ represents learnable parameters, $u_{c}^{(k)}$ is the $c$ th channel of $\mathbf{u}^{(k)}$.
+$\text{CGDecomp}_{\text{ext}}(\cdot)$ adds learnable parameters to $\text{CGDecomp}(\cdot)$ and expands its output from a single channel to multiple channels.
+To further enhance the model capacity, we also expand the output of $s_{\text{nonlin}}(\cdot)$ to multiple channels as follows: $\mathbf{z}^{(k)}=[z_{1}^{(k)},\ldots,z_{c}^{(k)},\ldots,z_{C}^{(k)}]$.
+We define $\mathbf{v}_{c}^{(k)}=\frac{\partial z_{c}^{(k)}}{\partial\mathbf{f}^{(k)}}$, and construct the new features $\mathbf{v}^{(k)}$ by $\mathbf{v}^{(k)}=\sum_{c}\mathbf{v}_{c}^{(k)}$.
+It is evident that these extensions maintain the SO(3)-invariance of $\mathbf{u}^{(k)}$ and $\mathbf{z}^{(k)}$, as well as the SO(3)-equivariance of $\mathbf{v}^{(k)}$.
+$\mathbf{v}^{(k)}$ is combined with $\mathbf{f}^{(k)}$ in a residual manner like $\textbf{o}^{(k)}=\textbf{f}^{(k)}+\textbf{v}^{(k)}$ as the output of the $k$ th encoding module.
+[…]
+5.2 Training
+The training loss function is shown as the following:
+$\begin{split}\min_{\theta}loss&=loss_{Q}+\mu(loss_{Q},loss_{T})\cdot loss_{T},\\ loss_{Q}&=Error(\widehat{\mathbf{Q}},\mathbf{Q}^{*}),\,\,\,\,loss_{T}=Error(\widehat{\mathbf{T}},\mathbf{T}^{*})\end{split}$ (7)
+where $\theta$ denotes all of the learnable parameters of our framework, $\widehat{\mathbf{Q}}$, $\widehat{\mathbf{T}}$ and $\mathbf{Q}^{*}$, $\mathbf{T}^{*}$ respectively denote the predictions and labels of $\mathbf{Q}$ and $\mathbf{T}$.
+In order to prevent the numerical disparity between the two loss terms and stabilize the training for both of the SO(3)-equivariant and SO(3)-invariant branches, we apply $\mu(loss_{Q},loss_{T})$, i.e., a coefficient to regularize the relative scale of the two loss terms:
+$\mu(loss_{Q},loss_{T})=\lambda\cdot No\_Grad(\frac{loss_{Q}}{loss_{T}})$ (8)
+where $No\_Grad(\cdot)$ denotes gradient discarding when calculating such coefficient, as this coefficient is only used for adjusting the relative scale between the two loss terms and should not itself be a source of training gradients, otherwise it would counteract the gradients from $loss_{T}$ in Eq.
+7.
+All of the encoding and decoding modules are trained jointly by Eq.
+7.
+For the implementation details of our method, including the specific design of the network modules, parameter settings, and training specifics, please refer to Appendix D.
+6 Experiments
+6.1 Experimental Conditions
+Table 1: Experimental results measured by the $MAE^{H}_{all}$, $MAE^{H}_{cha\_s}$ and $MAE^{H}_{cha\_b}$ metrics (meV) on the Monolayer Graphene ($MG$), Monolayer MoS2 ($MM$), Bilayer Graphene ($BG$), Bilayer Bismuthene ($BB$), Bilayer Bi2Te3 ($BT$) and Bilayer Bi2Se3 ($BS$) databases, where the superscripts $nt$ and $t$ respectively denote the non-twisted and twisted subsets.
 Methods $MG$ $MM$
@@
 DeepH-E3+TraceGrad 0.175 0.257 0.228 0.285 0.412 0.808
-Table 2: Experimental results measured by the $MAE^{H}_{all}$, $MAE^{H}_{cha\_s}$ and $MAE^{H}_{cha\_b}$ metrics (meV) on the Bilayer Graphene ($BG$), Bilayer Bismuthene ($BB$), Bilayer Bi2Te3 ($BT$) and Bilayer Bi2Se3 ($BS$) databases, where the superscripts $nt$ and $t$ respectively denote the non-twisted and twisted subsets.
-$\downarrow$ means lower values of the metrics correspond to better accuracy.
 Methods $BG^{nt}$ $BG^{t}$
-$MAE$ ()($\downarrow$)( ↓ )
-$MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$ $MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$
 DeepH-E3 (Baseline) 0.389 0.453 0.644 0.264 0.429 0.609
@@
 Methods $BB^{nt}$ $BB^{t}$
-$MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$ $MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$
 DeepH-E3 (Baseline) 0.274 0.304 1.042 0.468 0.602 2.399
@@
 Methods $BT^{nt}$ $BT^{t}$
-$MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$ $MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$
 DeepH-E3 (Baseline) 0.447 0.480 1.387 0.831 0.850 4.572
@@
 Methods $BS^{nt}$ $BS^{t}$
-$MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$ $MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$
 DeepH-E3 (Baseline) 0.397 0.424 0.867 0.370 0.390 0.875
 DeepH-E3+TraceGrad 0.300 0.332 0.644 0.291 0.302 0.674
-Figure 3: Visualization of $MAE^{H}_{block}$ on each basic block of the Hamiltonian matrices for representative datasets, where the superscripts $nt$ and $t$ respectively denote the non-twisted and twisted subsets.
-Lower value of $MAE^{H}_{block}$ corresponds to better accuracy.
-The $MAE^{H}_{block}$ for other datasets from DeepH series can be found in Supplementary Information E.
-Table 3: Experimental results measured by the $MAE^{H}_{all}$, $MAE^{H}_{diag}$, $MAE^{H}_{non\_diag}$, $MAE^{\epsilon}$, and $Sim(\psi)$ metrics on the QH9-Stable (QS) database using the ’ood’ split strategy [19].
-$\downarrow$ means lower values of the metrics correspond to better accuracy, while $\uparrow$ means higher value of the metric corresponds to better performance.
+We apply our theory and method to the electronic-structure Hamiltonian prediction task, and collect results on eight benchmark databases, i.e., Monolayer Graphene ($MG$), Monolayer MoS2 ($MM$), Bilayer Graphene ($BG$), Bilayer Bismuthene ($BB$), Bilayer Bi2Te3 ($BT$), Bilayer Bi2Se3 ($BS$), QH9-stable (QS), and QH9-dynamic (QD).
+The first six databases, consisting of periodic crystalline systems with elements like C, Mo, S, Bi, Te and Se, are from the DeepH benchmark series (Li et al., 2022; Gong et al., 2023).
+The last two databases are from the QH9 benchmark series (Yu et al., 2023a), composed of molecular systems with elements like C, H, O, N and F.
+These databases present diverse and complex challenges to the generality of a regression model.
+Regarding $MG$, $MM$, and QD, as their samples are prepared from an temperature environment at three-hundred Kelvin, the thermal motions lead to complex non-rigid deformations, increasing the difficulty of Hamiltonian prediction.
+For $BG$, $BB$, $BT$, and $BS$, the twisted structures, with an interplay of SO(3)-equivariant effects and van der Waals (vdW) force variations bring significant generalization challenges, which are further exacerbated by the absence of any twisted samples in the training sets.
+Besides, $BB$, $BT$, and $BS$ exhibit strong spin-orbit coupling (SOC) effects, which further increase the complexity of Hamiltonian modeling.
+For the $QS$ database, the ’ood’ strategy from the official settings is used to split the training, validation, and testing sets, ensuring that the atom number of samples do not overlap across the three subsets.
+For the $QD$ database, the ’mol’ strategy provided by Yu et al.
+(2023a) is applied to split the training, validation, and testing sets, ensuring that there are no thermal motion samples from the same temporal trajectory across the three subsets.
+The ’mol’ and ’ood’ strategies aim to assess the regression model’s extrapolation capability with respect to the number of atoms as well as the temporal trajectories, respectively.
+Detailed statistic information of these databases can be found in the Appendix C.
+Implementation details of our method for experiments on these databases are presented in Appendix D.
+Table 2: Experimental results measured by the $MAE^{H}_{all}$, $MAE^{H}_{diag}$, $MAE^{H}_{non\_diag}$, $MAE^{\epsilon}$, and $Sim(\psi)$ metrics on the QH9-stable (QS) and QH9-dynamic (QD) databases respectively using ’ood’ and ’mol’ split strategies (Yu et al., 2023a).
+$\downarrow$ means lower values correspond to better accuracy, while $\uparrow$ means higher values correspond to better performance.
 The units of MAE metrics are meV, while $Sim(\psi)$ is the cosine similarity which is dimensionless.
-Methods MAE ()($\downarrow$)( ↓ ) $Sim(\psi)$ ()($\uparrow$)( ↑ )
+Methods QS
+MAE ()($\downarrow$)( ↓ ) $Sim(\psi)$ ()($\uparrow$)( ↑ )
 $MAE^{H}_{all}$ $MAE^{H}_{diag}$ $MAE^{H}_{non\_diag}$ $MAE^{\epsilon}$
-QHNet 1.962 3.040 1.902 17.528 0.937
+QHNet (Baseline) 1.962 3.040 1.902 17.528 0.937
 QHNet+TraceGrad 1.191 2.125 1.139 8.579 0.948
-Table 4: Experimental results measured by the $MAE^{H}_{all}$, $MAE^{H}_{diag}$, $MAE^{H}_{non\_diag}$, $MAE^{\epsilon}$, and $Sim(\psi)$ metrics on the QH9-Dynamic (QD) database using the ’mol’ split strategy [19].
-$\downarrow$ means lower values of the metrics correspond to better accuracy, while $\uparrow$ means higher value of the metric corresponds to better performance.
-The units of MAE metrics are meV, while $Sim(\psi)$ is the cosine similarity which is dimensionless.
-Methods MAE ()($\downarrow$)( ↓ ) $Sim(\psi)$ ()($\uparrow$)( ↑ )
-$MAE^{H}_{all}$ $MAE^{H}_{diag}$ $MAE^{H}_{non\_diag}$ $MAE^{\epsilon}$
-QHNet 4.733 11.347 4.182 264.483 0.792
+Methods QD
+QHNet (Baseline) 4.733 11.347 4.182 264.483 0.792
 QHNet+TraceGrad 2.819 6.844 2.497 63.375 0.927
-We present experimental results of ablation study in Supplementary Information F.
+We use a comprehensive set of metrics to deeply evaluate the performance of deep learning electronic-structure Hamiltonian prediction models.
+The metrics can be broadly categorized into two major aspects:
+• Metrics on accuracy performance.
+On the databases from the DeepH benchmark series (Li et al., 2022; Gong et al., 2023), we follow Yin et al.
+(2024) to adopt a set of Mean Absolute Error (MAE) metrics between predicted and ground truth Hamiltonians, including $MAE^{H}_{all}$ for measuring average MAE of all samples and matrix elements, $MAE^{H}_{cha\_s}$ for measuring the MAE of challenging samples where the baseline model performs the worst, $MAE^{H}_{block}$ for measuring the MAE of different basic blocks in the Hamiltonian matrix, and $MAE^{H}_{cha\_b}$ for measuring the MAE on the most challenging Hamiltonian block where the baseline model shows the poorest performance (with the largest $MAE^{H}_{block}$).
+These metrics comprehensively reflect the accuracy performance, covering not only the average accuracy but also the accuracy on difficult samples and challenging blocks of the Hamiltonian matrices.
+On the two databases from the QH9 benchmark series, we adopt the metrics introduced by their original paper (Yu et al., 2023a), including MAE of Hamiltonian matrices, which are further subdivided into $MAE^{H}_{all}$ for measuring average MAE, $MAE^{H}_{diag}$ for measuring MAE of Hamiltonian matrix formed by an atom with itself, and $MAE^{H}_{non\_diag}$ for measuring MAE of Hamiltonian matrices formed by different atoms; as well as the MAE ($MAE^{\epsilon}$) of occupied orbital energies $\epsilon$ induced by the predicted Hamiltonians and compared to the ground truth ones, and the cosine similarity ($Sim(\psi)$) between the electronic wavefunctions $\psi$ induced by the predicted and ground truth Hamiltonians.
+$\epsilon$ and $\psi$ are crucial downstream physical quantities for determining multiple properties of the atomic systems as well as their dynamics, highly reflecting the application values of the Hamiltonian regression model.
+• Metrics on acceleration ratios for the convergence of traditional DFT algorithms.
+Despite the increasing ability of deep learning models to independently handle more electronic-structure computation tasks, there are still applications with extremely high numerical precision requirements and very low tolerance for error, where traditional DFT algorithms must perform the final calculations.
+In such cases, the predictions from deep models can be used as initial matrices to accelerate the convergence of traditional DFT algorithms.
+We evaluate the acceleration ratios brought by the proposed method for the convergence of classical DFT algorithms implemented by PySCF (Sun et al., 2018).
+Specifically, we adopt the two metrics defined in Yu et al.
+(2023a): the achieved ratio calculates the number of DFT optimization steps taken when initializing with the Hamiltonian matrices predicted by the deep model compared to using initial guess methods like minao and 1e.
+The error-level ratio measures the number of DFT optimization steps required, starting from random initialization, to reach the same error level as the deep model’s predictions, relative to the total number of steps in the DFT process.
+Table 3: The acceleration ratios of QHNet and QHNet+TraceGrad for DFT calculation.
+Both models are evaluated on a set of $50$ molecules chosen by Yu et al.
+(2023a), with the mean and standard deviation of the metrics across these samples reported.
+$\downarrow$ means lower values correspond to better accuracy, while $\uparrow$ means higher values correspond to better performance.
+Methods Training DFT Metric Ratio
+databases initialization
+QHNet (Baseline) QS 1e Achieved ratio $\downarrow$ $0.400\pm 0.030$
+Error-level ratio $\uparrow$ $0.620\pm 0.037$
+minao Achieved ratio $\downarrow$ $0.715\pm 0.033$
+Error-level ratio $\uparrow$ $0.406\pm 0.021$
+QD 1e Achieved ratio $\downarrow$ $0.512\pm 0.138$
+Error-level ratio $\uparrow$ $0.622\pm 0.048$
+minao Achieved ratio $\downarrow$ $0.882\pm 0.217$
+Error-level ratio $\uparrow$ $0.406\pm 0.066$
+QHNet+TraceGrad QS 1e Achieved ratio $\downarrow$ $\textbf{0.345}\pm 0.038$
+Error-level ratio $\uparrow$ $\textbf{0.685}\pm 0.037$
+minao Achieved ratio $\downarrow$ $\textbf{0.647}\pm 0.061$
+Error-level ratio $\uparrow$ $\textbf{0.466}\pm 0.035$
+QD 1e Achieved ratio $\downarrow$ $\textbf{0.440}\pm 0.101$
+Error-level ratio $\uparrow$ $\textbf{0.645}\pm 0.046$
+minao Achieved ratio $\downarrow$ $\textbf{0.761}\pm 0.167$
+Error-level ratio $\uparrow$ $\textbf{0.435}\pm 0.052$
+6.2 Results and Analysis
+We compare experimental results from two setups: the first one is the experimental results of the baseline SO(3)-equivariant regression model (Gong et al., 2023; Yu et al., 2023b) for Hamiltonian prediction, and the second one is the experimental results of extending the architecture and pipeline of the baseline model through the proposed TraceGrad method, which incorporates non-linear expressiveness into the SO(3)-equivariant features of the baseline model with the gradient operations of SO(3)-invariant non-linear features learned under the supervision of the trace targets.
+We choose DeepH-E3 (Gong et al., 2023) as the baseline model for databases from the DeepH benchmark series (Li et al., 2022; Gong et al., 2023); and we choose QHNet (Yu et al., 2023b) as the baseline model for databases from the QH9 benchmark series (Yu et al., 2023a).
+They are the respective state-of-the-art (SOTA) methods with strict SO(3)-equivariance on the corresponding databases.
+We list the results of DeepH-E3 and DeepH-E3+TraceGrad in Tables 1 for databases from the DeepH benchmark series, reporting the values of $MAE^{H}_{all}$, $MAE^{H}_{cha\_s}$, and $MAE^{H}_{cha\_b}$.
+The results of DeepH-E3 to be compared are copied from Yin et al.
+(2024).
+The results of DeepH-E3+TraceGrad are the average from 10 independent repeated experiments.
+Regarding the metric of $MAE^{H}_{block}$ for every Hamiltonian block, due to its large data volume, we just present its values for all databases from the DeepH series in Appendix 4.
+We use the same fixed random seed as adopted by DeepH-E3 for all random processes in experiments on these six databases.
+As a result, the standard deviation of the Hamiltonian prediction MAE across repeated experiments does not exceed $0.007$ meV for each of the six databases and is negligible.
+From results presented in Tables 1, we could find that the proposed TraceGrad method dramatically enhances the accuracy performance of the baseline method DeepH-E3, both on average and for challenging samples and blocks, both on the non-twisted samples and the twisted samples.
+Specifically, on the corresponding datasets, TraceGrad lowers down the $MAE^{H}_{all}$ and $MAE^{H}_{cha}$ of DeepH-E3 with relative ratios of up to $34\%$ and $35\%$, respectively.
+Furthermore, from the results included in Appendix 4, TraceGrad significantly improves the performance for the vast majority of basic blocks.
+Particularly, for the blocks where DeepH-E3 perform the worst, TraceGrad reduces the $MAE$ ($MAE^{H}_{cha\_b}$) by a maximum of $48\%$.
+The leading performance on the $MG$ and $MM$ databases prepared at three-hundred Kelvin temperature demonstrates the robustness of our method against thermal motion.
+The high accuracy on the $BB$, $BT$, and $BS$ databases, which have strong SOC effects, indicates our method’s strong capability to model such effects.
+The excellent performance on the $BG^{t}$, $BB^{t}$, $BT^{t}$, and $BS^{t}$ subsets showcases the method’s superior generalization to twisted structures, which are not present in the training data.
+The outstanding performance on such samples highlights the good potential for studying twist-related phenomena, a hot research topic that may bring new electrical and transport properties (Cao et al., 2018; Wang et al., 2024a; He et al., 2024).
+Additionally, the $BG^{t}$, $BB^{t}$, $BT^{t}$, and $BS^{t}$ subsets contain significantly larger unit cells compared to the training set (see Appendix C for statistics of their sizes), yet our method still excels on these subsets as measured by the multiple MAE metrics, demonstrating its good scalability on the sizes of atomic systems it handles.
+In Table 2, we present the results of QHNet and QHNet+TraceGrad under the metrics of $MAE^{H}_{all}$, $MAE^{H}_{diag}$, $MAE^{H}_{non\_diag}$, $MAE^{\epsilon}$, and $Sim(\psi)$ for the $QS$ and $QD$ databases.
+The results of QHNet to be compared are taken from their original paper (Yu et al., 2023a), and for the unification of MAE units, we convert the units of MAE from $10^{-6}$ Hartree ($E_{h}$) in the original paper to meV111 $27211.41E_{h}=$ 27211.4 meV.
+The results of QHNet+TraceGrad are the average from 10 independent repeated experiments.
+To ensure reproducibility, we use the same fixed random seeds as employed in QHNet for all random processes in the experiments on the $QS$ and $QD$ databases.
+As a result, the standard deviation of the Hamiltonian prediction MAE across repeated experiments is no greater than $0.009$ meV for both $QS$ and $QD$ and is also negligible.
+The results presented in Table 2 demonstrate that the proposed TraceGrad method significantly enhances the accuracy of the baseline QHNet model across all metrics on the $QS$ and $QD$ databases.
+Specifically, TraceGrad reduces $MAE^{H}_{all}$, $MAE^{H}_{non\_diag}$, $MAE^{H}_{diag}$, $MAE^{\epsilon}$, and $Sim(\psi)$ of QHNet with relative reductions of up to 40%, 39%, 40%, 76%, and 17%, respectively, on the corresponding databases.
+The significant accuracy improvements on the $QS$ database, partitioned using the ’ood’ split strategy (Yu et al., 2023a) without scale overlapping among the training, validation, and testing sets, once again demonstrate the method’s strong generalization capabilities across different scales of atomic systems.
+Meanwhile, performance on the $QD$ database under the ’mol’ strategy (Yu et al., 2023a), which partitions the training, validation, and testing sets with samples from completely different thermal motion trajectories, highlight our method’s robustness in generalizing to new thermal motion sequences.
+Furthermore, the substantial improvement in the prediction accuracy of $\epsilon$, i.e., occupied orbital energies crucial for determining electronic properties such as optical characteristics and conductivity in atomic systems, and $\psi$, i.e., the electronic wavefunctions essential for understanding electron distribution and interactions, underscores the potential values of our method for applications like material design, molecular pharmacology, and quantum computing.
+We also compare the acceleration ratios brought by the QHNet model and the QHNet+TraceGrad model for the convergence of classical DFT algorithms implemented by PySCF (Sun et al., 2018), with the experimental results reported in Table 3, using the achieved ratio and error-level ratio metrics introduced in Section 6.1.
+As observed from this table, TraceGrad also brings significant improvements to the baseline model QHNet, notably reducing the achieved ratio and enhancing the error-level ratio, indicating its substantial potential for accelerating traditional DFT algorithms.
+Due to the page limits, we present three groups of supplementary experimental results in the Appendices.
+The first group, included in Appendix H, is the results of ablation study.
 These fine-grained experiments assess the individual and combined contributions of the two core mechanisms of the proposed TraceGrad method on the six databases from the DeepH series.
@@
 Moreover, their combination provides even better performance.
-This is because, on one hand, with the gradient-based induction mechanism as a bridge, the SO(3)-invariant features learned from the trace label can be transformed into the SO(3)-equivariant representations during inference; on the other hand, with trace label, the SO(3)-invariant network branch has a strong supervisory signal, enabling it to learn the intrinsic symmetry and complexity of the regression targets, enhancing the quality of SO(3)-invariant features and ultimately benefits the encoding of SO(3)-equivariant features.
-The value of such complementarity has been fully demonstrated in the experimental results.
-5 Future Work
-In future research, various extensions are conceivable across theoretical, methodological, and application fields:
-First, from a theoretical perspective, we can generalize SO(3) group symmetry to a broader range and more complex groups.
-This will enable deep learning frameworks to accommodate richer mathematical structures.
-Second, from a methodological perspective, we plan to combine our framework with advanced Transformer [23] and KAN [24] architectures to continuously enhance the expressive capabilities of equivariant features.
-Third, from the application perspective, the current methodology extends beyond predicting electronic-structure Hamiltonians and is capable of generally predicting multiple physical quantities and properties that require equivariance.
-Furthermore, the methodology is not confined to physical research alone and may find applications in areas such as robotics, autonomous vehicles, and motion tracking systems.
-For instance, in the vision tasks of autonomous vehicles [25], where there is a high demand for robustness against coordinate transformations of cameras, the mainstream approach has been to achieve approximate SO(3)-equivariance through data augmentation.
-However, this does not guarantee absolute reliability and safety.
-Our work holds great potential to construct deep models with high generalization performance under the premise of strict SO(3)-equivariance, potentially revolutionizing techniques in such tasks.
-Looking forward to future endeavors, the foundational and pivotal role of this work is clear.
-6 Method
-6.1 Encoding and Decoding Framework
-As shown in Fig.
-2, our encoding framework corresponds to a total of $K$ encoding modules, which are sequentially connected to form a cascading stack.
-For the $k$-th module ($1\leq k\leq K$), it first introduces an SO(3)-equivariant base encoder, e.g., the encoder from DeepH-E3 [8] or QHNet [7] which is composed of recently developed equivariant operators [26, 27] like linear scaling, direct products, direct sums, Clebsch-Gordan decomposition, gated activation, equivariant normalization, and etc., to encode the physical system’s initial representations, such as spherical harmonics [28], or representations passed from the previous neural layers, as equivariant feature $\textbf{f}^{(k)}$ in the current hidden layer.
-Next, we construct the feature $\textbf{v}^{(k)}=g_{\text{nonlin}}(\textbf{f}^{(k)})$, to achieve sufficient non-linear expressiveness while maintaining SO(3)-equivariance, where $g_{\text{nonlin}}(\cdot)$ is the function defined in Eq.
-(5).
-$s_{nonlin}(\cdot)$, i.e., $g_{\text{nonlin}}(\cdot)$’s built-in non-linear mechanism, which can be any differentiable non-linear neural module, is implemented here as a feed-forward module with a hidden layer size of $1024$, using Layer Normalization for stabilization and regularization, and employing SiLU as the non-linear mapping.
-To facilitate learning and optimization, we connect the input and output of $g_{\text{nonlin}}(\cdot)$ in a residual manner [29], serving as the output of the current encoding module, i.e., $\textbf{o}^{(k)}=\textbf{f}^{(k)}+\textbf{v}^{(k)}$.
-$\textbf{o}^{(k)}$ is then fed into the next encoder module to yield $\textbf{f}^{(k+1)}$, $\textbf{o}^{(k+1)}$, …., until the last encoder module, which is followed by the decoder modules.
-For an expressive feature encoder handling a complex physical system, $\textbf{f}^{(k)}$ is usually not a feature of single degree but a direct-sum concatenation of series of components $\textbf{f}^{(k)l_{1}}$, $\textbf{f}^{(k)l_{2}}$, …, $\textbf{f}^{(k)l_{k}}$ at multiple degrees, i.e., $l_{1}$, $l_{2}$, …, $l_{k}$, where some components share the same degree while others differ.
-In this case, it becomes necessary to extend the decomposition operator, i.e., $\text{CGDecomp}(\cdot)$ in Eq.
-(5), to accommodate various components of degrees.
-Moreover, in the context of neural networks, introducing learnable parameters $W$ may improve the model capacity.
-Based on these considerations, when constructing the encoding module, the $\text{CGDecomp}(\cdot)$ operation in Eq.
-(5) can be expanded as $\text{CGDecomp}_{ext}(\cdot)$:
-$u^{(k)}=\text{CGDecomp}_{\text{ext}}(\textbf{f}^{(k)}\otimes\textbf{f}^{(k)},0,W)=\sum_{1\leq i,j\leq k,l_{i}=l_{j}}W_{ij}\cdot\text{CGDecomp}(\textbf{f}^{(k)l_{i}}\otimes\textbf{f}^{(k)l_{j}},0)$ (6)
-where $W$ represents learnable parameters.
-This expansion does not change the equivariance or invariance of the features.
-It is obvious that, the output ($u^{(k)}$) of $\text{CGDecomp}_{\text{ext}}(\cdot)$ is also SO(3)-invariant, just like the original $\text{CGDecomp}(\cdot)$.
-Thus, even with such an expansion, the feature $z^{(k)}=s_{\text{nonlin}}(u^{(k)})$ remains SO(3)-invariant and $\textbf{v}^{(k)}=\frac{\partial z^{(k)}}{\partial\mathbf{f}^{(k)}}$ remain SO(3)-equivariant.
-[…]
-Since complex non-linear transformations have already been incorporated in the encoding framework, the decoders can be simple: the SO(3)-equivariant decoder is implemented using an SO(3)-equivariant linear mapping [27], and the SO(3)-invariant decoder is implemented as a shallow feed-forward module.
-6.2 Training and Inference
-The training loss function is shown as the following:
-$\begin{split}\min_{\theta}loss&=loss_{Q}+\mu(loss_{Q},loss_{T})\cdot loss_{T},\\ loss_{Q}&=MSE(\widehat{\mathbf{Q}},\mathbf{Q}^{*}),\\ loss_{T}&=MSE(\widehat{\mathbf{T}},\mathbf{T}^{*})\end{split}$ (7)
-where $\theta$ denotes all of the learnable parameters of our framework, $\widehat{\mathbf{Q}}$, $\widehat{\mathbf{T}}$ and $\mathbf{Q}^{*}$, $\mathbf{T}^{*}$ respectively denote the predictions and labels of $\mathbf{Q}$ and $\mathbf{T}$.
-In order that the both of the SO(3)-equivariant and SO(3)-invariant branches can be stably converged, we introduce $\mu(loss_{Q},loss_{T})$, i.e., a function returning a coefficient to regularize the relative scale of the two loss terms:
-$\mu(loss_{Q},loss_{T})=\lambda\cdot No\_Grad(\frac{loss_{Q}}{loss_{T}})$ (8)
-where $No\_Grad(\cdot)$ denotes gradient discarding when calculating such coefficient, as this coefficient is only used for adjusting the relative scale of the two loss terms and should not itself be a source of training gradients, otherwise it would counteract $loss_{T}$ in Eq.
-(7).
-The introduction of $\mu(loss_{Q},loss_{T})$ effectively prevents the numerical disparity between the two loss terms to stabilize the training for both of them.
-All of the encoding and decoding modules are trained jointly by Eq.
-(7), and the training configurations are included in Supplementary Information D.
-In our framework, the SO(3)-invariant non-linear features $z^{(k)}(1\leq k\leq K)$ learned with the help of $\mathbf{T}$ contribute to the learning of SO(3)-equivariant representations $\textbf{v}^{(k)}(1\leq k\leq K)$ and the prediction of $\mathbf{Q}$.
-To achieve this goal, in the training phase, both the SO(3)-invariant feature encoder and the SO(3)-invariant decoder, used for predicting $\mathbf{T}$ from $z^{(k)}$, need to be trained.
-While in the inference phase, since $\mathbf{T}$ is not the final regression target, the SO(3)-invariant decoder can be removed, yet the SO(3)-invariant encoder still should be retained, used to induce SO(3)-equivariant features.
-6.3 Time Complexity
-Typically, one of the key advantage of deep learning algorithms for electronic-structure calculation lies in their low computational complexity which can asymptotically scale as $O(N)$, where $N$ is the number of atoms of a system.
-This is a significant efficiency advantage compared to the $O(N^{3})$ computational complexity of traditional DFT method, enabling efficient and low-cost prediction of electronic structures for large-scale atomic systems.
-Our method improves the accuracy of Hamiltonian predictions within the deep learning paradigm while maintaining this advantage of asymptotic complexity.
-The baseline models we select, whether DeepH-E3 or QHNet, are efficient models with an asymptotic computational complexity of $O(N)$ as $N$ increases.
-The neural module $g_{nonlin}(\cdot)$ we introduce efficiently operates on each individual feature f from the baseline models to obtain the corresponding non-linear SO(3)-equivariant feature v, rather than relying on complex interactions between multiple features to derive new ones.
-This ensures that the computational complexity remains asymptotically $O(N)$, with good potential to efficiently handle large atomic systems.
-7 Acknowledgements
-This work was supported by the National Natural Science Foundation of China (Grant Number 12134012) and the Strategic Priority Research Program of Chinese Academy of Sciences (Grant Number XDB0500201).
-Appendix A The Equivariance-Expressiveness Dilemma in Existing Deep Learning Methods
-In this section, we introduce existing efforts to harmonize the deep learning paradigm with SO(3)-equivariance, analyzing the dilemma between strict SO(3)-equivariance and non-linear expressive capabilities.
-This dilemma leads to an unsatisfactory trade-off between strict equivariance and non-linear expressiveness, ultimately limiting the achievable performance and scopes of application for these regression models.
-A.1 Approaches with Strict SO(3)-equivariance but Weak Expressiveness
-Most existing methods have focused on achieving SO(3)-equivariance at the neural mechanism level by applying group theory-guaranteed SO(3)-equivariant operators [26, 27].
-These operators, including linear scaling, element-wise sum, direct products, direct sums, and Clebsch-Gordan decomposition, have been used to build graph neural network architectures for tasks in 3D point cloud analysis [30], molecular dynamic simulation and property prediction [13, 12], as well as Hamiltonian prediction [31, 4, 8, 14].
-However, these operators could not provide sufficient expressiveness to fit complex non-linear mappings.
-The majority of these operators are linear operations, and the only one (direct product) that introduces some non-linearity can only approximate a simple squaring function, and cannot be extended to more complex non-linear forms due to the huge computational burden and difficulty in convergence.
-Even worse, directly applying highly-expressive non-linear activation functions, such as $Sigmoid$, $Softmax$, and $SiLU$, to SO(3)-equivariant features may result in the loss of equivariance.
-To alleviate this issue, these methods introduced a gated activation mechanism that processes SO(3)-invariant features ($l=0$) using non-linear activations and employs them as linear gating coefficients for SO(3)-equivariant features ($l\geq 1$), aiming to enhance their expressive power while maintaining strict equivariance.
-However, from the perspective of SO(3)-equivariant features, this mechanism is a combination of linear operations and cannot fundamentally increase their non-linear expressive capabilities.
-A.2 Approaches with Remarkable Expressiveness but Approximate SO(3)-equivariance
-Li et al.
-[6] proposed a local coordinate strategy to achieve SO(3)-equivariance.
-This strategy relied on the transformation relationship between local and global coordinates to achieve equivariance, without imposing restrictions on the neural network, allowing the use of network modules that possess non-linear expressiveness.
-However, this strategy, solely depending on coordinate transformations, only works under holistic rigid rotations of the atomic system.
-Without exploration of a neural mechanism for equivariance, when the local coordinates experience non-rigid perturbations such as thermal movements or bilayer twists, this strategy offers no guarantees on maintaining symmetry properties; Zitnick et al.
-[11] decomposed features into weighted sums of spherical harmonic basis functions, each corresponding to a discrete point on a unit sphere, through inner-product operations.
-Such discrete decomposition enables non-linear operations on the coefficients to enhance expressiveness.
-In subsequent developments [15, 16, 32], this approach has demonstrated a remarkable capacity to fit complex functions.
-However, the inner-product operations with discrete basis make such approach degenerate from continuous to discrete rotational equivariance, no longer maintaining strict equivariance to continuous rotational operations, as pointed out by Zhang et al.
-[1].
-Additionally, due to the high complexity of its discrete sampling strategies, tensor orders ($m$) corresponding to large degrees ($l$) are often partially pruned in architectures such as Equiformerv2 [16], further sacrificing the equivariance of features.
-In summary, although these methods with approximate SO(3)-equivariance achieve high numerical accuracy due to their extensive use of non-linear mechanisms, they are best suited for the applications that primarily require numerical approximations but not for tasks that demand theoretical depth in physics.
-In some applications, the demands for symmetry are extremely high.
-Even minor deviations from perfect symmetry can result in incorrect physical results.
-Combining these analyses, existing deep learning methods have compromised either strict SO(3)-equivariance or non-linear expressiveness when processing physical systems.
-In this paper, we develop a novel framework that theoretically combines strict SO(3)-equivariance with the non-linear expressiveness of neural networks.
-This solution resolves the equivariance-expressiveness dilemma by deeply exploring the intrinsic relationships between SO(3)-invariant and SO(3)-equivariant quantities and their representations.
-Appendix B Proof of Theory
-B.1 Proof of Theorem 1
-Proof.
+The second group of experiments, as included in Appendix I, test the extra time cost caused by TraceGrad compare to the baseline model.
+As we adopt a light-weight implementation of the non-linear modules introduced in the TraceGrad method, experimental results show that the extra time burden is slight relative to the baseline models DeepH-E3 and QHNet, indicating the good efficiency of our approach.
+The third group of experiments corresponds to the synergy of our method with an approximately SO(3)-equivariant methodology (Yin et al., 2024).
+As approximate equivariance is not the main focus of this paper, we just include these experimental results in Appendix J.
+The experimental results in Appendix J still demonstrate a significant accuracy improvement brought by our approach, further supporting the generality of our methodology.
+7 Conclusion
+We propose a theoretical and methodological framework to tackle the issue of reconciling non-linear expressiveness with SO(3)-equivariance in deep learning frameworks for physical system modeling, through deeply investigating the mathematical connections between SO(3)-invariant and SO(3)-equivariant quantities, as well as their representations.
+We first constructs SO(3)-invariant quantities from SO(3)-equivariant regression targets, using them to train informative SO(3)-invariant non-linear representations.
+From these, SO(3)-equivariant features are derived with gradient operations, achieving non-linear expressiveness while maintaining strict SO(3)-equivariance.
+We apply our theory and method to the challenging electronic-structure Hamiltonian prediction tasks, achieving dramatic promotions in prediction accuracy across eight benchmark databases.
+Experimental results demonstrate that this approach not only improves the accuracy of Hamiltonian prediction but also significantly enhances the prediction for downstream physical quantities, and also markedly improves the acceleration ratios for traditional DFT algorithms.
+Ethics Statement
+This work develops a representation learning method that exhibits strong non-linear expressive capabilities while strictly adhering to SO(3) equivariance.
+This method has demonstrated superior accuracy in predicting electronic-structure Hamiltonians and related physical quantities, showcasing its potential to accelerate research in materials science and molecular pharmacology.
+While we recognize that our research area has not yet revealed direct negative social or ethical implications, several issues warrant our vigilance.
+Currently, although our method yields accurate predictions, the decision-making processes of deep learning systems often lack transparency, hindering a comprehensive understanding of the learning outcomes and limiting our ability to gain deeper insights.
+We believe it is important to investigate the interpretability of such models, particularly in terms of how they apply physical knowledge in a comprehensible way.
+Additionally, it is crucial to continually improve the correctness and fairness of deep learning models on this area.
+Ensuring high-quality and diverse training data, implementing sound model designs, and performing ongoing validation and refinement are necessary to guarantee model accuracy and the broad applicability of their results.
+Appendices
+A Background
+Definition 1.
+Group. A set $G$, denoted as $G=\{\dots,g,\dots\}$, equipped with a binary operation denoted as $\cdot$, is called a group if it satisfies the following four conditions:
+  1. 1.
+    Closure: For any two elements $f,g\in G$, $f\cdot g\in G$.
+  2. 2.
+    Associativity: For all elements $f,g,h\in G$, the equality $(f\cdot g)\cdot h=f\cdot(g\cdot h)$ holds.
+  3. 3.
+    Existence of Identity Element: There exists a unique identity element $e\in G$ such that for all $f\in G$, $e\cdot f=f\cdot e=f$.
+  4. 4.
+    Existence of Inverse Element: For each element $f\in G$, there exists a unique inverse element $f^{-1}\in G$ such that $f\cdot f^{-1}=f^{-1}\cdot f=e$.
+Definition 2.
+SO(3) Group. The special orthogonal group $\text{SO}(3)$ is the group of all $3\times 3$ orthogonal matrices with determinant 1. Formally, it can be defined as:
+$\text{SO}(3)=\{\mathbf{R}\in\mathbb{R}^{3\times 3}\mid\mathbf{R}^{T}\mathbf{R}=I,\,\det(\mathbf{R})=1\}$
+where $\mathbf{R}^{T}$ denotes the transpose of $\mathbf{R}$ and $I$ is the $3\times 3$ identity matrix. The elements of $\text{SO}(3)$ represent rotations in three-dimensional Euclidean space.
+Definition 3.
+Group Representation.
+A representation of a group $G$ on a tensor space $T(V)$ is a homomorphism $\rho$ from $G$ to the general linear group $GL(T(V))$, the group of all invertible linear transformations on $T(V)$.
+Here, $T(V)$ represents the tensor space associated with the vector space $V$, encompassing all tensors that can be formed from elements of $V$.
+The homomorphism $\rho$ can be formalized as:
+$\rho:G\to GL(T(V))$
+such that for all $g_{1},g_{2}\in G$,
+$\rho(g_{1}g_{2})=\rho(g_{1})\rho(g_{2})$
+and $\rho(e)=I$, where $e$ is the identity element of $G$, and $I$ is the identity transformation on $T(V)$.
+Definition 4.
+SO(3) Group Representation.
+A representation of the special orthogonal group $\text{SO}(3)$ on a tensor space $T(V)$ is a homomorphism $\rho:\text{SO}(3)\to GL(T(V))$, where $GL(T(V))$ denotes the group of all invertible linear transformations on the tensor space $T(V)$.
+This representation can be expressed using Wigner-D matrices $\mathbf{D}^{l}(\mathbf{R})$, which are defined for a given degree $l$ corresponding to the representation of the rotation matrix $\mathbf{R}\in\text{SO}(3)$:
+$\rho(\mathbf{R})=\mathbf{D}^{l}(\mathbf{R}).$
+Definition 5.
+Equivariance with Respect to a Group.
+Let $G$ be a group, and let $\rho_{T(V)}:G\to GL(T(V))$ and $\rho_{T(W)}:G\to GL(T(W))$ be representations of $G$ on tensor spaces $T(V)$ and $T(W)$, respectively.
+A map $f:T(V)\to T(W)$ is said to be equivariant with respect to the group $G$ if the following condition holds:
+$f(\rho_{T(V)}(g)\circ v)=\rho_{T(W)}(g)\circ f(v)\quad\text{for all }v\in T(V)\,\,and\,\,g\in G.$
+where $\circ$ generally denotes the operation defined on the tensor space.
+Definition 6.
+Invariance with Respect to a Group.
+Let $G$ be a group, and let $\rho_{T(V)}:G\to GL(T(V))$ be a representation of $G$ on a tensor space $T(V)$.
+A function $f:T(V)\to T(W)$ is said to be invariant under the group $G$ if the following condition holds:
+$f(\rho_{T(V)}(g)\circ v)=f(v)\quad\text{for all }v\in T(V)\,\,and\,\,g\in G.$
+This definition indicates that the function $f$ remains unchanged under the action of the group $G$.
+B Proofs of Theorems
+Proof of Theorem 1.
 Under an SO(3) rotation represented by the rotational matrix $\mathbf{R}$, $\mathbf{Q}=\mathbf{Q}^{l_{p}\otimes l_{q}}$ is transformed as $\mathbf{Q}(\mathbf{R})$:
@@
 $\mathbf{Q}(\mathbf{R})^{\dagger}=\mathbf{D}^{l_{q}}(\mathbf{R})\cdot\mathbf{Q}^{\dagger}\cdot\mathbf{D}^{l_{p}}(\mathbf{R})^{\dagger}.$
-Thus, the product $\mathbf{Q}(\mathbf{R})\cdot\mathbf{Q}(\mathbf{R})^{\dagger}$ is:
-$\mathbf{Q}(\mathbf{R})\cdot\mathbf{Q}(\mathbf{R})^{\dagger}=(\mathbf{D}^{l_{p}}(\mathbf{R})\cdot\mathbf{Q}\cdot\mathbf{D}^{l_{q}}(\mathbf{R})^{\dagger})\cdot(\mathbf{D}^{l_{q}}(\mathbf{R})\cdot\mathbf{Q}^{\dagger}\cdot\mathbf{D}^{l_{p}}(\mathbf{R})^{\dagger}).$
 Using the cyclic property of the trace, which states that the trace of a product of matrices remains unchanged under cyclic permutations (i.e., $tr(ABC)=tr(BCA)=tr(CAB)$), and combining the properties that $\mathbf{D}^{l_{p}}(\mathbf{R})\cdot\mathbf{D}^{l_{p}}(\mathbf{R})^{\dagger}=\mathbf{I}$ and $\mathbf{D}^{l_{q}}(\mathbf{R})\cdot\mathbf{D}^{l_{q}}(\mathbf{R})^{\dagger}=\mathbf{I}$, we can rearrange the terms inside the trace as follows:
@@
 Therefore, $\mathbf{T}=tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})$ is invariant under SO(3) transformations, its SO(3)-invariance is proved. ∎
-B.2 Proof of Theorem 2
-Proof.
-The input feature $\mathbf{f}$ is SO(3)-equivariant, meaning that under an SO(3) rotation represented by $\mathbf{R}$, it transforms as follows:
+Proof of Theorem 2.
+Under the given condition, the input feature $\mathbf{f}$ in direct-sum state is SO(3)-equivariant, meaning that under an SO(3) rotation represented by $\mathbf{R}$, it transforms as follows:
 $\mathbf{f}(\mathbf{R})=\mathbf{D}^{l}(\mathbf{R})\cdot\mathbf{f}$
 where $\mathbf{D}^{l}(\mathbf{R})$ is the Wigner-D matrix corresponding to degree $l$.
-First, according to group theory, $u=\text{CGDecomp}(\mathbf{f}\otimes\mathbf{f},0)$ is an SO(3)-invariant scalar because the degree-zero component from the Clebsch-Gordan decomposition is invariant under rotations.
+First, according to group theory, $u=\text{CGDecomp}(\mathbf{f}\otimes\mathbf{f},0)$ is an SO(3)-invariant scalar as the degree-zero component from the Clebsch-Gordan decomposition is invariant under rotations.
 Since applying a non-linear operation to an SO(3)-invariant quantity does not change its invariance, $z=s_{\text{nonlin}}(u)$ is also SO(3)-invariant, independent to the specific form of $s_{\text{nonlin}}(\cdot)$.
 It formally holds that:
-$z=z(\mathbf{R})$ (9)
+$z(\mathbf{R})=z$ (9)
 Next, we apply the chain rule in Jacobian form.
-Considering $\mathbf{f}(\mathbf{R})$ is a column vector, to facilitate the application of the chain rule in vector form, we first transpose it into a row vector $\mathbf{f}(\mathbf{R})^{T}$, then differentiate:
+Considering $\mathbf{f}(\mathbf{R})$ is in the form of a column vector, to facilitate the application of the chain rule in vector form, we first transpose it into a row vector $\mathbf{f}(\mathbf{R})^{T}$, then differentiate:
 $\frac{\partial z(\mathbf{R})}{\partial\mathbf{f}^{T}(\mathbf{R})}=\frac{\partial z}{\partial\mathbf{f}^{T}(\mathbf{R})}=\frac{\partial z}{\partial\mathbf{f}^{T}}\frac{\partial\mathbf{f}^{T}}{\partial\mathbf{f}(\mathbf{R})^{T}}=\frac{\partial z}{\partial\mathbf{f}^{T}}\cdot\mathbf{D}^{l}(\mathbf{R})^{-1}=\frac{\partial z}{\partial\mathbf{f}^{T}}\cdot\mathbf{D}^{l}(\mathbf{R})^{T}$ (10)
-Here we utilize the property that $\mathbf{D}^{l}(\mathbf{R})^{-1}=\mathbf{D}^{l}(\mathbf{R})^{T}$.
-Since the representations of neural networks are generally real numbers, the Wigner-D matrix is also real unitary  [^3] In Theorem 1 and Theorem 2, the Wigner-D matrices are in the complex and real fields, respectively, because the output space may be complex while the internal representations of neural networks are typically real.
-Although the internal representations of neural network modules including $g_{\text{nonlin}}(\cdot)$ are in the real field, they are still able to predict targets in complex field with equivariance.
-Literature [8] has provided mechanisms for converting the network outputs in the real field into regression targets with real and imaginary parts..
+Here we utilize the property that $\mathbf{D}^{l}(\mathbf{R})^{-1}=\mathbf{D}^{l}(\mathbf{R})^{T}$  [^2] In Theorem 1 and Theorem 2, the Wigner-D matrices are in the complex and real fields, respectively, since the target quantity may be complex, whereas the internal representations of neural networks are typically in the real field.
+Nonetheless, neural network representations in the real field can still predict complex-valued targets with SO(3)-equivariance.
+Previous literature (Gong et al., 2023) has provided mechanisms for converting the network outputs in the real field into regression targets with real and imaginary parts..
+Since the representations of neural networks are generally real numbers, the corresponding Wigner-D matrix is also real unitary.
 Finally, we transpose the result back to a column vector:
@@
 ∎
-Appendix C Information of Experimental Databases
-In this part, we provide detailed information about the experimental databases, including the statistical information of the six databases from the DeepH series [6, 8], and the two databases from the QH9 series [19], listed in Table 5 and Table 6, respectively.
+C Information of Experimental Databases
+Figure 2: Illustration of a Hamiltonian matrix.
+Each Hamiltonian matrix contains multiple basic blocks, with the $\mathbf{Q}=\mathbf{Q}^{l_{p}\otimes l_{q}}$ defined in Section 3 corresponding to a basic block $\mathbf{H}^{l_{p}\otimes l_{q}}$ here.
+In this part, we provide detailed information about the experimental databases, including the statistical information of the six databases from the DeepH benchmark series (Li et al., 2022; Gong et al., 2023) and the two databases from the QH9 benchmark series (Yu et al., 2023a), listed in Table 4 and Table 5, respectively.
 Additionally, we visualize two types of challenging testing samples: samples with non-rigid deformation from thermal motions, as well as the bilayer samples with interlayer twists, which are shown in Fig.
 3 and Fig.
-5, respectively.
-Table 5: Statistic information of the six benchmark databases, i.e., Monolayer Graphene ($MG$), Monolayer MoS2 ($MM$), Bilayer Graphene ($BG$), Bilayer Bismuthene ($BB$), Bilayer Bi2Te3 ($BT$), Bilayer Bi2Se3 ($BS$), from the DeepH benchmark series [6, 8].
+4, respectively.
+D Implementation Details
+The hardware environment for our experiments is a server cluster equipped with Nvidia RTX A6000 GPUs, each with 48 GiB of memory.
+Other experimental details may differ across the DeepH and QH9 benchmark series, which we will describe separately.
+E Implementation Details on the DeepH Benchmark Series
+The software environment used is Pytorch 2.0.1 for experiments on the six crystalline databases from the DeepH benchmark series.
+When combining the proposed TraceGrad method with the DeepH-E3 architecture, the implementation of DeepH-E3 is based on the project  [^3] https://github.com/Xiaoxun-Gong/DeepH-E3 provided by Gong et al.
+(2023), keeping the architecture and hyperparameter configurations consistent with their setup.
+In our framework, we use the same number of encoding modules $K$ as DeepH-E3, which is set to $3$.
+In each encoding module, we apply the $g_{nonlin}(\cdot)$ module proposed in Section 4 and 5 to each SO(3)-equivariant edge feature, enabling non-linear expressiveness.
+We set the number of channels for the rotationally SO(3)-invariant feature $\textbf{u}^{(k)}$ $(1\leq k\leq 3)$ to $1024$.
+The neural network module $s_{nonlin}(\cdot)$ within $g_{nonlin}(\cdot)$ is implemented as a three-layers fully-connected module: the input size is set to $1024$, consistent with $\textbf{u}^{(k)}$; the hidden layer size is also $1024$, with SiLU as the non-linear activation function and LayerNorm as the normalization mechanism; and the output layer size (i.e., the dimensionality of $\textbf{z}^{(k)}$) is set to be equal to the number of basic blocks for a Hamiltonian matrix, which is $25$ for $MG$ and $BG$, $49$ for $MM$, and $196$ for $BB$, $BT$, and $BS$.
+It is worth noting that, while $s_{nonlin}(\cdot)$ can be implemented as any differentiable neural network module, we here implement it as a simple fully-connected module.
+This decision is made to avoid adding significant computational burden to the whole network.
+Meanwhile, as DeepH-E3 already incorporates complex graph network mechanisms for information aggregation and message-passing, there is no need for $s_{nonlin}(\cdot)$ to be overly complex.
+Its role is focusing on to filling in the gaps left by the existing equivariant mechanisms in DeepH-E3: to introduce a non-linear mapping mechanism that maintains equivariance, thereby activating and unleashing the expressive power of the overall network architecture through non-linearity.
+The SO(3)-equivariant decoder we adopt is the same as that of DeepH-E3; The SO(3)-invariant decoder we adopt is a four-layers fully-connected module: the input size is $3$ $(K)$ times of the dimensionality of $\textbf{z}^{k}$, e.g., $75$ for $MG$; the hidden layers have $1024$ neurons with SiLU as the non-linear activation function and LayerNorm as the normalization mechanism; the size of the output layer is the number of basic blocks for a Hamiltonian matrix.
+Since each basic block of the Hamiltonian matrix can compute a trace, the total number of trace variables corresponds to the number of basic blocks.
+Regarding the error metric in the loss function Eq.
+7, for the first term, we follow DeepH-E3 to use MSE (Mean Squared Error); for the second term, we choose between MSE and MAE based on performance on the validation sets, ultimately selecting MAE.
+$\lambda$ in the training loss function is set according to parameter selection on the validation sets, searching from $\{0.1,0.2,...,1.0\}$.
+Here, we aim to obtain a more general parameter setting for $\lambda$ on crystalline structures, and thus we determined $\lambda$ based on the overall performance on the validation sets of the six crystalline databases and the searched value is $0.3$.
+Other hyper-parameters and configurations are the same as DeepH-E3 (Gong et al., 2023): the initial learning rates for experiments on the $MG$, $MM$, $BG$, $BB$, $BT$, and $BS$ databases are set to $0.003$, $0.005$, $0.003$, $0.005$, $0.004$, and $0.005$, respectively; the training batch size is set as $1$; the optimizer is chosen as Adam; the scheduler is configured as a slippery slope scheduler.
+F Implementation Details on the QH9 Benchmark Series
+The software environment used is Pytorch 1.11.0 for experiments on the two molecular databases from the QH9 benchmark series.
+When combining the proposed TraceGrad method with the QHNet architecture, the implementation of QHNet is based on the project  [^4] https://github.com/divelab/AIRS provided by Yu et al.
+(2023b), keeping the architecture and hyperparameter configurations consistent with their setup.
+In our framework, we use the same number of encoding modules as QHNet: $5$ node feature encoding modules and $2$ edge feature encoding modules.
+We opt to apply the $g_{nonlin}(\cdot)$ module proposed in Section 4 and 5 to each SO(3)-equivariant edge feature.
+We set the number of channels for $\textbf{u}^{(k)}$ $(1\leq k\leq 2)$ as $1024$.
+The neural network module $s_{nonlin}(\cdot)$ within $g_{nonlin}(\cdot)$ is implemented as a three-layers fully-connected module: the input size is set to $1024$, consistent with $\textbf{u}^{(k)}$, the hidden layer size is also $1024$, with SiLU as the non-linear activation function and LayerNorm as the normalization mechanism, and the output layer size (i.e., the dimensionality of $\textbf{z}^{(k)}$) is set to be equal to the number of basic blocks for a Hamiltonian matrix, which is $36$ for $QS$ and $QD$ databases.
+The SO(3)-equivariant decoder we adopt is the same as that of QHNet; the SO(3)-invariant decoder we adopt is a four-layers fully-connected module: the input size is $2$ $(K)$ times of the dimensionality of $\textbf{z}^{(k)}$, e.g., $72$ for $QS$ and $QD$; the hidden layers have $1024$ neurons with SiLU as the non-linear activation function and LayerNorm as the normalization mechanism; the size of the output layer is the number of basic blocks for a Hamiltonian matrix.
+Regarding the error metric in the loss function Eq.
+7, for the first term, we follow QHNet to use a combination of MSE and MAE; for the second term, we choose between MSE and MAE based on performance on the validation sets, ultimately selecting MAE.
+$\lambda$ in the training loss function is set according to parameter selection on the validation sets, searching from $\{0.1,0.2,...,1.0\}$.
+Here, we aim to obtain a more general parameter setting for $\lambda$ on molecular structures, and thus we determined $\lambda$ based on the overall performance on the validation sets of the two molecular databases and the searched value is $0.2$.
+Other hyper-parameters and configurations are the same as QHNet: the initial learning rates for all experiments are set as $5\times 10^{-4}$, the training batch size is set as $32$, the optimizer is selected as AdamW, and a learning rate scheduler is implemented.
+To be more specific, the scheduler gradually increases the learning rate from 00 to a maximum value of $5\times 10^{-4}$ over the first $1,000$ warm-up steps.
+Subsequently, the scheduler linearly reduces the learning rate, ensuring it reaches $1\times 10^{-7}$ by the final step.
+G Visualization of Block-level MAE Statistics
+As shown in Fig.
+2, each Hamiltonian matrix consists of numerous basic block, with each basic block representing the direct product of two degrees.
+Here, we follow Yin et al.
+(2024) to measure the MAE performance of deep models on each basic block, denoted as $MAE^{H}_{block}$.
+The values of $MAE^{H}_{block}$ for the two setups, i.e., DeepH-E3 and DeepH-E3+TraceGrad, on different blocks of the Hamiltonian matrix for six databases from the DeepH benchmark series are illustrated in Fig.
+5 and 6.
+Fig.
+5 presents the results for monolayer structures, while Fig.
+6 focuses on bilayer structures.
+From these figures, it can be observed that our method, TraceGrad, brings significant accuracy improvements over the baseline method, DeepH-E3, across the vast majority of blocks of the Hamiltonian matrices, particularly on blocks where DeepH-E3 struggles with lower accuracy.
+Table 4: Statistical information of the six benchmark databases, i.e., Monolayer Graphene ($MG$), Monolayer MoS2 ($MM$), Bilayer Graphene ($BG$), Bilayer Bismuthene ($BB$), Bilayer Bi2Te3 ($BT$), Bilayer Bi2Se3 ($BS$), from the DeepH benchmark series (Li et al., 2022; Gong et al., 2023).
 SOC: effects of Spin-Orbit Coupling.
-$m$: number of samples in the current dataset; $a_{max}$: maximum number of atoms for a sample in the current dataset.
-$a_{min}$: minimum number of atoms for a sample in the current dataset.
+$m$: number of samples in the current dataset; $a_{max}$: maximum number of atoms from a unit cell in the current dataset.
+$a_{min}$: minimum number of atoms from a unit cell in the current dataset.
 $nt$: non-twisted samples.
@@
 $a_{min}$ 28 28 70 70
-Table 6: Basic statistics of the two benchmark databases, QH9-Stable ($QS$) and QH9-Dynamic ($QD$), from the QH9 benchmark series [19].
-For the $QS$ database, the ’ood’ strategy from the official settings is used to split the training, validation, and test sets, ensuring that the atom number of samples do not overlap across the three subsets.
-For the $QD$ database, the ’mol’ strategy from the official settings is employed to split the training, validation, and test sets, ensuring that there are no thermal motion samples from the same temporal trajectory across the three subsets.
-These split strategies aim to assess the regression model’s extrapolation capability with respect to the number of atoms as well as the temporal trajectories.
+Table 5: Statistical information of the two benchmark databases, QH9-stable ($QS$) and QH9-dynamic ($QD$), from the QH9 benchmark series (Yu et al., 2023a).
+The $QS$ database is split using the ’ood’ strategy, while the $QD$ database is split using the ’mol’ strategy.
 $m$: number of samples in the current dataset.
@@
 Figure 4: Visualization of testing samples with interlayer twists.
-Appendix D Training Configurations
-$\lambda$ in the training loss function is set as $0.3$ according to parameter selection on the validation sets.
-The initial learning rates for experiments on the $MG$, $MM$, $BG$, $BB$, $BT$ and $BS$ databases are set as $0.003$, $0.005$, $0.003$, $0.005$, $0.004$, and $0.005$, respectively, the training batch size for them is set as $1$ and the optimizer is selected as Adam; the initial learning rate for experiments on the $QS$ and $QD$ databases is set as $0.0005$, the training batch size is set as $16$, and the optimizer is selected as AdamW.
-Appendix E Visualization of Block-level MAE Statistics
-The MAE metrics (denoted as $MAE^{H}_{block}$) for the two setups, i.e., DeepH-E3 and DeepH-E3+TraceGrad, across different blocks of the Hamiltonian matrix for six databases from the DeepH benchmark series are illustrated in Fig.
-6 and 7.
-Fig.
-6 presents the results for monolayer structures, while Fig.
-7 focuses on bilayer structures.
-From these figures, it can be observed that our method, TraceGrad, brings significant accuracy improvements over the baseline method, i.e., DeepH-E3, on most of the blocks of the Hamiltonian matrix, especially on those blocks where DeepH-E3 performs poorly.
-Figure 6: Visualization of $MAE^{H}_{block}$ on each basic block of the Hamiltonian matrices for the $MG$ and $MM$ databases.
-Figure 7: Visualization of $MAE^{H}_{block}$ on each basic block of the Hamiltonian matrices for the non-twisted (marked with superscripts $nt$) and twisted (marked with superscripts $t$) structures of $BG$, $BB$, $BT$ and $BS$.
-Appendix F Ablation Study
+Figure 5: Visualization of $MAE^{H}_{block}$ on each basic block of the Hamiltonian matrices for the Monolayer Graphene ($MG$) and Monolayer MoS2 ($MM$) databases.
+Figure 6: Visualization of $MAE^{H}_{block}$ on each basic block of the Hamiltonian matrices for the non-twisted (marked with superscripts $nt$) and twisted (marked with superscripts $t$) testing subsets of Bilayer Graphene ($BG$), Bilayer Bismuthene ($BB$), Bilayer Bi2Te3 ($BT$), and Bilayer Bi2Se3 ($BS$).
+H Ablation Study
 We conduct fine-grained ablation study on the six databases from DeepH benchmark series, comparing results from the four setups:
-• DeepH-E3 [8]: the baseline model.
+• DeepH-E3 (Gong et al., 2023): the baseline model.
 • DeepH-E3+Trace: this experimental setup, an ablation term, only implements half part of our method.
 Specifically, it extends the architecture of DeepH-E3 by adding our SO(3)-invariant encoding and decoding branches and using the trace quantity $\mathbf{T}=\text{tr}(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})=\text{tr}(\mathbf{H}^{l_{p}\otimes l_{q}}\cdot(\mathbf{H}^{l_{p}\otimes l_{q}})^{\dagger})$ to train them.
-As for ablation study, this setup does not include the gradient-based mechanism for encoding SO(3)-equivariant non-linear features; instead, it directly uses the SO(3)-equivariant features outputted by DeepH-E3 for Hamiltonian regression.
+As for ablation study, this setup does not include the gradient-based mechanism delivering non-linear expressiveness from SO(3)-invariant features to encode SO(3)-equivariant features; instead, it directly uses the SO(3)-equivariant features outputted by DeepH-E3 for Hamiltonian regression.
+In this configuration, the SO(3)-invariant branches only contribute indirectly during the training phase by backpropagating the supervision signals from the trace quantity to the earlier layers.
 • DeepH-E3+Grad: this setup is also an ablation term and implements the other half part of our method in contrast to the previous ablation term.
-Specifically, it incorporates our SO(3)-invariant encoder branch as well as the gradient-induced operator to yield SO(3)-equivariant non-linear features based on the representations provided by the baseline encoders of DeepH-E3 for Hamiltonian prediction.
-As for ablation study, this setup continues to use the single-task training pipeline of DeepH-E3, predicting only the Hamiltonian without joint supervision training through the trace label.
+Specifically, it incorporates our SO(3)-invariant encoder branch as well as the gradient-induced operator to deliver non-linear expressiveness from SO(3)-invariant features to encode SO(3)-equivariant features.
+As for ablation study, this setup continues to use the single-task training pipeline of DeepH-E3, supervised only with the Hamiltonian label without joint supervised training through the trace of Hamiltonian.
 • DeepH-E3+TraceGrad: this is a complete implementation of our framework extending beyond of the architecture and training pipeline of DeepH-E3, at the label level, we introduce the trace quantity to guide the learning of SO(3)-invariant features; Meanwhile, at the representation level, we leverage the gradient operator to yield SO(3)-equivariant non-linear features for Hamiltonian prediction.
-Table 7: Ablation study MAE results (meV) on the $MG$ and $MM$ databases. $\downarrow$ means lower values of the metrics correspond to better accuracy.
+Table 6: Ablation study MAE results (meV) on the Monolayer Graphene ($MG$) and Monolayer MoS2 ($MM$) databases. $\downarrow$ means lower values of the metrics correspond to better accuracy.
 Methods $MG$ $MM$
@@
 DeepH-E3+TraceGrad 0.175 0.257 0.228 0.285 0.412 0.808
-Table 8: Ablation study MAE results (meV) on $BG$, $BB$, $BT$and $BS$.
+Table 7: Ablation study MAE results (meV) on the Bilayer Graphene ($BG$), Bilayer Bismuthene ($BB$), Bilayer Bi2Te3 ($BT$), and Bilayer Bi2Se3 ($BS$) databases.
 The superscripts $nt$ and $t$ respectively denote the non-twisted and twisted subsets.
@@
 DeepH-E3+TraceGrad 0.300 0.332 0.644 0.291 0.302 0.674
-Experimental results of the four setups are listed in Table 7 and 8.
+Experimental results of the four setups are listed in Table 6 and 7.
+Table 6 presents the results for monolayer structures, while Table 7 focuses on bilayer structures.
 From the results of ablation terms, we can obtain a more fine-grained experimental analysis.
@@
 Moreover, their combination provides even better performance.
+This is because, on one hand, with the gradient-based induction mechanism as a bridge, the non-linear expressiveness of SO(3)-invariant features learned from the trace label can be transformed into the SO(3)-equivariant representations during inference; on the other hand, with trace label, the SO(3)-invariant network branch has a strong supervisory signal, enabling it to learn the intrinsic symmetry and complexity of the regression targets, enhancing the quality of SO(3)-invariant features and ultimately benefits the encoding of SO(3)-equivariant features.
+The value of such complementarity has been fully demonstrated in the experimental results.
+I Discussion on Time Costs
+When the atomic system is small and any two atoms are within the cutoff radius of each other such that the Hamiltonian is defined, the number of Hamiltonian matrices is approximately proportional to $N^{2}$, where $N$ is the number of atoms in the system.
+Yet, as the atomic system grows, the system exhibits a dual nature of locality and non-locality.
+Locality refers to the fact that, when $N$ becomes large, the number of atoms within the cutoff radius of each atom remains roughly constant, and thus the growth in the number of Hamiltonian matrices asymptotically scales as $O(N)$.
+On the other hand, the non-locality arises from the fact that, although Hamiltonian matrices are defined only for pairs of atoms within each other’s cutoff radius, atoms outside the cutoff radius may still exert long-range effects on the Hamiltonian matrices within the cutoff radius
+The baseline models we select, whether DeepH-E3 (Gong et al., 2023) or QHNet (Yu et al., 2023b), are SO(3)-equivariant graph neural network models with efficient information aggregation and message-passing mechanisms, which cleverly balance the locality and non-locality of the system.
+As a result, the computational complexity asymptotically scales as $O(N)$ as $N$ grows to large.
+This is a significant efficiency advantage compared to the $O(N^{3})$ computational complexity of traditional DFT methods, enabling efficient and low-cost prediction of electronic structures for large-scale atomic systems.
+Our method improves the Hamiltonian prediction accuracy of DeepH-E3 and QHNet while maintaining this advantage of asymptotic complexity.
+The neural module $g_{nonlin}(\cdot)$ we introduce efficiently updates each edge feature $\textbf{f}^{(k)}$ $(1\leq k\leq K)$ from the baseline models to obtain the corresponding non-linear SO(3)-equivariant feature $\textbf{v}^{(k)}$, ensuring that the computational complexity remains asymptotically $O(N)$, with good potential to handle large atomic systems at acceptable computational costs.
+We test the inference time costs of the DeepH-E3 and DeepH-E3+TraceGrad models on an Nvidia RTX A6000 GPU, using the testing sets from two databases, namely $MG$ and $MM$, as examples; and test the inference time costs of the QHNet and QHNet+TraceGrad models on $QS$ and $QD$, on the same device.
+The batch size is set to 1 for all tests on inference time.
+Each sample is inferred 50 times, and the average time per inference is recorded.
+When using DeepH-E3, the average inference time per sample (s/sample) for $MG$ and $MM$ are as follows: $0.247$ and $0.256$, respectively; with DeepH-E3+TraceGrad, the average inference times are $0.264$ and $0.274$, correspondingly.
+When using QHNet, the average inference time per sample (s/sample) for $QS$ and $QD$ are as follows: $0.233$ and $0.174$, respectively; with QHNet+TraceGrad, the average inference times are $0.248$ and $0.187$, correspondingly.
+As we can see, the increases in computational costs due to the TraceGrad module are slight relative to that of the baseline models.
+Considering the significant accuracy improvements brought by the TraceGrad method as reported in Section 6.2, we believe that this minor increase in computational cost is acceptable and worthwhile in practical applications.
+J Empirical Study on Combining our Method with Approximately SO(3)-equivariance Framework
+While non-strict SO(3)-equivariance, which may limit the depth of theoretical exploration, is not the main focus of this study aiming at bridging rigorous SO(3)-equivariance with the non-linear expressive capabilities of neural networks, considering that they remain of interest in a few numerical computation applications where precision is highlighted over strict equivariance, we also conduct empirical study combining our method with approximately SO(3)-equivariant techniques.
+Taking the Monolayer MoS2 ($MM$) database as a case study, we evaluate the performance of combining our trace supervision and gradient induction method (TraceGrad) with the an approximately equivariant approach HarmoSE (Yin et al., 2024).
+We here take HarmoSE as the backbone encoder, and yields features by TraceGrad to enrich its representations.
+The experimental results in Table 8 and Fig.
+7 demonstrate that TraceGrad significantly enhances the accuracy of HarmoSE, surpassing DeepH-2 and achieving SOTA results.
+Both DeepH-2 and HarmoSE sacrificed strict SO(3)-equivariance to fully release the expressive capabilities of graph Transformers, aiming for the ultimate in prediction accuracy.
+Despite this, our method still manages to significantly exceed their accuracy, further confirming the superiority and generality of our method in learning expressive representations of physical systems.
+Figure 7: Comparison on the $MAE^{H}_{block}$ metric for HarmoSE and HarmoSE+TraceGrad on the $MM$ database.
+Table 8: MAE results (meV) for DeepH-2, HarmoSE, and HarmoSE+TraceGrad on the $MM$ database.
+$\downarrow$ means lower values of the metrics correspond to better accuracy.
+The results of the compared methods are taken from the corresponding literature (Wang et al., 2024b; Yin et al., 2024), where the empty items are due to the data not being provided in the original paper.
+Methods $MM$
+$MAE$ ()($\downarrow$)( ↓ )
+$MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$
+DeepH-2 (Wang et al., 2024b) 0.21 - -
+HarmoSE (Yin et al., 2024) 0.233 0.293 0.406
+HarmoSE+TraceGrad 0.178 0.228 0.296
+K Future Work
+In future research, various extensions are conceivable across theoretical, methodological, and application fields:
+First, from a theoretical and methodological perspective, we can generalize SO(3) group symmetry to a broader range and more complex groups in deep learning research.
+This will enable deep learning frameworks to accommodate richer mathematical structures and physical quantities.
+Second, from the application perspective, the current methodology extends beyond predicting electronic-structure Hamiltonians and is capable of generally predicting multiple physical quantities and properties that are equivariant.
+Furthermore, as our approach offers a promising general solution to the critical dilemma between equivariance and non-linear expressiveness for the deep learning paradigm, the methodology is not confined to physical research alone and may find applications in areas such as robotics, autonomous vehicles, and motion tracking systems.
+For instance, in the vision tasks of autonomous vehicles (Wang et al., 2023), where there is a high demand for robustness against coordinate transformations of cameras, the mainstream approach has been to achieve approximate SO(3)-equivariance through data augmentation.
+However, this does not guarantee absolute reliability and safety.
+Our work holds great potential to construct deep models with high generalization performance under the premise of strict SO(3)-equivariance, potentially revolutionizing techniques in such tasks.
+Looking forward to future endeavors, the foundational and pivotal role of this work is clear.

### v5

@@
-A Framework of SO(3)-equivariant Non-linear Representation Learning and its Application to Electronic-Structure Hamiltonian Prediction
+TraceGrad: a Framework Learning Expressive SO(3)-equivariant Non-linear Representations for Electronic-Structure Hamiltonian Prediction
 Abstract
-We propose both a theoretical and a methodological framework to address a critical challenge in applying deep learning to physical systems: the reconciliation of non-linear expressiveness with SO(3)-equivariance in predictions of SO(3)-equivariant quantities.
-Inspired by covariant theory in physics, we present a solution by exploring the mathematical relationships between SO(3)-invariant and SO(3)-equivariant quantities and their representations.
-We first construct theoretical SO(3)-invariant quantities derived from the SO(3)-equivariant regression targets, and use these invariant quantities as supervisory labels to guide the learning of high-quality SO(3)-invariant features.
-Given that SO(3)-invariance is preserved under non-linear operations, the encoding process for invariant features can extensively utilize non-linear mappings, thereby fully capturing the non-linear patterns inherent in physical systems.
+We propose a framework to combine strong non-linear expressiveness with strict SO(3)-equivariance in prediction of the electronic-structure Hamiltonian, by exploring the mathematical relationships between SO(3)-invariant and SO(3)-equivariant quantities and their representations.
+The proposed framework, called TraceGrad, first constructs theoretical SO(3)-invariant trace quantities derived from the Hamiltonian targets, and use these invariant quantities as supervisory labels to guide the learning of high-quality SO(3)-invariant features.
+Given that SO(3)-invariance is preserved under non-linear operations, the learning of invariant features can extensively utilize non-linear mappings, thereby fully capturing the non-linear patterns inherent in physical systems.
 Building on this, we propose a gradient-based mechanism to induce SO(3)-equivariant encodings of various degrees from the learned SO(3)-invariant features.
-This mechanism can incorporate non-linear expressive capabilities into SO(3)-equivariant representations, while theoretically preserving their equivariant properties as we prove, establishing a strong foundation for regressing complex SO(3)-equivariant targets.
-We apply our theory and method to the electronic-structure Hamiltonian prediction tasks, experimental results on eight benchmark databases covering multiple types of systems and challenging scenarios show substantial improvements on the state-of-the-art prediction accuracy of deep learning paradigm.
-Our method boosts Hamiltonian prediction accuracy by up to 40% and enhances downstream physical quantities, such as occupied orbital energy, by a maximum of 76%.
-Our method also significantly promotes the acceleration ratios for the convergence of traditional Density Functional Theory (DFT) methods.
+This mechanism can incorporate powerful non-linear expressive capabilities into SO(3)-equivariant features with consistency of physical dimensions to the regression targets, while theoretically preserving equivariant properties, establishing a strong foundation for predicting Hamiltonian.
+Our method achieves state-of-the-art performance in prediction accuracy across eight challenging benchmark databases on Hamiltonian prediction.
+Experimental results demonstrate that this approach not only improves the accuracy of Hamiltonian prediction but also significantly enhances the prediction for downstream physical quantities, and also markedly improves the acceleration performance for the traditional Density Functional Theory algorithms.
+Electronic-structure Hamiltonian prediction; SO(3)-equivariant representation learning; Non-linear expressiveness
 1 Introduction
-With the advantages in computational complexity and generalization capabilities, deep learning paradigm has vigorously driven advancements in physics research (Zhang et al., 2023).
-For example, in predicting electronic-structure Hamiltonians, traditional Density Functional Theory (DFT) methods (Hohenberg & Kohn, 1964; Kohn & Sham, 1965) suffer from computational complexities of $\mathcal{O}(N^{3})$, where $N$ is the number of atoms in a system.
-In contrast, deep learning approaches have significantly reduced this complexity to $\mathcal{O}(N)$ (Unke et al., 2021; Gu et al., 2022; Li et al., 2022; Yu et al., 2023b; Gong et al., 2023), opening up new possibilities for analyzing extremely large atomic systems, enabling efficient materials simulation and design, as well as groundbreaking molecular pharmaceutical research that was previously unimaginable.
-However, deep learning methods still face substantial challenges when processing physical systems.
-To align with fundamental physical laws, these methods must strictly adhere to symmetry principles.
-For example, physical quantities such as force fields and electronic-structure Hamiltonians must be equivariant under 3D rotational transformations, i.e., elements from the SO(3) group.
-Besides, the calculation of physical quantities calls for high numerical accuracy, necessitating that the neural networks possess strong capabilities to express the complex non-linear mappings from atomic structures to the regression targets.
-However, it is challenging for deep learning methods to simultaneously ensure strict SO(3)-equivariance and high numerical accuracy on modeling physical systems.
+Electronic structure calculations provide crucial insights into the behavior of electrons in condensed matter, forming the foundation for predicting material properties such as conductivity, magnetism, optical response, and chemical reactivity, playing a key role in applications ranging from semiconductors to renewable energy and catalysis, ultimately helping to reshape our lives.
+The core of electronic structure calculations lies in solving the Hamiltonian matrices, from which critical physical quantities such as orbital energy, band structure and electronic wavefunction are induced.
+However, traditional Density Functional Theory (DFT) methods (Hohenberg & Kohn, 1964; Kohn & Sham, 1965) suffer from the extremely time-consuming self-consistent field (SCF) algorithms used to obtain the Hamiltonians.
+These algorithms involve the exhaustive iterative diagonalization of large matrices, with each diagonalization scaling with a high computational complexity of $\mathcal{O}(N^{3})$, where $N$ is the number of atoms in the system.
+With advantages in computational complexity and generalization capabilities, the deep learning paradigm has significantly advanced physics research (Zhang et al., 2023).
+In the task of predicting the Hamiltonians, deep learning approaches (Unke et al., 2021; Yu et al., 2023b; Gong et al., 2023; Zhang et al., 2024; Wang et al., 2024c) bypass the SCF process, dramatically reducing the computational complexity of solving Hamiltonians, opening up new possibilities for analyzing extremely large atomic systems, enabling efficient materials simulation and design that was previously unimaginable.
+See Appendix B for more details.
+To align with fundamental physical laws, deep learning methods for Hamiltonian prediction must strictly adhere to equivariance under 3D rotational transformations, i.e., elements from the SO(3) group.
+However, it remains challenging for deep learning methods to simultaneously ensure strict SO(3)-equivariance and high numerical accuracy on predicting Hamiltonians.
 The root cause of this problem lies in the conflict between SO(3)-equivariance and non-linear expressiveness: specifically, directly applying non-linear activation functions on SO(3)-equivariant features (with degree $l\geq 1$) may lead to the loss of equivariance, while bypassing non-linear mappings severely restricts the network’s expressive capabilities and thereby lowering down the achievable accuracy.
-This issue is commonly found in physics-oriented machine learning tasks that demand both strict equivariance and fine-grained generalization performance, as analyzed by Zitnick et al.
+This issue is commonly encountered in physics-oriented machine learning tasks that require both strict equivariance and fine-grained generalization performance, as discussed by Zitnick et al.
 (2022).
-Several recent efforts have been made to alleviate this issue (Zitnick et al., 2022; Passaro & Zitnick, 2023; Wang et al., 2024c; Yin et al., 2024), however, they often compromised strict SO(3)-equivariance.
-To address the equivariance-expressiveness dilemma, we make theoretical and methodological explorations on unifying strict SO(3)-equivariance with strong non-linear expressiveness within the realm of deep representation learning.
-We are inspired by the insight that invariant quantities in transformation often reflect the mathematical nature of physical laws and can induce other quantities with equivariant properties.
-For example, in special relativity (Resnick, 1991), the spacetime interval between events is invariant under Lorentz transformations, and is fundamental in formulating physical laws that involve equivariant quantities.
-For another example, the total electronic energy of an atomic system is SO(3)-invariant, while the derived force fields are SO(3)-equivariant.
-We aim to extend the relationship between invariance and equivariance from specific physical quantities to a more general representation learning framework.
+Several recent methods, such as the gated mechanism (Weiler et al., 2018), spherical channel-based approaches (Zitnick et al., 2022; Liao et al., 2024; Wang et al., 2024c), local coordinate strategies (Li et al., 2022), and hybrid regression methods (Yin et al., 2024), have been proposed to address this challenge.
+However, combining strong non-linear expressiveness with strict SO(3)-equivariance remains challenging for the Hamiltonian prediction task, as analyzed in detail in Section 2.
+To address the equivariance-expressiveness dilemma, we make theoretical and methodological explorations on unifying strict SO(3)-equivariance with strong non-linear expressiveness within the realm of deep representation learning for electronic-structure Hamiltonian prediction.
+We are inspired by the insight that invariant quantities in transformation (Resnick, 1991) often reflect the mathematical nature of physical laws and can induce other quantities with equivariant properties, and aim to extend the relationship between invariance and equivariance from specific physical quantities to hidden representations of neural networks.
 From the perspective of deep representation learning, the attribute that invariance is preserved under non-linear operations is a significant advantage, given its compatibility with non-linear expressive capabilities.
-Built upon these insights, we propose a solution to the equivariance-expressiveness dilemma by intensively exploring and making use of the intrinsic relationships between SO(3)-invariant and SO(3)-equivariant quantities and representations: we first dedicate efforts to learning high-quality SO(3)-invariant features with ample non-linear expressiveness, and subsequently, we derive SO(3)-equivariant non-linear representations and the target quantities from these SO(3)-invariant ones.
+Built upon these insights, we propose a solution to the equivariance-expressiveness dilemma by intensively exploring and making use of the intrinsic relationships between SO(3)-invariant and SO(3)-equivariant quantities and representations: we first dedicate efforts to supervising high-quality SO(3)-invariant features with ample non-linear expressiveness, and subsequently, we derive SO(3)-equivariant non-linear representations and the target quantities from these SO(3)-invariant ones.
 Specifically:
-First, we propose a theoretical construct of SO(3)-invariant quantities, namely $tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})$, where $tr(\cdot)$ signifies the trace operation, $\dagger$ denotes the conjugate transpose operation, and $\mathbf{Q}$ denotes the SO(3)-equivariant regression targets we aim to predict.
-A significant advantage of these SO(3)-invariant quantities lies in the fact that they are directly derived from the SO(3)-equivariant target labels and can serve as unique supervision labels for the effective learning of informative SO(3)-invariant features that capture the intrinsic symmetry properties of the mathematical structure of $\mathbf{Q}$ without requiring additional labeling resources.
-Second, we propose a gradient-based mechanism to induce SO(3)-equivariant representations for predicting the regression target $\mathbf{Q}$ in the inference phase from high-quality non-linear SO(3)-invariant features learned under the supervision of $tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})$ in the training phase.
-Taking SO(3)-invariant features as a bridge, this mechanism can incorporate non-linear expressive capabilities into SO(3)-equivariant representations while preserving their equivariant properties, as we prove, laying a solid foundation for accurately inferring complex SO(3)-equivariant targets.
-We develop our theory into a unified SO(3)-equivariant non-linear representation learning method, and apply it to the computation of electronic-structure Hamiltonian, which occupies a central position in quantum mechanics to induce key physical quantities such as electronic wavefuction, orbital energy, and band structure, bringing deep understanding of the electrical, optical, magnetic, and transport properties of atomic systems, while posing significant difficulties for machine learning techniques due to its intrinsic symmetry with respect to the SO(3) group as well as the high-dimensional complexity, as pointed out by Yin et al.
-(2024).
-Our method significantly improves the state-of-the-art performance in Hamiltonian prediction on eight databases from the well-known DeepH and QH9 benchmark series (Li et al., 2022; Gong et al., 2023; Yu et al., 2023a).
+First, we propose a theoretical construct of SO(3)-invariant quantities, namely $tr(\mathbf{H}\cdot\mathbf{H}^{\dagger})$, where $tr(\cdot)$ signifies the trace operation, $\dagger$ denotes the conjugate transpose operation, and $\mathbf{H}$ denotes the SO(3)-equivariant regression targets, i.e., the block of Hamiltonian matrix.
+A significant advantage of these SO(3)-invariant quantities lies in the fact that they are directly derived from the SO(3)-equivariant target labels and can serve as unique supervision labels for the effective learning of informative SO(3)-invariant features that capture the intrinsic symmetry properties of the mathematical structure of $\mathbf{H}$ without requiring additional labeling resources.
+Second, we propose a gradient-based mechanism to induce SO(3)-equivariant representations for predicting the regression target $\mathbf{H}$ in the inference phase from high-quality non-linear SO(3)-invariant features learned under the supervision of $tr(\mathbf{H}\cdot\mathbf{H}^{\dagger})$ in the training phase.
+Taking SO(3)-invariant features as a bridge, this mechanism can incorporate powerful non-linear expressive capabilities into SO(3)-equivariant representations while preserving their equivariant properties, as we prove.
+Compared to the gated mechanism (Weiler et al., 2018), the proposed mechanism naturally follows the correspondence of physical dimensions between the SO(3)-equivariant Hamiltonian and the SO(3)-invariant trace quantity, laying a solid foundation for regressing Hamiltonian with the help of supervision signals from the trace quantity.
+Experimental results demonstrate that, our TraceGrad framework significantly improves the state-of-the-art performance in Hamiltonian prediction on eight databases from the well-known DeepH and QH9 benchmark series (Li et al., 2022; Gong et al., 2023; Yu et al., 2023a).
 It demonstrates excellent generalization performance to both crystalline and molecular systems, covering challenging scenarios such as thermal motions, bilayer twists, scale variations, and new trajectories.
@@
 2 Related Work
-The SO(3)-equivariant representation learning paradigm (Thomas et al., 2018; Geiger & Smidt, 2022) typically developed group theory-based symmetry operators, such as linear scaling, element-wise sum, direct products, direct sums, Clebsch-Gordan decomposition, and equivariant normalization, to encode equivariant features.
-These operators have been used to construct graph neural network architectures for tasks in 3D point cloud analysis (Fuchs et al., 2020), molecular property prediction and dynamic simulation (Musaelian et al., 2023; Liao & Smidt, 2023), as well as Hamiltonian prediction (Schütt et al., 2019; Unke et al., 2021; Gong et al., 2023; Zhong et al., 2023).
-However, as non-linear activation functions may result in the loss of equivariance, they are restricted when applied to SO(3)-equivariant features with degree $l$ greater than one.
+The SO(3)-equivariant representation learning paradigm (Thomas et al., 2018; Geiger & Smidt, 2022) typically developed group theory-based tensor operators, such as linear scaling, element-wise sum, direct products, direct sums, Clebsch-Gordan decomposition, and equivariant normalization, to encode equivariant features.
+These operators have been used to construct graph neural network architectures for tasks in 3D point cloud analysis (Fuchs et al., 2020), energy and force prediction (Liao & Smidt, 2023), as well as Hamiltonian prediction (Schütt et al., 2019; Unke et al., 2021; Gong et al., 2023; Zhong et al., 2023).
+However, as non-linear activation functions may result in the loss of equivariance, they are restricted when applied to SO(3)-equivariant features with degree $l\geq 1$.
 This restriction severely limits the network’s capability to model complex non-linear mappings.
-To alleviate this issue, methods like DeepH-E3 (Gong et al., 2023), QHNet (Yu et al., 2023b) and Equiformer (Liao & Smidt, 2023) introduced a gated activation mechanism that feeds SO(3)-invariant features ($l=0$) into non-linear activation functions and uses these features as linear gating coefficients for SO(3)-equivariant features ($l\geq 1$), aiming to enhance their expressive power while maintaining strict equivariance.
-Nevertheless, multiplying SO(3)-equivariant features with linear coefficients may not fundamentally improve their non-linear expressiveness.
+To alleviate this issue, Hamiltonian prediction methods like DeepH-E3 (Gong et al., 2023) and QHNet (Yu et al., 2023b) introduced a gated activation mechanism (Weiler et al., 2018) that fed SO(3)-invariant features ($l=0$) into non-linear activation functions and used these features as linear gating coefficients for SO(3)-equivariant features ($l\geq 1$), aiming to enhance their expressive power while maintaining strict equivariance.
+However, these methods only indirectly supervised the learning of SO(3)-invariant features through SO(3)-equivariant Hamiltonians taking SO(3)-equivariant representations as a bridge, and the lack of direct SO(3)-invariant supervision signals may result in insufficient learning of SO(3)-invariant features, limiting their expressiveness to assist SO(3)-equivariant features.
+Furthermore, the gated mechanism, which directly multiplies SO(3)-equivariant features with SO(3)-invariant features to yield new SO(3)-equivariant features, lacks consideration of the physical dimensional relationship between the actual SO(3)-equivariant and SO(3)-invariant physical quantities, making it difficult to achieve a perfect coupling of their representations.
+See Remark 4.3 for a more comprehensive analysis.
 In order to improve non-linear expressiveness, Zitnick et al.
-(2022) decomposed SO(3)-equivariant features into SO(3)-invariant coefficients of spherical harmonic basis functions.
+(2022) proposed spherical channel methods, which decomposed SO(3)-equivariant features into SO(3)-invariant coefficients of spherical harmonic basis functions.
 These SO(3)-invariant coefficients were processed by non-linear neural networks to enhance expressiveness, with equivariance regained by recombining the updated coefficients with the basis functions.
-In subsequent developments (Passaro & Zitnick, 2023; Liao et al., 2024; Wang et al., 2024b; c), this approach has demonstrated a remarkable capacity to fit complex functions.
+In subsequent developments (Passaro & Zitnick, 2023; Liao et al., 2024; Wang et al., 2024b, c), this approach has demonstrated a remarkable capacity to fit complex functions.
 However, as pointed out by existing literature (Zhang et al., 2023), this approach degenerates from continuous to discrete rotational equivariance, losing strict equivariance to continuous rotational transformations due to the decomposition based on inner-product operations with discrete basis functions; Li et al.
@@
 However, this strategy is effective only for global rigid rotations and fails to maintain symmetry under non-rigid perturbations like thermal fluctuations or bilayer twists, as it lacks a neural mechanism to enforce equivariance; Yin et al.
-(2024) proposed a hybrid framework consisting of both group theory-guaranteed SO(3)-equivariant mechanisms and non-linear mechanisms to regress Hamiltonians.
-In this framework, the non-linear mechanisms showed remarkable capability at learning SO(3)-equivariance from the data with the help of the theoretically SO(3)-equivariant mechanisms, and released powerful non-linear expressive capabilities to achieve more numerical accuracy.
+(2024) proposed a hybrid regression framework, where the non-linear mechanisms showed remarkable capability at learning SO(3)-equivariance from the data with the help of the theoretically SO(3)-equivariant mechanisms, and released powerful non-linear expressive capabilities to achieve more numerical accuracy.
 However, the equivariance achieved through data-driven methods does not have strict theoretical guarantee, even with rotational data augmentation.
-In some applications, the demands for symmetry are extremely high.
+In many applications of electronic structure calculations, the demands for symmetry are extremely high.
 Even minor deviations from perfect SO(3)-equivariance can result in incorrect physical results.
-In this paper, we propose a framework that theoretically combines strict SO(3)-equivariance with the non-linear expressiveness of neural networks to resolve the equivariance-expressiveness dilemma.
-3 Problem Formalization
-For a background introduction on groups, group representations, equivariance, and invariance, please refer to Appendix A.
-Here, we directly focus on the equivariance of physical quantities under 3D rotational operations that form the SO(3) group.
-Let ${\bf{Q}}^{l_{p}\otimes l_{q}}$ denote an SO(3)-equivariant quantity in direct-product state formed by $l_{p}\otimes l_{q}$, i.e., the direct product between degrees $l_{p}$ and $l_{q}$.
-It obeys the following SO(3)-equivariant law:
-${\bf{Q}}({\bf{R}})^{l_{p}\otimes l_{q}}={\bf{D}}^{l_{p}}({\bf{R}})\cdot{\bf{Q}}^{l_{p}\otimes l_{q}}\cdot{({\bf{D}}^{l_{q}}({\bf{R}}))^{\dagger}}$ (1)
-where $\dagger$ denotes the conjugate transpose operation.
-${\bf{R}}\in\mathbb{R}^{3\times 3}$ is the rotational matrix, ${\bf{D}}^{l_{p}}({\bf{R}})\in\mathbb{R}^{(2l_{p}+1)\times(2l_{p}+1)}$ and ${\bf{D}}^{l_{q}}({\bf{R}})\in\mathbb{R}^{(2l_{q}+1)\times(2l_{q}+1)}$ are the Wigner-D matrices of degrees $l_{p}$ and $l_{q}$, respectively; ${\bf{Q}}({\bf{R}})^{l_{p}\otimes l_{q}}\in\mathbb{R}^{(2l_{p}+1)\times(2l_{q}+1)}$ denotes the transformed results of ${\bf{Q}}^{l_{p}\otimes l_{q}}\in\mathbb{R}^{(2l_{p}+1)\times(2l_{q}+1)}$ through the rotational operation by ${\bf{R}}$.
-${\bf{Q}}^{l_{p}\otimes l_{q}}$ in the direct-product state can be further decomposed into a series of direct-sum state components, i.e., ${\bf{q}}^{l}(|l_{p}-l_{q}|\leq l\leq l_{p}+l_{q})$, which follows SO(3)-equivariant law mathematically equivalent to Eq.
-1 but with a simpler form:
-${\bf{q}}({\bf{R}})^{l}={\bf{D}}^{l}({\bf{R}})\cdot{\bf{q}}^{l},|l_{p}-l_{q}|\leq l\leq l_{p}+l_{q}$ (2)
-where ${\bf{q}}^{l}\in\mathbb{R}^{2l+1}$ and ${\bf{q}}({\bf{R}})^{l}\in\mathbb{R}^{2l+1}$ respectively denote the components with degree $l$ before and after the rotational operation by ${\bf{R}}$.
-For ease of processing, the internal representations of SO(3)-equivariant neural networks (Gong et al., 2023; Liao & Smidt, 2023) are typically in the direct-sum form.
-To obey the equivariant law of Eq.
-2 for the regression target, these hidden representations must also satisfy the same form of equivariance:
-${\bf{f}}({\bf{R}})^{(k)l}={\bf{D}}^{l}({\bf{R}})\cdot{\bf{f}}^{(k)l}$ (3)
-where ${\bf{f}}^{(k)l}\in\mathbb{R}^{2l+1}$ and ${\bf{f}}({\bf{R}})^{(k)l}\in\mathbb{R}^{2l+1}$ respectively denote one channel of hidden features with degree $l$ before and after the rotational operation by ${\bf{R}}$, at the $k$ th hidden layer.
-Due to the intrinsic complexity and non-linearity of physic quantities, neural networks on regressing these quantities are supposed to equip with non-linear mappings to fully capture the intrinsic patterns of the physical quantities, which is crucial for precise and generalizable prediction performance.
-Meanwhile, the non-linear mappings, denoted as $g_{nonlin}(\cdot)$, must also preserve SO(3)-equivariance, which is expressed as:
-${\bf{f}}({\bf{R}})^{(k+1)l}={\bf{D}}^{l}({\bf{R}})\cdot{\bf{f}}^{(k+1)l},\,\,\,\,\text{subject to}\,\,{\bf{f}}^{(k+1)l}=g_{nonlin}({\bf{f}}^{(k)l})$ (4)
-However, directly implementing $g_{nonlin}(\cdot)$ as neural network module with non-linear activation functions, such as $Sigmoid$, $Softmax$ and $SiLU$, may result in the destruction of strict equivariance.
-How to make $g_{nonlin}(\cdot)$ both theoretically SO(3)-equivariant and capable of non-linear expressiveness, and effectively apply it to the prediction of SO(3)-equivariant complex physical quantities, i.e., electronic-structure Hamiltonian in this context, is the core problem this paper aims at solving.
+In this paper, we seek for combining strict SO(3)-equivariance with the powerful non-linear neural expressiveness to resolve the equivariance-expressiveness dilemma for electronic-structure Hamiltonian prediction.
+3 Preliminary
+In order to better understand our theory and methods, we recommend that the readers first refer to the preliminary in Appendices: Appendix A provides the definitions of foundational concepts relevant to the research topic; Appendix B describes the application tasks, specifically the electronic-structure Hamiltonian prediction task; Appendix C formalizes Hamiltonian prediction as an SO(3)-equivariant non-linear representation learning problem.
 4 Theory
-Theorem 1.
-The quantity $\mathbf{T}=tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})$ is SO(3)-invariant, where $\mathbf{Q}$ is the simplified representation (without superscripts) of $\mathbf{Q}^{l_{p}\otimes l_{q}}$ defined in Section 3, and $\dagger$ denotes the conjugate transpose operation, $tr(\cdot)$ is the trace operation.
-Theorem 2.
+Theorem 4.1.
+The quantity $\mathbf{T}=tr(\mathbf{H}\cdot\mathbf{H}^{\dagger})$ is SO(3)-invariant, where $\mathbf{H}$ is the simplified representation (without superscripts) of the basic block $\mathbf{H}^{l^{i}_{p}\otimes l^{j}_{q}}$ of Hamiltonian matrix defined in Appendix B, and $\dagger$ denotes the conjugate transpose operation, $tr(\cdot)$ is the trace operation.
+Theorem 4.2.
 The non-linear neural mapping $g_{nonlin}(\cdot)$ defined as the following is SO(3)-equivariant:
-$\mathbf{v}=g_{nonlin}(\mathbf{f})=\frac{\partial z}{\partial\mathbf{f}},\quad\text{subject to}\quad z=s_{nonlin}(u),u=\text{CGDecomp}(\mathbf{f}\otimes\mathbf{f},0)$ (5)
-where $\mathbf{f}$ is an input SO(3)-equivariant feature with degree $l$ in the direct-sum state, $\otimes$ denotes the direct-product operation of tensors, $\text{CGDecomp}(\cdot,0)$ refers to performing a Clebsch-Gordan decomposition of the tensor and returning the scalar component of degree 0, $s_{nonlin}(\cdot)$ represents arbitrary differentiable non-linear neural modules, and $\mathbf{v}$ is the outputted feature encoded by $g_{nonlin}(\cdot)$.
-The proofs of the two theorems above are presented in Appendix B.
-Figure 1: Methodological framework for learning SO(3)-equivariant representations with non-linear expressiveness to regress complex SO(3)-equivariant targets.
+$\mathbf{v}=g_{nonlin}(\mathbf{f})=\frac{\partial z}{\partial\mathbf{f}},\quad$ (1)
+$\text{subject to}\quad z=s_{nonlin}(u),u=\text{CGDecomp}(\mathbf{f}\otimes\mathbf{f},0)$
+where $\mathbf{f}$ is an input SO(3)-equivariant feature with degree $l$ in the direct-sum state, $\otimes$ denotes the direct-product operation of tensors, $\text{CGDecomp}(\cdot,0)$ refers to performing a Clebsch-Gordan decomposition of the tensor and returning the SO(3)-invariant component corresponding to degree 0, $s_{nonlin}(\cdot)$ represents arbitrary differentiable non-linear neural modules, $u$ and $z$ are SO(3)-invariant features, and $\mathbf{v}$ is the outputted feature encoded by $g_{nonlin}(\cdot)$.
+The proofs of the two theorems above are presented in Appendix D.
+Remark 4.3.
+In physics, dimensions describe the fundamental properties of physical quantities.
+The seven base dimensions are mass $[M]$, time $[T]$, length $[L]$, electric current $[I]$, temperature $[\Theta]$, amount of substance $[N]$, and luminous intensity $[J]$.
+Other physical quantities can be derived from these base dimensions.
+For instance, the dimension of energy $E$ is given by $\dim E=ML^{2}T^{-2}$.
+The quantities $\mathbf{H}$ and $\mathbf{T}$ presented in Theorem 4.1 correspond to the physical dimensions of $\dim E$ and $\dim E^{2}$, respectively, where their units can be taken as meV and (meV)2.
+These two physical dimensions are related through: $\dim E=\frac{\dim E^{2}}{\dim E}$.
+As we developed in Section 5, in the feature space, the SO(3)-invariant feature $z$ is used to regress $\mathbf{T}$, and the SO(3)-equivariant features $\mathbf{f}$ and $\mathbf{v}$ are used to regress $\mathbf{H}$.
+Our gradient-based mechanism $\mathbf{v}=\frac{\partial z}{\partial\mathbf{f}}$ in Theorem 4.2 naturally preserves the dimensional correspondence between $\mathbf{H}$ and $\mathbf{T}$.
+On the other hand, the existing gated activation function (Weiler et al., 2018), which can be expressed as $\mathbf{v}=z\cdot\mathbf{f}$, does not enforce such a dimensional correspondence and could not offer a strong guarantee of a balanced coupling between the SO(3)-invariant and SO(3)-equivariant representations.
+In contrast, our approach, with adherence to the dimensional correspondence between $\mathbf{H}$ and $\mathbf{T}$, ensures that the learned features are more physically grounded and expressive, effectively capturing the underlying relationships between SO(3)-invariant and SO(3)-equivariant quantities.
+See Appendix H for an empirical study comparing between the gradient-based mechanism and the gated mechanism.
+Figure 1: The proposed framework for learning SO(3)-equivariant representations with strong non-linear expressiveness to regress the SO(3)-equivariant electronic-structure Hamiltonian.
 5 Method
@@
 1, taking $z$ in Eq.
-5 serving as the bridge between Theorem 1 and Theorem 2, we propose a general method for learning non-linear representations that satisfy the SO(3)-equivariance property outlined in Eq.
-4.
+(1) serving as the bridge between Theorem 4.1 and Theorem 4.2, we propose a method for learning expressive non-linear representations that satisfy the SO(3)-equivariance property outlined in Eq.
+(10).
 The core of our method can be abstractly referred to as TraceGrad.
-At the label level, it incorporates the SO(3)-invariant trace quantity $tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})$ introduced in Theorem 1 as a supervisory signal for learning $z$, i.e., the SO(3)-invariant internal representations of $g_{\text{nonlin}}(\cdot)$ in Theorem 2.
+At the label level, it incorporates the SO(3)-invariant trace quantity $tr(\mathbf{H}\cdot\mathbf{H}^{\dagger})$ introduced in Theorem 4.1 as a supervisory signal for learning $z$, i.e., the SO(3)-invariant internal representations of $g_{\text{nonlin}}(\cdot)$ in Theorem 4.2.
 Given the attribute that invariance is preserved under non-linear operations, $z$ is encoded by $s_{nonlin}(\cdot)$ to obtain non-linear expressive capabilities.
 At the representation level, by taking the gradient of $z$ with respect to $\mathbf{f}$, the non-linear expressiveness of $z$ is transferred to the equivariant feature $\mathbf{v}$, while maintaining strict SO(3)-equivariance as we prove.
-Subsequently, merging $\mathbf{v}$ and $\mathbf{f}$, and applying $g_{\text{nonlin}}(\cdot)$ in a stacked manner can yield rich SO(3)-equivariant non-linear representations for inferring the SO(3)-equivariant regression target.
+Subsequently, merging $\mathbf{v}$ and $\mathbf{f}$, and applying $g_{\text{nonlin}}(\cdot)$ in a stacked manner can yield rich SO(3)-equivariant non-linear representations for inferring the SO(3)-equivariant electronic-structure Hamiltonians.
 5.1 Encoding and Decoding Framework
@@
 Next, we construct the feature $\textbf{v}^{(k)}=g_{\text{nonlin}}(\textbf{f}^{(k)})$, to achieve sufficient non-linear expressiveness while maintaining SO(3)-equivariance, where $g_{\text{nonlin}}(\cdot)$, $s_{nonlin}(\cdot)$ are the non-linear functions defined in Eq.
-5.
+(1).
 The function $s_{nonlin}(\cdot)$ can be implemented as any differentiable non-linear neural network module, such as feed-forward layers with non-linear activation functions like $SiLU$ and normalization operations like $Layernorm$.
@@
 In this case, it becomes necessary to extend the decomposition operator, i.e., $\text{CGDecomp}(\cdot)$ in Eq.
-5, to accommodate various components of degrees.
+(1), to accommodate various components of degrees.
 Moreover, in the context of neural networks, introducing learnable parameters $W$ and more feature channels $C$ may improve the model capacity.
 Based on these considerations, when constructing the encoding module, the $\text{CGDecomp}(\cdot)$ operation in Eq.
-5 can be expanded as $\text{CGDecomp}_{ext}(\cdot)$, and its outputs can be expanded as $\textbf{u}^{(k)}$, as shown in Eq.
-6:
+(1) can be expanded as $\text{CGDecomp}_{ext}(\cdot)$, and its outputs can be expanded as $\textbf{u}^{(k)}$, as shown in Eq.
+(2):
 $\textbf{u}^{(k)}$ $=[u_{1}^{(k)},...,u_{c}^{(k)},...,u_{C}^{(k)}],$ (2)
-$u_{c}^{(k)}$ $=\text{CGDecomp}_{\text{ext}}(\textbf{f}^{(k)}\otimes\textbf{f}^{(k)},0,W)=\sum_{l_{i},l_{j}\in L^{(k)},l_{i}=l_{j}}W^{c}_{ij}\cdot\text{CGDecomp}(\textbf{f}^{(k)l_{i}}\otimes\textbf{f}^{(k)l_{j}},0)$
+$u_{c}^{(k)}$ $=\text{CGDecomp}_{\text{ext}}(\textbf{f}^{(k)}\otimes\textbf{f}^{(k)},0,W)$
+$=\sum_{l_{i},l_{j}\in L^{(k)},l_{i}=l_{j}}W^{c}_{ij}\cdot\text{CGDecomp}(\textbf{f}^{(k)l_{i}}\otimes\textbf{f}^{(k)l_{j}},0)$
 where $W^{c}_{ij}$ represents learnable parameters, $u_{c}^{(k)}$ is the $c$ th channel of $\mathbf{u}^{(k)}$.
@@
 $\mathbf{v}^{(k)}$ is combined with $\mathbf{f}^{(k)}$ in a residual manner like $\textbf{o}^{(k)}=\textbf{f}^{(k)}+\textbf{v}^{(k)}$ as the output of the $k$ th encoding module.
-[…]
+We follow previous literature (Gong et al., 2023; Yu et al., 2023b) to send the features from the last layer of the encoder, i.e., $\textbf{o}^{(K)}$ in our framework, into the SO(3)-equivariant decoder to regress the predictions of $\mathbf{H}$.
+For this, we can directly utilize the mature design of the SO(3)-equivariant decoders in DeepH-E3 or QHNet.
+The new part in the decoding phase introduced by our method is the SO(3)-invariant decoder, consisting of feed-forward layers taking $\textbf{z}=[\textbf{z}^{(1)},\ldots,\textbf{z}^{(k)},\ldots,\textbf{z}^{(K)}]$ as the input to predict $\mathbf{T}$.
 5.2 Training
 The training loss function is shown as the following:
-$\begin{split}\min_{\theta}loss&=loss_{Q}+\mu(loss_{Q},loss_{T})\cdot loss_{T},\\ loss_{Q}&=Error(\widehat{\mathbf{Q}},\mathbf{Q}^{*}),\,\,\,\,loss_{T}=Error(\widehat{\mathbf{T}},\mathbf{T}^{*})\end{split}$ (7)
-where $\theta$ denotes all of the learnable parameters of our framework, $\widehat{\mathbf{Q}}$, $\widehat{\mathbf{T}}$ and $\mathbf{Q}^{*}$, $\mathbf{T}^{*}$ respectively denote the predictions and labels of $\mathbf{Q}$ and $\mathbf{T}$.
-In order to prevent the numerical disparity between the two loss terms and stabilize the training for both of the SO(3)-equivariant and SO(3)-invariant branches, we apply $\mu(loss_{Q},loss_{T})$, i.e., a coefficient to regularize the relative scale of the two loss terms:
-$\mu(loss_{Q},loss_{T})=\lambda\cdot No\_Grad(\frac{loss_{Q}}{loss_{T}})$ (8)
+$\begin{split}\min_{\theta}loss&=loss_{H}+\mu(loss_{H},loss_{T})\cdot loss_{T},\\ loss_{H}&=Error(\widehat{\mathbf{H}},\mathbf{H}^{*}),\,\,\,\,loss_{T}=Error(\widehat{\mathbf{T}},\mathbf{T}^{*})\end{split}$ (3)
+where $\theta$ denotes all of the learnable parameters of our framework, $\widehat{\mathbf{H}}$, $\widehat{\mathbf{T}}$ and $\mathbf{H}^{*}$, $\mathbf{T}^{*}$ respectively denote the predictions and labels of $\mathbf{H}$ and $\mathbf{T}$.
+In order to prevent the numerical disparity between the two loss terms and stabilize the training for both of the SO(3)-equivariant and SO(3)-invariant branches, we apply $\mu(loss_{H},loss_{T})$, i.e., a coefficient to regularize the relative scale of the two loss terms:
+$\mu(loss_{H},loss_{T})=\lambda\cdot No\_Grad(\frac{loss_{H}}{loss_{T}})$ (4)
 where $No\_Grad(\cdot)$ denotes gradient discarding when calculating such coefficient, as this coefficient is only used for adjusting the relative scale between the two loss terms and should not itself be a source of training gradients, otherwise it would counteract the gradients from $loss_{T}$ in Eq.
-7.
+(3).
 All of the encoding and decoding modules are trained jointly by Eq.
-7.
-For the implementation details of our method, including the specific design of the network modules, parameter settings, and training specifics, please refer to Appendix D.
+(3).
 6 Experiments
@@
 The last two databases are from the QH9 benchmark series (Yu et al., 2023a), composed of molecular systems with elements like C, H, O, N and F.
-These databases present diverse and complex challenges to the generality of a regression model.
+These databases present diverse and complex challenges to a regression model.
 Regarding $MG$, $MM$, and QD, as their samples are prepared from an temperature environment at three-hundred Kelvin, the thermal motions lead to complex non-rigid deformations, increasing the difficulty of Hamiltonian prediction.
@@
 (2023a) is applied to split the training, validation, and testing sets, ensuring that there are no thermal motion samples from the same temporal trajectory across the three subsets.
-The ’mol’ and ’ood’ strategies aim to assess the regression model’s extrapolation capability with respect to the number of atoms as well as the temporal trajectories, respectively.
-Detailed statistic information of these databases can be found in the Appendix C.
-Implementation details of our method for experiments on these databases are presented in Appendix D.
+The ’mol’ and ’ood’ strategies aim to assess the regression model’s extrapolation capability with respect to the number of atoms and the temporal trajectories, respectively.
+Detailed statistic information of these databases can be found in the Appendix E.
+Implementation details of our method for experiments on these databases are presented in Appendix F.
 Table 2: Experimental results measured by the $MAE^{H}_{all}$, $MAE^{H}_{diag}$, $MAE^{H}_{non\_diag}$, $MAE^{\epsilon}$, and $Sim(\psi)$ metrics on the QH9-stable (QS) and QH9-dynamic (QD) databases respectively using ’ood’ and ’mol’ split strategies (Yu et al., 2023a).
@@
 QHNet+TraceGrad 2.819 6.844 2.497 63.375 0.927
-We use a comprehensive set of metrics to deeply evaluate the performance of deep learning electronic-structure Hamiltonian prediction models.
-The metrics can be broadly categorized into two major aspects:
-• Metrics on accuracy performance.
+We use a comprehensive set of metrics to deeply evaluate the accuracy performance of deep learning electronic-structure Hamiltonian prediction models.
 On the databases from the DeepH benchmark series (Li et al., 2022; Gong et al., 2023), we follow Yin et al.
@@
 $\epsilon$ and $\psi$ are crucial downstream physical quantities for determining multiple properties of the atomic systems as well as their dynamics, highly reflecting the application values of the Hamiltonian regression model.
-• Metrics on acceleration ratios for the convergence of traditional DFT algorithms.
-Despite the increasing ability of deep learning models to independently handle more electronic-structure computation tasks, there are still applications with extremely high numerical precision requirements and very low tolerance for error, where traditional DFT algorithms must perform the final calculations.
-In such cases, the predictions from deep models can be used as initial matrices to accelerate the convergence of traditional DFT algorithms.
-We evaluate the acceleration ratios brought by the proposed method for the convergence of classical DFT algorithms implemented by PySCF (Sun et al., 2018).
-Specifically, we adopt the two metrics defined in Yu et al.
-(2023a): the achieved ratio calculates the number of DFT optimization steps taken when initializing with the Hamiltonian matrices predicted by the deep model compared to using initial guess methods like minao and 1e.
-The error-level ratio measures the number of DFT optimization steps required, starting from random initialization, to reach the same error level as the deep model’s predictions, relative to the total number of steps in the DFT process.
-Table 3: The acceleration ratios of QHNet and QHNet+TraceGrad for DFT calculation.
-Both models are evaluated on a set of $50$ molecules chosen by Yu et al.
-(2023a), with the mean and standard deviation of the metrics across these samples reported.
-$\downarrow$ means lower values correspond to better accuracy, while $\uparrow$ means higher values correspond to better performance.
-Methods Training DFT Metric Ratio
-databases initialization
-QHNet (Baseline) QS 1e Achieved ratio $\downarrow$ $0.400\pm 0.030$
-Error-level ratio $\uparrow$ $0.620\pm 0.037$
-minao Achieved ratio $\downarrow$ $0.715\pm 0.033$
-Error-level ratio $\uparrow$ $0.406\pm 0.021$
-QD 1e Achieved ratio $\downarrow$ $0.512\pm 0.138$
-Error-level ratio $\uparrow$ $0.622\pm 0.048$
-minao Achieved ratio $\downarrow$ $0.882\pm 0.217$
-Error-level ratio $\uparrow$ $0.406\pm 0.066$
-QHNet+TraceGrad QS 1e Achieved ratio $\downarrow$ $\textbf{0.345}\pm 0.038$
-Error-level ratio $\uparrow$ $\textbf{0.685}\pm 0.037$
-minao Achieved ratio $\downarrow$ $\textbf{0.647}\pm 0.061$
-Error-level ratio $\uparrow$ $\textbf{0.466}\pm 0.035$
-QD 1e Achieved ratio $\downarrow$ $\textbf{0.440}\pm 0.101$
-Error-level ratio $\uparrow$ $\textbf{0.645}\pm 0.046$
-minao Achieved ratio $\downarrow$ $\textbf{0.761}\pm 0.167$
-Error-level ratio $\uparrow$ $\textbf{0.435}\pm 0.052$
 6.2 Results and Analysis
@@
 They are the respective state-of-the-art (SOTA) methods with strict SO(3)-equivariance on the corresponding databases.
-We list the results of DeepH-E3 and DeepH-E3+TraceGrad in Tables 1 for databases from the DeepH benchmark series, reporting the values of $MAE^{H}_{all}$, $MAE^{H}_{cha\_s}$, and $MAE^{H}_{cha\_b}$.
+We list the results of DeepH-E3 and DeepH-E3+TraceGrad in Table 1 for databases from the DeepH benchmark series, reporting the values of $MAE^{H}_{all}$, $MAE^{H}_{cha\_s}$, and $MAE^{H}_{cha\_b}$.
 The results of DeepH-E3 to be compared are copied from Yin et al.
@@
 The results of DeepH-E3+TraceGrad are the average from 10 independent repeated experiments.
-Regarding the metric of $MAE^{H}_{block}$ for every Hamiltonian block, due to its large data volume, we just present its values for all databases from the DeepH series in Appendix 4.
+Regarding the metric of $MAE^{H}_{block}$ for every Hamiltonian block, due to its large data volume, we just present its values for all databases from the DeepH series in Appendix G.
 We use the same fixed random seed as adopted by DeepH-E3 for all random processes in experiments on these six databases.
-As a result, the standard deviation of the Hamiltonian prediction MAE across repeated experiments does not exceed $0.007$ meV for each of the six databases and is negligible.
+As a result, for each Hamiltonian MAE result presented in Table 1, the standard deviation across independent repeated experiments does not exceed $0.007$ meV and is negligible.
 From results presented in Tables 1, we could find that the proposed TraceGrad method dramatically enhances the accuracy performance of the baseline method DeepH-E3, both on average and for challenging samples and blocks, both on the non-twisted samples and the twisted samples.
 Specifically, on the corresponding datasets, TraceGrad lowers down the $MAE^{H}_{all}$ and $MAE^{H}_{cha}$ of DeepH-E3 with relative ratios of up to $34\%$ and $35\%$, respectively.
-Furthermore, from the results included in Appendix 4, TraceGrad significantly improves the performance for the vast majority of basic blocks.
+Furthermore, from the results included in Appendix G, TraceGrad significantly improves the performance for the vast majority of basic blocks.
 Particularly, for the blocks where DeepH-E3 perform the worst, TraceGrad reduces the $MAE$ ($MAE^{H}_{cha\_b}$) by a maximum of $48\%$.
@@
 The outstanding performance on such samples highlights the good potential for studying twist-related phenomena, a hot research topic that may bring new electrical and transport properties (Cao et al., 2018; Wang et al., 2024a; He et al., 2024).
-Additionally, the $BG^{t}$, $BB^{t}$, $BT^{t}$, and $BS^{t}$ subsets contain significantly larger unit cells compared to the training set (see Appendix C for statistics of their sizes), yet our method still excels on these subsets as measured by the multiple MAE metrics, demonstrating its good scalability on the sizes of atomic systems it handles.
+Additionally, the $BG^{t}$, $BB^{t}$, $BT^{t}$, and $BS^{t}$ subsets contain significantly larger unit cells compared to the training set (see Appendix E for statistics of their sizes), yet our method still excels on these subsets as measured by the multiple MAE metrics, demonstrating its good scalability on the sizes of atomic systems it handles.
 In Table 2, we present the results of QHNet and QHNet+TraceGrad under the metrics of $MAE^{H}_{all}$, $MAE^{H}_{diag}$, $MAE^{H}_{non\_diag}$, $MAE^{\epsilon}$, and $Sim(\psi)$ for the $QS$ and $QD$ databases.
-The results of QHNet to be compared are taken from their original paper (Yu et al., 2023a), and for the unification of MAE units, we convert the units of MAE from $10^{-6}$ Hartree ($E_{h}$) in the original paper to meV111 $27211.41E_{h}=$ 27211.4 meV.
+The results of QHNet to be compared are taken from their original paper (Yu et al., 2023a), and for the unification of MAE units, we convert the units of MAE from $10^{-6}$ Hartree ($E_{h}$) in the original paper to meV ($27211.41E_{h}=$ 27211.4 meV).
 The results of QHNet+TraceGrad are the average from 10 independent repeated experiments.
 To ensure reproducibility, we use the same fixed random seeds as employed in QHNet for all random processes in the experiments on the $QS$ and $QD$ databases.
-As a result, the standard deviation of the Hamiltonian prediction MAE across repeated experiments is no greater than $0.009$ meV for both $QS$ and $QD$ and is also negligible.
+As a result, for each Hamiltonian MAE result presented in Table 2, the standard deviation across repeated experiments is no greater than $0.009$ meV and is also negligible.
 The results presented in Table 2 demonstrate that the proposed TraceGrad method significantly enhances the accuracy of the baseline QHNet model across all metrics on the $QS$ and $QD$ databases.
-Specifically, TraceGrad reduces $MAE^{H}_{all}$, $MAE^{H}_{non\_diag}$, $MAE^{H}_{diag}$, $MAE^{\epsilon}$, and $Sim(\psi)$ of QHNet with relative reductions of up to 40%, 39%, 40%, 76%, and 17%, respectively, on the corresponding databases.
+Specifically, TraceGrad reduces $MAE^{H}_{all}$, $MAE^{H}_{non\_diag}$, $MAE^{H}_{diag}$, $MAE^{\epsilon}$, and $Sim(\psi)$ of QHNet with relative reductions by a maximum of 40%, 39%, 40%, 76%, and 17%, respectively, on the corresponding databases.
 The significant accuracy improvements on the $QS$ database, partitioned using the ’ood’ split strategy (Yu et al., 2023a) without scale overlapping among the training, validation, and testing sets, once again demonstrate the method’s strong generalization capabilities across different scales of atomic systems.
@@
 Furthermore, the substantial improvement in the prediction accuracy of $\epsilon$, i.e., occupied orbital energies crucial for determining electronic properties such as optical characteristics and conductivity in atomic systems, and $\psi$, i.e., the electronic wavefunctions essential for understanding electron distribution and interactions, underscores the potential values of our method for applications like material design, molecular pharmacology, and quantum computing.
-We also compare the acceleration ratios brought by the QHNet model and the QHNet+TraceGrad model for the convergence of classical DFT algorithms implemented by PySCF (Sun et al., 2018), with the experimental results reported in Table 3, using the achieved ratio and error-level ratio metrics introduced in Section 6.1.
-As observed from this table, TraceGrad also brings significant improvements to the baseline model QHNet, notably reducing the achieved ratio and enhancing the error-level ratio, indicating its substantial potential for accelerating traditional DFT algorithms.
-Due to the page limits, we present three groups of supplementary experimental results in the Appendices.
-The first group, included in Appendix H, is the results of ablation study.
-These fine-grained experiments assess the individual and combined contributions of the two core mechanisms of the proposed TraceGrad method on the six databases from the DeepH series.
-These mechanisms are the SO(3)-invariant trace supervision mechanism (Trace) at the label level and the gradient-based induction mechanism (Grad) at the representation layer.
-Results indicate that each mechanism can contribute individually to the performance.
+We also evaluate the acceleration performance of our method for DFT computations and report the results in Appendix K .
+Experimental results show that while combining TraceGrad introduces only a slight increase from the inference time of QHNet, it delivers significant improvements in accelerating the convergence of DFT methods.
+7 Summary of Appendices
+Due to the page limit of the main manuscript, we have to provide some important contents in the Appendices, summarized as follows:
+Appendix A provides the definitions of foundational concepts relevant to the problem addressed in this paper; Appendix B describes the application task in this paper, i.e., the electronic-structure Hamiltonian prediction task; Appendix C formalizes Hamiltonian prediction as an SO(3)-equivariant non-linear representation learning problem; Appendix D provides the proofs for all of the proposed theorems; Appendix E presents the detailed information of the experimental databases; Appendix F presents the implementation details of the experiments; Appendix G compares the block-level MAE statistics ($MAE^{H}_{block}$ and $MAE^{H}_{cha\_b}$) for DeepH-E3 and DeepH-E3+TraceGrad; Appendix H reports the results of ablation study for each individual mechanism of our framework and the quantitative comparison between the proposed gradient-based mechanism (Grad) with the existing gated activation mechanism (Gate).
+Experimental results indicate that each individual mechanism of our method can contribute individually to the performance.
 Moreover, their combination provides even better performance.
-The second group of experiments, as included in Appendix I, test the extra time cost caused by TraceGrad compare to the baseline model.
-As we adopt a light-weight implementation of the non-linear modules introduced in the TraceGrad method, experimental results show that the extra time burden is slight relative to the baseline models DeepH-E3 and QHNet, indicating the good efficiency of our approach.
-The third group of experiments corresponds to the synergy of our method with an approximately SO(3)-equivariant methodology (Yin et al., 2024).
-As approximate equivariance is not the main focus of this paper, we just include these experimental results in Appendix J.
-The experimental results in Appendix J still demonstrate a significant accuracy improvement brought by our approach, further supporting the generality of our methodology.
+Experimental results also demonstrate better accuracy performance of the proposed gradient-based mechanism compared to the gated mechanism; Appendix I provides a theoretical analysis of the computational complexity advantage of our method over traditional DFT calculations; Appendix J makes a joint discussion on GPU time costs and performance gains brought by our TraceGrad method.
+It underscores the superiority of the TraceGrad method in improving accuracy performance while maintaining time efficiency; Appendix K reports the acceleration performance of our method for traditional DFT algorithms; Appendix L corresponds to the synergy of our method with an approximately SO(3)- equivariant methodology (Yin et al., 2024).
 8 Conclusion
-We propose a theoretical and methodological framework to tackle the issue of reconciling non-linear expressiveness with SO(3)-equivariance in deep learning frameworks for physical system modeling, through deeply investigating the mathematical connections between SO(3)-invariant and SO(3)-equivariant quantities, as well as their representations.
+We propose a theoretical and methodological framework to tackle the issue of reconciling powerful non-linear expressiveness with SO(3)-equivariance in deep learning paradigm for electronic-structure Hamiltonian prediction, through deeply investigating the mathematical connections between SO(3)-invariant and SO(3)-equivariant quantities, as well as their representations.
 We first constructs SO(3)-invariant quantities from SO(3)-equivariant regression targets, using them to train informative SO(3)-invariant non-linear representations.
-From these, SO(3)-equivariant features are derived with gradient operations, achieving non-linear expressiveness while maintaining strict SO(3)-equivariance.
-We apply our theory and method to the challenging electronic-structure Hamiltonian prediction tasks, achieving dramatic promotions in prediction accuracy across eight benchmark databases.
-Experimental results demonstrate that this approach not only improves the accuracy of Hamiltonian prediction but also significantly enhances the prediction for downstream physical quantities, and also markedly improves the acceleration ratios for traditional DFT algorithms.
-Ethics Statement
-This work develops a representation learning method that exhibits strong non-linear expressive capabilities while strictly adhering to SO(3) equivariance.
-This method has demonstrated superior accuracy in predicting electronic-structure Hamiltonians and related physical quantities, showcasing its potential to accelerate research in materials science and molecular pharmacology.
-While we recognize that our research area has not yet revealed direct negative social or ethical implications, several issues warrant our vigilance.
-Currently, although our method yields accurate predictions, the decision-making processes of deep learning systems often lack transparency, hindering a comprehensive understanding of the learning outcomes and limiting our ability to gain deeper insights.
-We believe it is important to investigate the interpretability of such models, particularly in terms of how they apply physical knowledge in a comprehensible way.
-Additionally, it is crucial to continually improve the correctness and fairness of deep learning models on this area.
-Ensuring high-quality and diverse training data, implementing sound model designs, and performing ongoing validation and refinement are necessary to guarantee model accuracy and the broad applicability of their results.
+From these, SO(3)-equivariant features are derived with gradient operations, achieving non-linear expressiveness while maintaining strict SO(3)-equivariance and physical dimensional conistency to the target labels.
+Experimental results on eight benchmark databases covering multiple types of systems and challenging scenarios show substantial improvements on the state-of-the-art prediction accuracy of deep learning paradigm on electronic-structure Hamiltonian prediction.
+Our method also significantly promotes the acceleration performance for the convergence of traditional DFT methods.
 Appendices
-A Background
-Definition 1.
-Group. A set $G$, denoted as $G=\{\dots,g,\dots\}$, equipped with a binary operation denoted as $\cdot$, is called a group if it satisfies the following four conditions:
+A Definition of Concepts
+This section provides definitions of foundational concepts relevant to the problem addressed in this paper. For additional background, please refer to the book by Dresselhaus et al. (2007).
+Definition .1.
+Group. A set $G$, denoted as $G=\{\dots,g,\dots\}$, equipped with a binary operation $\cdot$, is called a group if it satisfies the following four axioms:
   1. 1.
-    Closure: For any two elements $f,g\in G$, $f\cdot g\in G$.
+    Closure: For any $f,g\in G$, the result of their operation $f\cdot g$ is also an element of $G$: $f\cdot g\in G$.
   2. 2.
-    Associativity: For all elements $f,g,h\in G$, the equality $(f\cdot g)\cdot h=f\cdot(g\cdot h)$ holds.
+    Associativity: For all $f,g,h\in G$, the operation satisfies $(f\cdot g)\cdot h=f\cdot(g\cdot h)$.
   3. 3.
-    Existence of Identity Element: There exists a unique identity element $e\in G$ such that for all $f\in G$, $e\cdot f=f\cdot e=f$.
+    Identity Element: There exists a unique element $e\in G$ (called the identity) such that for all $f\in G$, $e\cdot f=f\cdot e=f$.
   4. 4.
-    Existence of Inverse Element: For each element $f\in G$, there exists a unique inverse element $f^{-1}\in G$ such that $f\cdot f^{-1}=f^{-1}\cdot f=e$.
-Definition 2.
-SO(3) Group. The special orthogonal group $\text{SO}(3)$ is the group of all $3\times 3$ orthogonal matrices with determinant 1. Formally, it can be defined as:
-$\text{SO}(3)=\{\mathbf{R}\in\mathbb{R}^{3\times 3}\mid\mathbf{R}^{T}\mathbf{R}=I,\,\det(\mathbf{R})=1\}$
-where $\mathbf{R}^{T}$ denotes the transpose of $\mathbf{R}$ and $I$ is the $3\times 3$ identity matrix. The elements of $\text{SO}(3)$ represent rotations in three-dimensional Euclidean space.
-Definition 3.
+    Inverse Element: For each $f\in G$, there exists a unique element $f^{-1}\in G$ (called the inverse) such that $f\cdot f^{-1}=f^{-1}\cdot f=e$.
+Definition .2.
+SO(3) Group. The special orthogonal group $\text{SO}(3)$ is the group of all $3\times 3$ orthogonal matrices with determinant 1. Formally, it is defined as:
+$\text{SO}(3)=\{\mathbf{R}\in\mathbb{R}^{3\times 3}\mid\mathbf{R}^{T}\mathbf{R}=I,\,\det(\mathbf{R})=1\},$
+where $\mathbf{R}^{T}$ denotes the transpose of $\mathbf{R}$, and $I$ is the $3\times 3$ identity matrix. The elements of $\text{SO}(3)$ represent rotations in three-dimensional Euclidean space.
+Definition .3.
 Group Representation.
@@
 Here, $T(V)$ represents the tensor space associated with the vector space $V$, encompassing all tensors that can be formed from elements of $V$.
-The homomorphism $\rho$ can be formalized as:
+Formally, the homomorphism $\rho$ is defined as:
 $\rho:G\to GL(T(V))$
 such that for all $g_{1},g_{2}\in G$,
-$\rho(g_{1}g_{2})=\rho(g_{1})\rho(g_{2})$
+$\rho(g_{1}g_{2})=\rho(g_{1})\rho(g_{2}),$
 and $\rho(e)=I$, where $e$ is the identity element of $G$, and $I$ is the identity transformation on $T(V)$.
-Definition 4.
-SO(3) Group Representation.
-A representation of the special orthogonal group $\text{SO}(3)$ on a tensor space $T(V)$ is a homomorphism $\rho:\text{SO}(3)\to GL(T(V))$, where $GL(T(V))$ denotes the group of all invertible linear transformations on the tensor space $T(V)$.
-This representation can be expressed using Wigner-D matrices $\mathbf{D}^{l}(\mathbf{R})$, which are defined for a given degree $l$ corresponding to the representation of the rotation matrix $\mathbf{R}\in\text{SO}(3)$:
-$\rho(\mathbf{R})=\mathbf{D}^{l}(\mathbf{R}).$
-Definition 5.
+Definition .4.
+Irreducible Representation.
+A representation $\rho:G\to GL(V)$ of a group $G$ on a vector space $V$ is said to be irreducible if there is no proper subspace $W\subset V$ such that $\rho(g)W\subset W$ for all $g\in G$.
+In other words, the only invariant subspaces under the group action are the trivial subspaces $\{0\}$ and $V$ itself.
+If such a nontrivial invariant subspace exists, the representation is said to be reducible.
+Definition .5.
+SO(3) Group Representation. A representation of the special orthogonal group $\text{SO}(3)$ on a vector space $V$ is a homomorphism
+$\rho:\text{SO}(3)\to GL(V),$
+where $GL(V)$ denotes the group of all invertible linear transformations on $V$.
+The irreducible representation of $\text{SO}(3)$, each labeled by a degree $l$, often corresponds to the quantum number for angular momentum in quantum physics.
+Definition .6.
+Wigner-D Matrices.
+One of the irreducible representations of the rotation group $\text{SO}(3)$ is given by the Wigner-D matrices $\mathbf{D}^{l}(\mathbf{R})$, where $l$ is the degree of the representation, typically associated with angular momentum quantum number in quantum mechanics.
+The Wigner-D matrices are defined as:
+$D^{l}_{m^{\prime}m}(\mathbf{R})=\langle l,m^{\prime}|\mathbf{R}|l,m\rangle,$
+where $|l,m\rangle$ are the eigenstates of the angular momentum operator $\hat{L}^{2}$ and its $z$-component $\hat{L}_{z}$, with $m$ and $m^{\prime}$ being the magnetic quantum numbers corresponding to the eigenvalue of $\hat{L}_{z}$.
+The Wigner-D matrices $\mathbf{D}^{l}(\mathbf{R})$ encode how the angular momentum states $|l,m\rangle$ transform under the action of a rotation.
+Definition .7.
 Equivariance with Respect to a Group.
@@
 where $\circ$ generally denotes the operation defined on the tensor space.
-Definition 6.
+Definition .8.
 Invariance with Respect to a Group.
@@
 This definition indicates that the function $f$ remains unchanged under the action of the group $G$.
-B Proofs of Theorems
+Definition .9.
+Direct-Product State.
+Let $V_{1}$ and $V_{2}$ be two vector spaces, and let $|v_{1}\rangle\in V_{1}$ and $|v_{2}\rangle\in V_{2}$ be arbitrary elements of these spaces.
+The direct-product state of $|v_{1}\rangle$ and $|v_{2}\rangle$ is defined as the element $|v_{1}\rangle\otimes|v_{2}\rangle$ in the direct-product space $V_{1}\otimes V_{2}$.
+The direct-product state represents all possible combinations of the elements of $V_{1}$ and $V_{2}$, and forms a new vector space that captures the joint state of two systems.
+The action of a group on the direct-product state is defined by the direct product of the individual actions on $|v_{1}\rangle$ and $|v_{2}\rangle$, i.e., the effect of the group action on the direct-product space is the direct product of the actions on each individual space:
+$g\cdot(|v_{1}\rangle\otimes|v_{2}\rangle)=(g\cdot|v_{1}\rangle)\otimes(g\cdot|v_{2}\rangle),\quad\forall g\in G.$
+Definition .10.
+Physical Quantity Formed by the Direct-Product of Two Degrees.
+In quantum mechanics, direct-product spaces are used to describe the combined states of systems with distinct degrees of freedom, such as angular momentum.
+The combined state captures both the independent action of each degree and their joint transformation under rotations governed by the $\text{SO}(3)$ group.
+Let degrees $l_{p}$ and $l_{q}$ represent the angular momentum quantum numbers of two systems.
+A direct-product state $\mathbf{Q}^{l_{p}\otimes l_{q}}\in\mathbb{R}^{(2l_{p}+1)\times(2l_{q}+1)}$ combines these degrees and transforms under the $\text{SO}(3)$ group according to:
+$\mathbf{Q}(\mathbf{R})^{l_{p}\otimes l_{q}}=\mathbf{D}^{l_{p}}(\mathbf{R})\cdot\mathbf{Q}^{l_{p}\otimes l_{q}}\cdot(\mathbf{D}^{l_{q}}(\mathbf{R}))^{\dagger},$
+where $\mathbf{D}^{l_{p}}(\mathbf{R})\in\mathbb{R}^{(2l_{p}+1)\times(2l_{p}+1)}$ and $\mathbf{D}^{l_{q}}(\mathbf{R})\in\mathbb{R}^{(2l_{q}+1)\times(2l_{q}+1)}$ are the Wigner-D matrices for the degrees $l_{p}$ and $l_{q}$, respectively, and $\dagger$ denotes the conjugate transpose, ensuring unitary transformations.
+Definition .11.
+Direct-Sum State.
+Let $V_{1}$ and $V_{2}$ be two vector spaces, and let $|v_{1}\rangle\in V_{1}$ and $|v_{2}\rangle\in V_{2}$ be arbitrary elements of these spaces.
+The direct-sum state of $|v_{1}\rangle$ and $|v_{2}\rangle$ is defined as the element $|v_{1}\rangle\oplus|v_{2}\rangle$ in the direct-sum space $V_{1}\oplus V_{2}$.
+The direct-sum space $V_{1}\oplus V_{2}$ consists of ordered pairs of elements, where each element is drawn from one of the original spaces.
+The operations of vector addition and scalar multiplication in $V_{1}\oplus V_{2}$ are defined component-wise:
+$(a_{1},a_{2})+(b_{1},b_{2})=(a_{1}+b_{1},a_{2}+b_{2}),\quad c\cdot(a_{1},a_{2})=(c\cdot a_{1},c\cdot a_{2}),$
+for $(a_{1},a_{2}),(b_{1},b_{2})\in V_{1}\oplus V_{2}$ and $c\in\mathbb{F}$, where $\mathbb{F}$ is a field, typically either the real numbers $\mathbb{R}$ or the complex numbers $\mathbb{C}$.
+The direct-sum state $|v_{1}\rangle\oplus|v_{2}\rangle$ represents a combination where the components remain in their respective vector spaces and do not interact with each other.
+The action of a group $G$ on the direct-sum state is defined by its independent action on each component:
+$g\cdot(|v_{1}\rangle\oplus|v_{2}\rangle)=(g\cdot|v_{1}\rangle)\oplus(g\cdot|v_{2}\rangle),\quad\forall g\in G.$
+Definition .12.
+Clebsch-Gordan Decomposition for $\text{SO}(3)$ Group.
+In quantum mechanics, the Clebsch-Gordan decomposition is used to describe the coupling and decoupling of two angular momentum systems.
+For two angular momentum quantum numbers $l_{p}$ and $l_{q}$, their combined state can be decomposed into series components in the direct-sum state.
+Taking $\mathbf{Q}^{l_{p}\otimes l_{q}}$ defined in Definition .10 as an example, it can be decomposed as follows:
+$CGDecomp(\mathbf{Q}^{l_{p}\otimes l_{q}})=\bigoplus_{l=|l_{p}-l_{q}|}^{l_{p}+l_{q}}\mathbf{q}^{l},\,\,\,q_{m}^{l}=\sum_{m_{p},m_{q}}C_{m,m_{p},m_{q}}^{l,l_{p},l_{q}}Q^{l_{p}\otimes l_{q}}_{m_{p},m_{q}}$
+where $CGDecomp(\cdot)$ denotes the Clebsch-Gordan decomposition operator, $Q^{l_{p}\otimes l_{q}}_{m_{p},m_{q}}$ are the elements of the quantity $\mathbf{Q}^{l_{p}\otimes l_{q}}$ in the direct-product state, $C_{m,m_{p},m_{q}}^{l,l_{p},l_{q}}$ are the Clebsch-Gordan coefficients, and $\mathbf{q}^{l}$ denotes the decomposed results, with $q_{m}^{l}$ representing the component under quantum numbers $l$ and $m$.
+Figure 2: Illustration of atomic pairwise Hamiltonian matrices partitioned from the complete Hamiltonian matrix of the whole system.
+Each Hamiltonian matrix contains multiple basic blocks, denoted as $\mathbf{H}^{l^{i}_{p}\otimes l^{j}_{q}}$.
+B Application Task Description: Electronic-structure Hamiltonian Calculation
+Density Functional Theory (DFT) (Hohenberg & Kohn, 1964; Kohn & Sham, 1965) has become a cornerstone of modern electronic structure theory, playing a pivotal role in condensed matter physics, quantum chemistry, and materials science.
+Introduced in the 1960s through the foundational work of Hohenberg, Kohn, and Sham, DFT provides a framework for studying many-electron systems by replacing the computationally expensive many-body wavefunction with the electron density $\rho(\mathbf{r})$ as the fundamental variable.
+This reformulation significantly reduces computational complexity while preserving essential quantum mechanical effects, enabling researchers to investigate systems of practical interest with manageable computational resources.
+Over the decades, DFT has proven to be instrumental in calculating orbital energies and band structures, optimizing structural geometries, and exploring a wide range of material properties, underscoring its versatility and importance across diverse scientific disciplines.
+At the heart of DFT lies the Kohn-Sham equation (Kohn & Sham, 1965), which simplifies the many-body problem into a set of single-particle equations. These equations are expressed as:
+$\hat{H}\psi_{i}(\mathbf{r})=\epsilon_{i}\psi_{i}(\mathbf{r}),\quad\text{subject to}\quad\hat{H}=-\frac{\hbar^{2}}{2m}\nabla^{2}+V_{\text{ext}}(\mathbf{r})+V_{\text{HXC}}[\rho](\mathbf{r})$ (5)
+where $\hat{H}$ is the Hamiltonian operator, the Hartree-exchange-correlation potential $V_{\text{HXC}}[\rho](\mathbf{r})=V_{\text{H}}[\rho](\mathbf{r})+V_{\text{XC}}[\rho](\mathbf{r})$ is a functional of the electron density $\rho(\mathbf{r})$.
+The electron density is obtained from the Kohn-Sham orbitals $\psi_{i}(\mathbf{r})$ as:
+$\rho(\mathbf{r})=\sum_{i=1}^{M}|\psi_{i}(\mathbf{r})|^{2},$
+where $M$ represents the total number of electrons.
+Atomic orbitals offer a computationally efficient basis for electronic structure calculations because they require fewer basis functions compared to other types of basis sets (Lin et al., 2023).
+Additionally, their inherently localized nature makes them particularly advantageous for large systems.
+When expressed in an atomic orbital basis set, the Kohn-Sham equations can be formulated into as a generalized eigenvalue problem:
+$\mathbf{H}\mathbf{C}=\mathbf{\epsilon}\mathbf{S}\mathbf{C},$ (6)
+where $\mathbf{H}$ is the Hamiltonian matrix incorporating the contributions from $V_{\text{ext}}$ and $V_{\text{HXC}}$, $\mathbf{S}$ is the overlap matrix, $\mathbf{C}$ contains the orbital coefficients, and $\mathbf{\epsilon}$ represents the eigenvalues of the system.
+The atomic orbitals used as the basis functions typically take the form:
+$\phi_{\mu}(\mathbf{r})=R_{\mu}(r)Y_{l_{\mu}m_{\mu}}(\theta,\phi),$
+where $R_{\mu}(r)$ is the radial part of the wavefunction, and $Y_{l_{\mu}m_{\mu}}(\theta,\phi)$ are the spherical harmonics describing the angular dependence.
+These orbitals are localized around their respective atomic centers, decay rapidly as $r$ increases, and are truncated to zero beyond a cutoff radius $r_{c}$.
+The matrix elements of the Hamiltonian and overlap matrices, $H_{\mu\nu}$ and $S_{\mu\nu}$, are defined as:
+[…]
+The localized nature of atomic orbitals ensures that matrix elements are non-zero only when the orbitals $\phi_{\mu}$ and $\phi_{\nu}$ overlap according to their cutoff radius.
+Beyond this range, the corresponding matrix elements are treated as zero, leading to sparsity in the Hamiltonian and overlap matrices.
+This sparsity significantly reduces computational effort and memory usage, especially for large systems.
+Solving $H_{\mu\nu}$ involves an iterative self-consistent process: starting with an initial guess for $\rho^{(1)}(\mathbf{r})$, the potentials $V_{\text{HXC}}[\rho](\mathbf{r})$ are updated, and the equations are solved repeatedly until $\rho(\mathbf{r})$ converges.
+This process can be illustrated as $\rho^{(1)}(\mathbf{r})\rightarrow V^{(1)}_{\text{HXC}}[\rho](\mathbf{r})\rightarrow H^{(1)}_{\mu\nu}\rightarrow\psi^{(1)}_{\mu}(\mathbf{r})\rightarrow\rho^{(2)}(\mathbf{r})\rightarrow\ldots\rightarrow\rho^{(T)}(\mathbf{r})$, until the charge density $\rho(\mathbf{r})$ converges.
+At that point, the Hamiltonian matrix is outputted by $\rho^{(T)}(\mathbf{r})\rightarrow H^{(T)}_{\mu\nu}$, from which downstream physical properties such as orbital energies and band structures are induced, determining the electronic, magnetic, and transport characteristics of the electron system.
+Despite the remarkable success of Kohn-Sham DFT in advancing fields such as materials science, energy, and biomedicine over recent decades (Nagy, 1998; Jones, 2015), the challenge of high computational complexity remains unresolved.
+The main computational bottleneck lies in iterative solving the Kohn-Sham equation, especially in the matrix diagonalization of Eq.
+(6), which has a complexity of $\mathcal{O}(N^{3})$, where $N$ is the number of atoms in the system.
+To address this challenge, recent approaches (Li et al., 2022; Yu et al., 2023b; Gong et al., 2023) have applied the deep graph learning paradigm to predict the self-consistent Hamiltonian.
+These methods use the Hamiltonian matrix $\mathbf{H}^{(T)}$, calculated from traditional DFT methods, as labels.
+This matrix can be partitioned into a series of atomic pairwise Hamiltonian matrices $\{\mathbf{H}^{ij}\mid j\in\Omega(i)\}$, as shown in Fig.
+2, where $i$ and $j$ represent two atoms in the system.
+Assuming that the angular momentum quantum numbers of the orbitals of atom $i$ are $l^{i}_{p}\in\{l^{i}_{1},l^{i}_{2},...,l^{i}_{P}\}$, and those of atom $j$ are $l^{j}_{q}\in\{l^{j}_{1},l^{j}_{2},...,l^{j}_{Q}\}$, the basic block of $\mathbf{H}^{ij}$ can be expressed as ${\bf{H}}^{l^{i}_{p}\otimes l^{j}_{q}}$, which is formed by the direct product between the $l^{i}_{p}$ and $l^{j}_{q}$.
+These methods train an efficient graph neural network to directly predict each $\mathbf{H}^{ij}$ from the 3D structure of the atomic system, thereby circumventing the extremely time-consuming self-consistent iterations.
+During the inference phase, these methods successfully reduced the computational complexity of calculating the Hamiltonian matrices to $\mathcal{O}(N)$, while showing good potential in generalizing to larger atomic systems, even though the training set, constrained by the expensive DFT labels, only includes smaller systems.
+Once the Hamiltonian matrices are obtained, many downstream physical properties can be efficiently calculated with $\mathcal{O}(N)$ complexity.
+A detailed analysis of the computational complexity is provided in Appendix I.
+This paper aims to tackle the challenge of predicting electronic Hamiltonians with high accuracy and reliability, while rigorously maintaining SO(3)-equivariance.
+C Formulation of Hamiltonian Prediction as an SO(3)-Equivariant Non-linear Representation Learning Problem
+As shown in Fig.
+2, a basic block of Hamiltonian can be denoted as ${\bf{H}}^{l^{i}_{p}\otimes l^{j}_{q}}$, which is formed by the direct product between degrees $l^{i}_{p}$ and $l^{j}_{q}$.
+It obeys the following SO(3)-equivariant law:
+${\bf{H}}({\bf{R}})^{l^{i}_{p}\otimes l^{j}_{q}}={\bf{D}}^{l^{i}_{p}}({\bf{R}})\cdot{\bf{H}}^{l^{i}_{p}\otimes l^{j}_{q}}\cdot{({\bf{D}}^{l^{j}_{q}}({\bf{R}}))^{\dagger}}$ (7)
+where $\dagger$ denotes the conjugate transpose operation.
+${\bf{R}}\in\mathbb{R}^{3\times 3}$ is the rotational matrix, ${\bf{D}}^{l^{i}_{p}}({\bf{R}})\in\mathbb{R}^{(2l^{i}_{p}+1)\times(2l^{i}_{p}+1)}$ and ${\bf{D}}^{l^{j}_{q}}({\bf{R}})\in\mathbb{R}^{(2l^{j}_{q}+1)\times(2l^{j}_{q}+1)}$ are the Wigner-D matrices of degrees $l^{i}_{p}$ and $l^{j}_{q}$, respectively; ${\bf{H}}({\bf{R}})^{l^{i}_{p}\otimes l^{j}_{q}}\in\mathbb{R}^{(2l^{i}_{p}+1)\times(2l^{j}_{q}+1)}$ denotes the transformed results of ${\bf{H}}^{l^{i}_{p}\otimes l^{j}_{q}}\in\mathbb{R}^{(2l^{i}_{p}+1)\times(2l^{j}_{q}+1)}$ through the rotational operation by ${\bf{R}}$.
+${\bf{H}}^{l^{i}_{p}\otimes l^{j}_{q}}$ in the direct-product state can be further decomposed into a series of direct-sum state components, i.e., ${\bf{h}}^{l}(|l^{i}_{p}-l^{j}_{q}|\leq l\leq l^{i}_{p}+l^{j}_{q})$, which follows SO(3)-equivariant law mathematically equivalent to Eq.
+(7) but with a simpler form:
+${\bf{h}}({\bf{R}})^{l}={\bf{D}}^{l}({\bf{R}})\cdot{\bf{h}}^{l},|l^{i}_{p}-l^{j}_{q}|\leq l\leq l^{i}_{p}+l^{j}_{q}$ (8)
+where ${\bf{h}}^{l}\in\mathbb{R}^{2l+1}$ and ${\bf{h}}({\bf{R}})^{l}\in\mathbb{R}^{2l+1}$ respectively denote the components with degree $l$ before and after the rotational operation by ${\bf{R}}$.
+For ease of processing, the internal representations of SO(3)-equivariant neural networks (Gong et al., 2023; Liao & Smidt, 2023) are typically in the direct-sum form.
+To obey the equivariant law of Eq.
+(8) for regressing Hamiltonian, these hidden representations must also satisfy the same form of equivariance:
+${\bf{f}}({\bf{R}})^{(k)l}={\bf{D}}^{l}({\bf{R}})\cdot{\bf{f}}^{(k)l}$ (9)
+where ${\bf{f}}^{(k)l}\in\mathbb{R}^{2l+1}$ and ${\bf{f}}({\bf{R}})^{(k)l}\in\mathbb{R}^{2l+1}$ respectively denote one channel of hidden representations with degree $l$ before and after the rotational operation by ${\bf{R}}$, at the $k$ th hidden layer.
+Due to the intrinsic complexity and non-linearity of Hamiltonians, the regression neural networks are supposed to equip with powerful non-linear mappings as feature encoders to fully capture the intrinsic patterns of the Hamiltonians, which is crucial for precise and generalizable prediction performance.
+Meanwhile, the non-linear mappings, denoted as $g_{nonlin}(\cdot)$, must also preserve SO(3)-equivariance, which is expressed as:
+${\bf{f}}({\bf{R}})^{(k+1)l}={\bf{D}}^{l}({\bf{R}})\cdot{\bf{f}}^{(k+1)l},\,\,\,\text{subject to}\,\,{\bf{f}}^{(k+1)l}=g_{nonlin}({\bf{f}}^{(k)l})$ (10)
+However, directly implementing $g_{nonlin}(\cdot)$ as neural network module with non-linear activation functions, such as $Sigmoid$, $Softmax$ and $SiLU$, may result in the destruction of strict equivariance.
+How to make $g_{nonlin}(\cdot)$ both theoretically SO(3)-equivariant and capable of powerful non-linear expressiveness, and effectively apply it to the prediction of SO(3)-equivariant complex physical quantities, i.e., electronic-structure Hamiltonians in this context, is the core problem this paper aims at solving.
+D Proofs of Theorems
 Proof of Theorem 1.
-Under an SO(3) rotation represented by the rotational matrix $\mathbf{R}$, $\mathbf{Q}=\mathbf{Q}^{l_{p}\otimes l_{q}}$ is transformed as $\mathbf{Q}(\mathbf{R})$:
-$\mathbf{Q}(\mathbf{R})=\mathbf{D}^{l_{p}}(\mathbf{R})\cdot\mathbf{Q}\cdot\mathbf{D}^{l_{q}}(\mathbf{R})^{\dagger},$
-where $\mathbf{D}^{l_{p}}(\mathbf{R})$ and $\mathbf{D}^{l_{q}}(\mathbf{R})$ are the Wigner-D matrices for the degrees of $l_{p}$ and $l_{q}$, respectively, corresponding to the rotation $\mathbf{R}$.
+Under an SO(3) rotation represented by the rotational matrix $\mathbf{R}$, $\mathbf{H}=\mathbf{H}^{l^{i}_{p}\otimes l^{j}_{q}}$ is transformed as $\mathbf{H}(\mathbf{R})$:
+$\mathbf{H}(\mathbf{R})=\mathbf{D}^{l^{i}_{p}}(\mathbf{R})\cdot\mathbf{H}\cdot\mathbf{D}^{l^{j}_{q}}(\mathbf{R})^{\dagger},$
+where $\mathbf{D}^{l^{i}_{p}}(\mathbf{R})$ and $\mathbf{D}^{l^{j}_{q}}(\mathbf{R})$ are the Wigner-D matrices for the degrees of $l^{i}_{p}$ and $l^{j}_{q}$, respectively, corresponding to the rotation $\mathbf{R}$.
 The conjugate transpose of the transformed quantity is:
-$\mathbf{Q}(\mathbf{R})^{\dagger}=\mathbf{D}^{l_{q}}(\mathbf{R})\cdot\mathbf{Q}^{\dagger}\cdot\mathbf{D}^{l_{p}}(\mathbf{R})^{\dagger}.$
-Using the cyclic property of the trace, which states that the trace of a product of matrices remains unchanged under cyclic permutations (i.e., $tr(ABC)=tr(BCA)=tr(CAB)$), and combining the properties that $\mathbf{D}^{l_{p}}(\mathbf{R})\cdot\mathbf{D}^{l_{p}}(\mathbf{R})^{\dagger}=\mathbf{I}$ and $\mathbf{D}^{l_{q}}(\mathbf{R})\cdot\mathbf{D}^{l_{q}}(\mathbf{R})^{\dagger}=\mathbf{I}$, we can rearrange the terms inside the trace as follows:
-$\mathbf{T}(\mathbf{R})=tr(\mathbf{Q}(\mathbf{R})\cdot\mathbf{Q}(\mathbf{R})^{\dagger})=tr((\mathbf{D}^{l_{p}}(\mathbf{R})\cdot\mathbf{Q}\cdot\mathbf{D}^{l_{q}}(\mathbf{R})^{\dagger})\cdot(\mathbf{D}^{l_{q}}(\mathbf{R})\cdot\mathbf{Q}^{\dagger}\cdot\mathbf{D}^{l_{p}}(\mathbf{R})^{\dagger}))$
-$=tr(\mathbf{D}^{l_{p}}(\mathbf{R})\cdot\mathbf{Q}\cdot\mathbf{Q}^{\dagger}\cdot\mathbf{D}^{l_{p}}(\mathbf{R})^{\dagger})=tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger}\cdot\mathbf{D}^{l_{p}}(\mathbf{R})^{\dagger}\cdot\mathbf{D}^{l_{p}}(\mathbf{R}))=tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})=\mathbf{T}.$
-Therefore, $\mathbf{T}=tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})$ is invariant under SO(3) transformations, its SO(3)-invariance is proved. ∎
+$\mathbf{H}(\mathbf{R})^{\dagger}=\mathbf{D}^{l^{j}_{q}}(\mathbf{R})\cdot\mathbf{H}^{\dagger}\cdot\mathbf{D}^{l^{i}_{p}}(\mathbf{R})^{\dagger}.$
+Using the cyclic property of the trace, which states that the trace of a product of matrices remains unchanged under cyclic permutations (i.e., $tr(ABC)=tr(BCA)=tr(CAB)$), and combining the properties that $\mathbf{D}^{l^{i}_{p}}(\mathbf{R})\cdot\mathbf{D}^{l^{i}_{p}}(\mathbf{R})^{\dagger}=\mathbf{I}$ and $\mathbf{D}^{l^{j}_{q}}(\mathbf{R})\cdot\mathbf{D}^{l^{j}_{q}}(\mathbf{R})^{\dagger}=\mathbf{I}$, we can rearrange the terms inside the trace as follows:
+$\mathbf{T}(\mathbf{R})=tr(\mathbf{H}(\mathbf{R})\cdot\mathbf{H}(\mathbf{R})^{\dagger})=tr((\mathbf{D}^{l^{i}_{p}}(\mathbf{R})\cdot\mathbf{H}\cdot\mathbf{D}^{l^{j}_{q}}(\mathbf{R})^{\dagger})\cdot(\mathbf{D}^{l^{j}_{q}}(\mathbf{R})\cdot\mathbf{H}^{\dagger}\cdot\mathbf{D}^{l^{i}_{p}}(\mathbf{R})^{\dagger}))$
+$=tr(\mathbf{D}^{l^{i}_{p}}(\mathbf{R})\cdot\mathbf{H}\cdot\mathbf{H}^{\dagger}\cdot\mathbf{D}^{l^{i}_{p}}(\mathbf{R})^{\dagger})=tr(\mathbf{H}\cdot\mathbf{H}^{\dagger}\cdot\mathbf{D}^{l^{i}_{p}}(\mathbf{R})^{\dagger}\cdot\mathbf{D}^{l^{i}_{p}}(\mathbf{R}))=tr(\mathbf{H}\cdot\mathbf{H}^{\dagger})=\mathbf{T}.$
+Therefore, $\mathbf{T}=tr(\mathbf{H}\cdot\mathbf{H}^{\dagger})$ is invariant under SO(3) transformations, its SO(3)-invariance is proved. ∎
 Proof of Theorem 2.
@@
 ∎
-C Information of Experimental Databases
-Figure 2: Illustration of a Hamiltonian matrix.
-Each Hamiltonian matrix contains multiple basic blocks, with the $\mathbf{Q}=\mathbf{Q}^{l_{p}\otimes l_{q}}$ defined in Section 3 corresponding to a basic block $\mathbf{H}^{l_{p}\otimes l_{q}}$ here.
+E Information of Experimental Databases
 In this part, we provide detailed information about the experimental databases, including the statistical information of the six databases from the DeepH benchmark series (Li et al., 2022; Gong et al., 2023) and the two databases from the QH9 benchmark series (Yu et al., 2023a), listed in Table 3 and Table 4, respectively.
@@
 4, respectively.
-D Implementation Details
+F Implementation Details
 The hardware environment for our experiments is a server cluster equipped with Nvidia RTX A6000 GPUs, each with 48 GiB of memory.
 Other experimental details may differ across the DeepH and QH9 benchmark series, which we will describe separately.
-E Implementation Details on the DeepH Benchmark Series
+F.1 Implementation Details on the DeepH Benchmark Series
 The software environment used is Pytorch 2.0.1 for experiments on the six crystalline databases from the DeepH benchmark series.
 When combining the proposed TraceGrad method with the DeepH-E3 architecture, the implementation of DeepH-E3 is based on the project  [^2] https://github.com/Xiaoxun-Gong/DeepH-E3 provided by Gong et al.
-(2023), keeping the architecture and hyperparameter configurations consistent with their setup.
+(2023), keeping the architecture and model hyperparameters consistent with their setup.
 In our framework, we use the same number of encoding modules $K$ as DeepH-E3, which is set to $3$.
 In each encoding module, we apply the $g_{nonlin}(\cdot)$ module proposed in Section 4 and 5 to each SO(3)-equivariant edge feature, enabling non-linear expressiveness.
-We set the number of channels for the rotationally SO(3)-invariant feature $\textbf{u}^{(k)}$ $(1\leq k\leq 3)$ to $1024$.
+We set the number of channels for the SO(3)-invariant feature $\textbf{u}^{(k)}$ $(1\leq k\leq 3)$ to $1024$.
 The neural network module $s_{nonlin}(\cdot)$ within $g_{nonlin}(\cdot)$ is implemented as a three-layers fully-connected module: the input size is set to $1024$, consistent with $\textbf{u}^{(k)}$; the hidden layer size is also $1024$, with SiLU as the non-linear activation function and LayerNorm as the normalization mechanism; and the output layer size (i.e., the dimensionality of $\textbf{z}^{(k)}$) is set to be equal to the number of basic blocks for a Hamiltonian matrix, which is $25$ for $MG$ and $BG$, $49$ for $MM$, and $196$ for $BB$, $BT$, and $BS$.
@@
 Its role is focusing on to filling in the gaps left by the existing equivariant mechanisms in DeepH-E3: to introduce a non-linear mapping mechanism that maintains equivariance, thereby activating and unleashing the expressive power of the overall network architecture through non-linearity.
-The SO(3)-equivariant decoder we adopt is the same as that of DeepH-E3; The SO(3)-invariant decoder we adopt is a four-layers fully-connected module: the input size is $3$ $(K)$ times of the dimensionality of $\textbf{z}^{k}$, e.g., $75$ for $MG$; the hidden layers have $1024$ neurons with SiLU as the non-linear activation function and LayerNorm as the normalization mechanism; the size of the output layer is the number of basic blocks for a Hamiltonian matrix.
-Since each basic block of the Hamiltonian matrix can compute a trace, the total number of trace variables corresponds to the number of basic blocks.
+The SO(3)-equivariant decoder we adopt is the same as that of DeepH-E3; The SO(3)-invariant decoder we adopt is a four-layers fully-connected module: the input size is $3$ $(K)$ times of the dimensionality of $\textbf{z}^{k}$, e.g., $75$ for $MG$; the hidden layers have $1024$ neurons with SiLU as the non-linear activation function and LayerNorm as the normalization mechanism; the size of the output layer is the number of basic blocks for an atomic pairwise Hamiltonian matrix.
+Since each basic block of the Hamiltonian matrix can compute a trace, the total number of trace variables corresponds to the number of basic matrix blocks.
 Regarding the error metric in the loss function Eq.
-7, for the first term, we follow DeepH-E3 to use MSE (Mean Squared Error); for the second term, we choose between MSE and MAE based on performance on the validation sets, ultimately selecting MAE.
+(3), for the first term, we follow DeepH-E3 to use MSE (Mean Squared Error); for the second term, we choose between MSE and MAE based on performance on the validation sets, ultimately selecting MAE.
 $\lambda$ in the training loss function is set according to parameter selection on the validation sets, searching from $\{0.1,0.2,...,1.0\}$.
 Here, we aim to obtain a more general parameter setting for $\lambda$ on crystalline structures, and thus we determined $\lambda$ based on the overall performance on the validation sets of the six crystalline databases and the searched value is $0.3$.
+To ensure the convergence of the TraceGrad method, we set the maximum training epochs to $5,000$.
 Other hyper-parameters and configurations are the same as DeepH-E3 (Gong et al., 2023): the initial learning rates for experiments on the $MG$, $MM$, $BG$, $BB$, $BT$, and $BS$ databases are set to $0.003$, $0.005$, $0.003$, $0.005$, $0.004$, and $0.005$, respectively; the training batch size is set as $1$; the optimizer is chosen as Adam; the scheduler is configured as a slippery slope scheduler.
-F Implementation Details on the QH9 Benchmark Series
+F.2 Implementation Details on the QH9 Benchmark Series
 The software environment used is Pytorch 1.11.0 for experiments on the two molecular databases from the QH9 benchmark series.
 When combining the proposed TraceGrad method with the QHNet architecture, the implementation of QHNet is based on the project  [^3] https://github.com/divelab/AIRS provided by Yu et al.
-(2023b), keeping the architecture and hyperparameter configurations consistent with their setup.
+(2023b), keeping the architecture and network configurations consistent with their setup.
 In our framework, we use the same number of encoding modules as QHNet: $5$ node feature encoding modules and $2$ edge feature encoding modules.
@@
 The neural network module $s_{nonlin}(\cdot)$ within $g_{nonlin}(\cdot)$ is implemented as a three-layers fully-connected module: the input size is set to $1024$, consistent with $\textbf{u}^{(k)}$, the hidden layer size is also $1024$, with SiLU as the non-linear activation function and LayerNorm as the normalization mechanism, and the output layer size (i.e., the dimensionality of $\textbf{z}^{(k)}$) is set to be equal to the number of basic blocks for a Hamiltonian matrix, which is $36$ for $QS$ and $QD$ databases.
-The SO(3)-equivariant decoder we adopt is the same as that of QHNet; the SO(3)-invariant decoder we adopt is a four-layers fully-connected module: the input size is $2$ $(K)$ times of the dimensionality of $\textbf{z}^{(k)}$, e.g., $72$ for $QS$ and $QD$; the hidden layers have $1024$ neurons with SiLU as the non-linear activation function and LayerNorm as the normalization mechanism; the size of the output layer is the number of basic blocks for a Hamiltonian matrix.
+The SO(3)-equivariant decoder we adopt is the same as that of QHNet; the SO(3)-invariant decoder we adopt is a four-layers fully-connected module: the input size is $2$ $(K)$ times of the dimensionality of $\textbf{z}^{(k)}$, e.g., $72$ for $QS$ and $QD$; the hidden layers have $1024$ neurons with SiLU as the non-linear activation function and LayerNorm as the normalization mechanism; the size of the output layer is the number of basic blocks for an atomic pairwise Hamiltonian matrix.
 Regarding the error metric in the loss function Eq.
-7, for the first term, we follow QHNet to use a combination of MSE and MAE; for the second term, we choose between MSE and MAE based on performance on the validation sets, ultimately selecting MAE.
+(3), for the first term, we follow QHNet to use a combination of MSE and MAE; for the second term, we choose between MSE and MAE based on performance on the validation sets, ultimately selecting MAE.
 $\lambda$ in the training loss function is set according to parameter selection on the validation sets, searching from $\{0.1,0.2,...,1.0\}$.
 Here, we aim to obtain a more general parameter setting for $\lambda$ on molecular structures, and thus we determined $\lambda$ based on the overall performance on the validation sets of the two molecular databases and the searched value is $0.2$.
-Other hyper-parameters and configurations are the same as QHNet: the initial learning rates for all experiments are set as $5\times 10^{-4}$, the training batch size is set as $32$, the optimizer is selected as AdamW, and a learning rate scheduler is implemented.
-To be more specific, the scheduler gradually increases the learning rate from 00 to a maximum value of $5\times 10^{-4}$ over the first $1,000$ warm-up steps.
+Other hyper-parameters and configurations are the same as QHNet: the maximum training steps are set as $300,000$ for $QS$ and $260,000$ for $QD$, the initial learning rates for all experiments are set as $5\times 10^{-4}$, the training batch size is set as $32$, the optimizer is set as AdamW, and a learning rate scheduler is implemented: the scheduler gradually increases the learning rate from 00 to a maximum value of $5\times 10^{-4}$ over the first $1,000$ warm-up steps.
 Subsequently, the scheduler linearly reduces the learning rate, ensuring it reaches $1\times 10^{-7}$ by the final step.
@@
 As shown in Fig.
-2, each Hamiltonian matrix consists of numerous basic block, with each basic block representing the direct product of two degrees.
-Here, we follow Yin et al.
-(2024) to measure the MAE performance of deep models on each basic block, denoted as $MAE^{H}_{block}$.
+2, each Hamiltonian matrix consists of numerous basic blocks, with each basic block representing the direct product of two degrees.
+Here, we follow (Yin et al., 2024) to measure the MAE performance of deep models on each basic block, denoted as $MAE^{H}_{block}$.
 The values of $MAE^{H}_{block}$ for the two setups, i.e., DeepH-E3 and DeepH-E3+TraceGrad, on different blocks of the Hamiltonian matrix for six databases from the DeepH benchmark series are illustrated in Fig.
@@
 Figure 5: Visualization of $MAE^{H}_{block}$ on each basic block of the Hamiltonian matrices for the Monolayer Graphene ($MG$) and Monolayer MoS2 ($MM$) databases.
-Figure 6: Visualization of $MAE^{H}_{block}$ on each basic block of the Hamiltonian matrices for the non-twisted (marked with superscripts $nt$) and twisted (marked with superscripts $t$) testing subsets of Bilayer Graphene ($BG$), Bilayer Bismuthene ($BB$), Bilayer Bi2Te3 ($BT$), and Bilayer Bi2Se3 ($BS$).
-H Ablation Study
+Figure 6: Visualization of $MAE^{H}_{block}$ on each basic block of the Hamiltonian matrices for the non-twisted (marked with superscripts $nt$) and twisted (marked with superscripts $t$) testing subsets of the Bilayer Graphene ($BG$), Bilayer Bismuthene ($BB$), Bilayer Bi2Te3 ($BT$), and Bilayer Bi2Se3 ($BS$) databases.
+H Ablation Study and Comparison between the Proposed Gradient-based Mechanism with the Gated Activation Mechanism
 We conduct fine-grained ablation study on the six databases from DeepH benchmark series, comparing results from the four setups:
@@
 • DeepH-E3+Trace: this experimental setup, an ablation term, only implements half part of our method.
-Specifically, it extends the architecture of DeepH-E3 by adding our SO(3)-invariant encoding and decoding branches and using the trace quantity $\mathbf{T}=\text{tr}(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})=\text{tr}(\mathbf{H}^{l_{p}\otimes l_{q}}\cdot(\mathbf{H}^{l_{p}\otimes l_{q}})^{\dagger})$ to train them.
-As for ablation study, this setup does not include the gradient-based mechanism delivering non-linear expressiveness from SO(3)-invariant features to encode SO(3)-equivariant features; instead, it directly uses the SO(3)-equivariant features outputted by DeepH-E3 for Hamiltonian regression.
+Specifically, it extends the architecture of DeepH-E3 by adding our SO(3)-invariant encoding and decoding branches and using the trace quantity $\mathbf{T}=\text{tr}(\mathbf{H}\cdot\mathbf{H}^{\dagger})=\text{tr}(\mathbf{H}^{l^{i}_{p}\otimes l^{j}_{q}}\cdot(\mathbf{H}^{l^{i}_{p}\otimes l^{j}_{q}})^{\dagger})$ to train them.
+As for ablation study, this setup does not include the gradient-based mechanism delivering non-linear expressiveness from SO(3)-invariant features to encode SO(3)-equivariant features; instead, it directly uses the SO(3)-equivariant features outputted by the SO(3)-equivariant encoders of DeepH-E3 for Hamiltonian regression.
 In this configuration, the SO(3)-invariant branches only contribute indirectly during the training phase by backpropagating the supervision signals from the trace quantity to the earlier layers.
@@
 • DeepH-E3+TraceGrad: this is a complete implementation of our framework extending beyond of the architecture and training pipeline of DeepH-E3, at the label level, we introduce the trace quantity to guide the learning of SO(3)-invariant features; Meanwhile, at the representation level, we leverage the gradient operator to yield SO(3)-equivariant non-linear features for Hamiltonian prediction.
-Table 6: Ablation study MAE results (meV) on the Monolayer Graphene ($MG$) and Monolayer MoS2 ($MM$) databases. $\downarrow$ means lower values of the metrics correspond to better accuracy.
+Table 5: MAE results (meV) of multiple experimental settings on the Monolayer Graphene ($MG$) and Monolayer MoS2 ($MM$) databases.
+$\downarrow$ means lower values of the metrics correspond to better accuracy.
 Methods $MG$ $MM$
@@
 DeepH-E3+Trace 0.230 0.344 0.348 0.378 0.537 1.091
+DeepH-E3+Gate 0.232 0.340 0.351 0.364 0.527 1.086
 DeepH-E3+Grad 0.185 0.269 0.258 0.308 0.453 0.924
+DeepH-E3+TraceGate 0.203 0.295 0.305 0.346 0.491 0.912
 DeepH-E3+TraceGrad 0.175 0.257 0.228 0.285 0.412 0.808
-Table 7: Ablation study MAE results (meV) on the Bilayer Graphene ($BG$), Bilayer Bismuthene ($BB$), Bilayer Bi2Te3 ($BT$), and Bilayer Bi2Se3 ($BS$) databases.
+Table 6: MAE results (meV) of multiple experimental settings on the Bilayer Graphene ($BG$), Bilayer Bismuthene ($BB$), Bilayer Bi2Te3 ($BT$), and Bilayer Bi2Se3 ($BS$) databases.
 The superscripts $nt$ and $t$ respectively denote the non-twisted and twisted subsets.
@@
 DeepH-E3+Trace 0.362 0.417 0.593 0.251 0.401 0.480
+DeepH-E3+Gate 0.368 0.441 0.626 0.241 0.405 0.602
 DeepH-E3+Grad 0.320 0.356 0.511 0.222 0.389 0.446
+DeepH-E3+TraceGate 0.354 0.403 0.580 0.239 0.388 0.464
 DeepH-E3+TraceGrad 0.291 0.323 0.430 0.198 0.372 0.406
@@
 DeepH-E3+Trace 0.259 0.285 0.928 0.429 0.570 1.782
+DeepH-E3+Gate 0.268 0.301 0.991 0.450 0.593 2.276
 DeepH-E3+Grad 0.243 0.272 0.824 0.406 0.542 1.431
+DeepH-E3+TraceGate 0.252 0.279 0.908 0.417 0.561 1.740
 DeepH-E3+TraceGrad 0.226 0.256 0.740 0.384 0.503 1.284
@@
 DeepH-E3+Trace 0.406 0.462 1.239 0.784 0.812 4.520
+DeepH-E3+Gate 0.423 0.469 1.196 0.813 0.829 4.494
 DeepH-E3+Grad 0.342 0.365 0.750 0.742 0.786 4.463
+DeepH-E3+TraceGate 0.368 0.377 0.982 0.778 0.783 4.489
 DeepH-E3+TraceGrad 0.295 0.312 0.718 0.735 0.755 4.418
@@
 DeepH-E3+Trace 0.382 0.397 0.843 0.351 0.367 0.838
+DeepH-E3+Gate 0.385 0.401 0.852 0.369 0.373 0.820
 DeepH-E3+Grad 0.343 0.365 0.696 0.324 0.339 0.746
+DeepH-E3+TraceGate 0.364 0.386 0.765 0.337 0.348 0.798
 DeepH-E3+TraceGrad 0.300 0.332 0.644 0.291 0.302 0.674
-Experimental results of the four setups are listed in Table 6 and 7.
+Furthermore, to compare our proposed gradient-based mechanism with the gated activation mechanism (Weiler et al., 2018) which is widely used in SO(3)-equivariant neural networks (Gong et al., 2023; Liao & Smidt, 2023), the following experimental setups are also included:
+• DeepH-E3+Gate: This variant modifies the experimental setup from DeepH-E3+Grad by replacing the gradient mechanism, which constructs equivariant features as $\mathbf{v}=\frac{\partial z}{\partial\mathbf{f}}$, with the gated activation mechanism, constructing equivariant features as $\mathbf{v}=z\cdot\mathbf{f}$.
+All other aspects remain the same.
+• DeepH-E3+TraceGate: This variant modifies the experimental setup from DeepH-E3+TraceGrad by replacing the gradient mechanism ($\mathbf{v}=\frac{\partial z}{\partial\mathbf{f}}$) with the gated activation mechanism ($\mathbf{v}=z\cdot\mathbf{f}$).
+All other aspects remain the same.
+Experimental results for all of the six setups are listed in Table 5 and 6.
 Table 5 presents the results for monolayer structures, while Table 6 focuses on bilayer structures.
-From the results of ablation terms, we can obtain a more fine-grained experimental analysis.
-By comparing among the results of DeepH-E3, DeepH-E3+Trace, DeepH-E3+Grad, and DeepH-E3+TraceGrad, we can conclude that the two core mechanisms of our method, i.e., the SO(3)-invariant trace supervision mechanism (Trace) at the label level as well as the gradient-based induction mechanism (Grad) at the representation layer, can contribute to the performance individually.
+We have the following observations:
+First, by comparing among the results of DeepH-E3, DeepH-E3+Trace, DeepH-E3+Grad, and DeepH-E3+TraceGrad, we can conclude that the two core mechanisms of our method, i.e., the SO(3)-invariant trace supervision mechanism (Trace) at the label level as well as the gradient-based induction mechanism (Grad) at the representation level, can contribute to the performance individually.
 Moreover, their combination provides even better performance.
@@
 The value of such complementarity has been fully demonstrated in the experimental results.
-I Discussion on Time Costs
-When the atomic system is small and any two atoms are within the cutoff radius of each other such that the Hamiltonian is defined, the number of Hamiltonian matrices is approximately proportional to $N^{2}$, where $N$ is the number of atoms in the system.
-Yet, as the atomic system grows, the system exhibits a dual nature of locality and non-locality.
-Locality refers to the fact that, when $N$ becomes large, the number of atoms within the cutoff radius of each atom remains roughly constant, and thus the growth in the number of Hamiltonian matrices asymptotically scales as $O(N)$.
-On the other hand, the non-locality arises from the fact that, although Hamiltonian matrices are defined only for pairs of atoms within each other’s cutoff radius, atoms outside the cutoff radius may still exert long-range effects on the Hamiltonian matrices within the cutoff radius
-The baseline models we select, whether DeepH-E3 (Gong et al., 2023) or QHNet (Yu et al., 2023b), are SO(3)-equivariant graph neural network models with efficient information aggregation and message-passing mechanisms, which cleverly balance the locality and non-locality of the system.
-As a result, the computational complexity asymptotically scales as $O(N)$ as $N$ grows to large.
-This is a significant efficiency advantage compared to the $O(N^{3})$ computational complexity of traditional DFT methods, enabling efficient and low-cost prediction of electronic structures for large-scale atomic systems.
-Our method improves the Hamiltonian prediction accuracy of DeepH-E3 and QHNet while maintaining this advantage of asymptotic complexity.
-The neural module $g_{nonlin}(\cdot)$ we introduce efficiently updates each edge feature $\textbf{f}^{(k)}$ $(1\leq k\leq K)$ from the baseline models to obtain the corresponding non-linear SO(3)-equivariant feature $\textbf{v}^{(k)}$, ensuring that the computational complexity remains asymptotically $O(N)$, with good potential to handle large atomic systems at acceptable computational costs.
-We test the inference time costs of the DeepH-E3 and DeepH-E3+TraceGrad models on an Nvidia RTX A6000 GPU, using the testing sets from two databases, namely $MG$ and $MM$, as examples; and test the inference time costs of the QHNet and QHNet+TraceGrad models on $QS$ and $QD$, on the same device.
-The batch size is set to 1 for all tests on inference time.
-Each sample is inferred 50 times, and the average time per inference is recorded.
-When using DeepH-E3, the average inference time per sample (s/sample) for $MG$ and $MM$ are as follows: $0.247$ and $0.256$, respectively; with DeepH-E3+TraceGrad, the average inference times are $0.264$ and $0.274$, correspondingly.
-When using QHNet, the average inference time per sample (s/sample) for $QS$ and $QD$ are as follows: $0.233$ and $0.174$, respectively; with QHNet+TraceGrad, the average inference times are $0.248$ and $0.187$, correspondingly.
-As we can see, the increases in computational costs due to the TraceGrad module are slight relative to that of the baseline models.
-Considering the significant accuracy improvements brought by the TraceGrad method as reported in Section 6.2, we believe that this minor increase in computational cost is acceptable and worthwhile in practical applications.
-J Empirical Study on Combining our Method with Approximately SO(3)-equivariance Framework
+Second, it is evident that both DeepH-E3+Gate and DeepH-E3+TraceGate show improvement over the baseline DeepH-E3, as these configurations introduce an additional neural network module $s_{nonlin}(\cdot)$ in learning the feature $z$, which increases model capacity.
+Moreover, DeepH-E3+TraceGate also incorporates our proposed trace supervision signal, which guides the learning of SO(3)-invariant features.
+However, their performance falls short of DeepH-E3+Grad and DeepH-E3+TraceGrad, respectively, indicating that the gated activation mechanism may not fully capture the system’s non-linearity patterns, whereas the proposed gradient-based mechanism demonstrates stronger generalization performance on Hamiltonian prediction and may be a better choice in terms of expressive capability.
+I Theoretical Analysis on Computational Complexity
+The total number of non-zero Hamiltonian matrix elements to be calculated is proportional to the number of local atomic pairs in the system, with a complexity of $\mathcal{O}(N\overline{E})$, where $N$ is the total number of atoms and $\overline{E}$ is the average number of neighboring atoms within the cutoff radius per atom.
+Since the atomic orbital basis set has a finite range, the non-zero matrix elements vanish beyond a certain distance for an given atom.
+In small systems, where all atoms lie within each other’s cutoff radius, $\overline{E}$ scales with $N$, resulting in a total number of non-zero elements proportional to $N^{2}$.
+However, in sufficiently large systems, the finite range of the atomic orbitals ensures that $\overline{E}$ remains constant, independent of $N$.
+As a result, for large atomic systems, the total number of non-zero elements simplifies to $\mathcal{O}(N)$.
+The baseline models we select, whether DeepH-E3 or QHNet, are SO(3)-equivariant graph neural network models with efficient information aggregation and message-passing mechanisms.
+These models cleverly balance the locality of Hamiltonian definitions with the long-range interactions present in the system.
+As a result, the computational complexity asymptotically scales as $\mathcal{O}(N)$ as $N$ increases, which is consistent with the growth of the scales of non-zero Hamiltonian matrix elements.
+The proposed TraceGrad method directly updates each SO(3)-equivariant feature of the baseline models, and the computational amount is proportinal to the number of features of the baseline models.
+Therefore, combining TraceGrad, the computational complexity also scales as $\mathcal{O}(N)$.
+Traditional DFT methods require $T$ iterations of diagonalizing $N\times N$ matrices, each with a time complexity of $\mathcal{O}(N^{3})$, because all occupied states are needed to compute the charge density.
+As $N$ increases, this cubic complexity leads to significant computational overhead, making it challenging to simulate large atomic systems within a reasonable time frame.
+In contrast, our deep learning framework enables the efficient and accurate prediction of the Hamiltonian for large atomic systems with a linear time complexity of $\mathcal{O}(N)$, eliminating the need for self-consistent iterations.
+Moreover, since most physical properties, such as transport, optical, and topological properties, depend only on the energy bands near the Fermi level, it is unnecessary to solve for the eigenfunctions of all occupied states once the Hamiltonian is known.
+Since the Hamiltonian matrix is sparse and only a limited number of bands near the Fermi level are needed, these eigenstates can be efficiently computed using methods like the shift-invert approach available in the ARPACK package (Lehoucq et al., 1998), with a computational complexity of $\mathcal{O}(N)$.
+J A Joint Discussion on GPU Time Costs and Performance Gains
+We here provide a joint comparison of the GPU time cost and corresponding accuracy of different models across four databases as representative: Monolayer Graphene ($MG$), Monolayer MoS2 ($MM$), QH9-stable ($QS$), and QH9-dynamic ($QD$).
+This comparison includes the average inference time per sample for each model, with the test batch size set as 1 and hardware environment set as Nvidia RTX A6000 GPU in single-task mode without computational sharing with other processes.
+For $MG$ and $MM$, we compare among DeepH-E3, DeepH-E3+TraceGrad, and DeepH-E3×2, where DeepH-E3×2 refers to a model obtained by doubling the number of encoding blocks in DeepH-E3 and training it from scratch until convergence.
+For $QS$ and $QD$, we compare among QHNet, QHNet+TraceGrad, and QHNet×2, where QHNet×2 refers to a model obtained by doubling the number of encoding blocks in QHNet and training it from scratch until convergence.
+The experimental results are documented in the Table 7.
+Table 7: Average inference time (denoted as Time, in seconds) and MAE performance ($MAE^{H}_{all}$, in meV) for testing samples from the Monolayer Graphene ($MG$), Monolayer MoS2 ($MM$), QH9-stable ($QS$), and QH9-dynamic ($QD$) databases.
+$\downarrow$ indicates that lower values of the metrics correspond to better accuracy.
+All models are tested individually on a single Nvidia RTX A6000 in single-task mode.
+Methods $MG$ $MM$
+$Time$ ()($\downarrow$)( ↓ ) $MAE^{H}_{all}$ ()($\downarrow$)( ↓ ) $Time$ ()($\downarrow$)( ↓ ) $MAE^{H}_{all}$ ()($\downarrow$)( ↓ )
+DeepH-E3 (Baseline) 0.247 0.251 0.256 0.406
+DeepH-E3×2 0.483 0.244 0.510 0.387
+DeepH-E3+TraceGrad 0.264 0.175 0.274 0.285
+Methods $QS$ $QD$
+$Time$ ()($\downarrow$)( ↓ ) $MAE^{H}_{all}$ ()($\downarrow$)( ↓ ) $Time$ ()($\downarrow$)( ↓ ) $MAE^{H}_{all}$ ()($\downarrow$)( ↓ )
+QHNet (Baseline) 0.233 1.962 0.174 4.733
+QHNet×2 0.497 1.845 0.385 4.532
+QHNet+TraceGrad 0.248 1.191 0.187 2.819
+From this Table, we find that adding the TraceGrad module results in only a slight increase in inference time compared to the baseline models.
+Given the substantial accuracy improvements introduced by the TraceGrad method, we consider this minor increase in computational time acceptable for practical applications.
+In contrast, simply increasing the depth of DeepH-E3 or QHNet results in a significant rise in inference time but yields only limited accuracy improvements.
+In contrast, DeepH-E3+TraceGrad demonstrates significantly better accuracy performance than DeepH-E3×2, and similarly, QHNet+TraceGrad achieves notably higher accuracy than QHNet×2.
+Furthermore, the inference time of DeepH-E3+TraceGrad and QHNet+TraceGrad are both lower than their respective DeepH-E3×2 and QHNet×2 counterparts.
+These results underscore the superiority of the TraceGrad method in enhancing expressive capability and improving accuracy performance, while maintaining time efficiency.
+K Acceleration Performance for the Convergence of Traditional DFT Algorithms
+Despite the increasing ability of deep learning models to independently handle more electronic-structure computation tasks, there are still applications with extremely high numerical precision requirements and very low tolerance for error, where traditional DFT algorithms must perform the final calculations.
+In such cases, the predictions from deep models can be used as initial matrices to accelerate the convergence of traditional DFT algorithms.
+We evaluate the acceleration performance brought by the proposed method for the convergence of classical DFT algorithms implemented by PySCF (Sun et al., 2018).
+Specifically, we adopt the two groups of metrics on acceleration performance:
+The first group of metrics are defined by Yu et al. (2023a), as follows:
+  • Achieved ratio.
+This metric calculates the number of DFT optimization steps taken when initializing with the Hamiltonian matrices predicted by the deep model compared to using initial guess methods like minao and 1e.
+  • Error-level ratio.
+This metric measures the number of DFT optimization steps required, starting from random initialization, to reach the same error level as the deep model’s predictions, relative to the total number of steps in the DFT process.
+Table 8: The acceleration ratios of QHNet and QHNet+TraceGrad for DFT calculation.
+Both models are evaluated on a set of $50$ molecules chosen by (Yu et al., 2023a), with the mean and standard deviation of the metrics across these samples reported.
+$\downarrow$ means lower values correspond to better accuracy, while $\uparrow$ means higher values correspond to better performance.
+Methods Training DFT Metric Ratio
+databases initialization
+QHNet (Baseline) QS 1e Achieved ratio $\downarrow$ $0.400\pm 0.030$
+Error-level ratio $\uparrow$ $0.620\pm 0.037$
+minao Achieved ratio $\downarrow$ $0.715\pm 0.033$
+Error-level ratio $\uparrow$ $0.406\pm 0.021$
+QD 1e Achieved ratio $\downarrow$ $0.512\pm 0.138$
+Error-level ratio $\uparrow$ $0.622\pm 0.048$
+minao Achieved ratio $\downarrow$ $0.882\pm 0.217$
+Error-level ratio $\uparrow$ $0.406\pm 0.066$
+QHNet+TraceGrad QS 1e Achieved ratio $\downarrow$ $\textbf{0.345}\pm 0.038$
+Error-level ratio $\uparrow$ $\textbf{0.685}\pm 0.037$
+minao Achieved ratio $\downarrow$ $\textbf{0.647}\pm 0.061$
+Error-level ratio $\uparrow$ $\textbf{0.466}\pm 0.035$
+QD 1e Achieved ratio $\downarrow$ $\textbf{0.440}\pm 0.101$
+Error-level ratio $\uparrow$ $\textbf{0.645}\pm 0.046$
+minao Achieved ratio $\downarrow$ $\textbf{0.761}\pm 0.167$
+Error-level ratio $\uparrow$ $\textbf{0.435}\pm 0.052$
+Table 9: Average wall time costs per sample (/s) for three experimental settings: $t1$ represents the wall time required for DFT calculation with a random initial guess like 1e of minao; $t2$ denotes the wall time for inference of deep learning methods (i.e., QHNet or QHNet+TraceGrad); $t3$ is the total time of the process including deep learning inference and the DFT calculation that uses the outputs of deep learning methods as initialization.
+Both models are evaluated on a set of $50$ molecules chosen by (Yu et al., 2023a), with the mean and standard deviation of the metrics across these samples reported.
+$\downarrow$ indicates that lower values correspond to better performance.
+In order to ensure fairness in comparison, all experimental settings including DFT calculation and deep learning inference are measured on a single thread of an Intel(R) Xeon(R) Gold 6330 @ 2.00GHz CPU in single-task mode.
+Methods Training DFT Metric Time
+databases initialization
+QHNet (Baseline) QS 1e t1 $120.896\pm 9.134$
+t2 $\downarrow$ $1.724\pm 0.025$
+t3 $\downarrow$ $48.604\pm 7.741$
+minao t1 $63.193\pm 5.335$
+t2 $\downarrow$ $1.724\pm 0.025$
+t3 $\downarrow$ $48.604\pm 7.741$
+QD 1e t1 $87.161\pm 12.075$
+t2 $\downarrow$ $1.280\pm 0.019$
+t3 $\downarrow$ $44.146\pm 9.342$
+minao t1 $51.396\pm 5.870$
+t2 $\downarrow$ $1.280\pm 0.019$
+t3 $\downarrow$ $44.146\pm 9.342$
+QHNet+TraceGrad QS 1e t1 $\downarrow$ $120.896\pm 9.134$
+t2 $\downarrow$ $1.852\pm 0.020$
+t3 $\downarrow$ $\bm{41.941\pm 6.783}$
+minao t1 $63.193\pm 5.335$
+t2 $\downarrow$ $1.852\pm 0.020$
+t3 $\downarrow$ $\bm{41.941\pm 6.783}$
+QD 1e t1 $87.161\pm 12.075$
+t2 $\downarrow$ $1.361\pm 0.010$
+t3 $\downarrow$ $\bm{39.712\pm 9.076}$
+minao t1 $51.396\pm 5.870$
+t2 $\downarrow$ $1.361\pm 0.010$
+t3 $\downarrow$ $\bm{39.712\pm 9.076}$
+Experimental results on these metrics are recorded in Table 8, where the results for the compared method QHNet are taken from Yu et al.
+(2023a), while the results of QHNet+TraceGrad, come from our experiments.
+In our experiments, the DFT calculation settings follow those in Section 4 of Yu et al.
+(2023a) (the DFT parameters and the $50$ testing samples) , except for the CPU environment, where we use a single thread of an Intel(R) Xeon(R) Gold 6330 @ 2.00GHz CPU in single-task mode.
+It is worth noting that this does not affect the fairness of the comparison, as the achieved ratio and error-level ratio measure the ratio of iteration counts rather than runtime, and differences in CPU computation times are negligible in these metrics.
+From Table 8, we could observe that the proposed TraceGrad method brings significant improvements to the baseline model QHNet on the acceleration ratio of DFT calculation, notably reducing the achieved ratio and enhancing the error-level ratio.
+The second group of metrics measures the wall time savings that deep learning methods contribute to DFT calculations, specifically quantifying the incremental time savings brought by our proposed TraceGrad method in accelerating DFT calculations.
+For a fair comparison, we report the average wall time costs of testing samples (/s) across three metrics:
+  • $t1$: The wall time required for a DFT calculation initialized with a random guess initialization, such as 1e or minao.
+  • $t2$: The wall time for inference using deep learning methods (i.e., QHNet or QHNet+TraceGrad).
+  • $t3$: The total wall time for the combined process, including both deep learning inference and the subsequent DFT calculation initialized with the deep model’s outputs.
+$t3$ here provides a more comprehensive evaluation of the actual time savings achieved when incorporating deep learning methods.
+Experimental results on these metrics are recorded in Table 9.
+Here all time-related measurements are conducted on a single thread of an Intel(R) Xeon(R) Gold 6330 @ 2.00GHz CPU, including the experiments for QHNet, which are reproduced under the same conditions to ensure fairness.
+Unlike the time measurements in Appendix J, which are performed on GPUs, all deep learning models here are evaluated on the CPU thread to maintain consistency in the comparison with DFT software.
+From the experimental results, we observe three key findings:
+  • Comparing $t2$ and $t1$, deep models are significantly faster than DFT calculations, achieving speeds tens of times greater for the testing samples.
+It is worth noting that, given that the testing samples here are all small molecular systems, deep models have already demonstrated a significant time efficiency advantage compared to DFT.
+Based on the computational complexity analysis of deep learning methods compared to DFT methods in Appendix I, it is reasonable to infer that for larger atomic systems, the disparity between $t2$ and $t1$ will expand rapidly.
+  • Comparing $t2$ values between QHNet and QHNet+TraceGrad, the additional runtime introduced by TraceGrad is relatively minor on the CPU, consistent with the GPU-based results reported in Appendix J.
+  • Comparing $t3$ and $t1$, the total runtime of using a deep model to predict initial values and then running DFT calculations is significantly lower than performing DFT calculations from a random guess initialization.
+Particularly, comparing row 4 and row 16, row 7 and row 19, row 10 and row 22, as well as row 13 and row 25 in Table 9, it can be observed that combing TraceGrad further reduces $t_{3}$, demonstrating that the time saved by TraceGrad in DFT calculations far exceeds the minimal additional time required for its inference.
+L Empirical Study on Combining our Method with Approximately SO(3)-equivariance Framework
 While non-strict SO(3)-equivariance, which may limit the depth of theoretical exploration, is not the main focus of this study aiming at bridging rigorous SO(3)-equivariance with the non-linear expressive capabilities of neural networks, considering that they remain of interest in a few numerical computation applications where precision is highlighted over strict equivariance, we also conduct empirical study combining our method with approximately SO(3)-equivariant techniques.
@@
 The experimental results in Table 10 and Fig.
-7 demonstrate that TraceGrad significantly enhances the accuracy of HarmoSE, surpassing DeepH-2 and achieving SOTA results.
+7 demonstrate that TraceGrad significantly enhances the accuracy of HarmoSE, surpassing DeepH-2 (Wang et al., 2024b) and achieving SOTA results.
 Both DeepH-2 and HarmoSE sacrificed strict SO(3)-equivariance to fully release the expressive capabilities of graph Transformers, aiming for the ultimate in prediction accuracy.
-Despite this, our method still manages to significantly exceed their accuracy, further confirming the superiority and generality of our method in learning expressive representations of physical systems.
+Despite this, our method still manages to significantly exceed their accuracy, further confirming the superiority of our method in learning expressive representations of physical systems for Hamiltonian prediction.
 Figure 7: Comparison on the $MAE^{H}_{block}$ metric for HarmoSE and HarmoSE+TraceGrad on the $MM$ database.
@@
 HarmoSE+TraceGrad 0.178 0.228 0.296
-K Future Work
-In future research, various extensions are conceivable across theoretical, methodological, and application fields:
-First, from a theoretical and methodological perspective, we can generalize SO(3) group symmetry to a broader range and more complex groups in deep learning research.
-This will enable deep learning frameworks to accommodate richer mathematical structures and physical quantities.
-Second, from the application perspective, the current methodology extends beyond predicting electronic-structure Hamiltonians and is capable of generally predicting multiple physical quantities and properties that are equivariant.
-Furthermore, as our approach offers a promising general solution to the critical dilemma between equivariance and non-linear expressiveness for the deep learning paradigm, the methodology is not confined to physical research alone and may find applications in areas such as robotics, autonomous vehicles, and motion tracking systems.
-For instance, in the vision tasks of autonomous vehicles (Wang et al., 2023), where there is a high demand for robustness against coordinate transformations of cameras, the mainstream approach has been to achieve approximate SO(3)-equivariance through data augmentation.
-However, this does not guarantee absolute reliability and safety.
-Our work holds great potential to construct deep models with high generalization performance under the premise of strict SO(3)-equivariance, potentially revolutionizing techniques in such tasks.
-Looking forward to future endeavors, the foundational and pivotal role of this work is clear.